A Complete Bibliography of Publications in Annals of Statistics (1980–1989)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 March 2019 Version 1.02

Title word cross-reference

(k − 1) [Vos80]. 2 [Bed82]. 2^{m} [Gho80, Gho91]. 2^{m}4^{n} [Wu89]. 2^{n} [Che80a]. _{ϵ} [BB86]. A [HM85b, SS89a, Stu87]. α [DuM83]. w̄ [Ker85]. C [Sam80]. C_{L} [Li86]. C_{p}, C_{L} [Li87]. χ^{2} [Por89, Loh89]. χ^{2}U [dWR87]. D [GK80, GK82]. D_{s} [LS88b, Stu80]. E [Con81, Con82]. F [Mar82a]. Γ [Mie81]. H_{0} [Neu82]. K [Gra88, Pol81, TP82, WSS86, BM87b, MP87b]. k + 1 [Hoe81]. K − S [Bha83]. L [BNJ88, CW89, LM85, Mas83, PK89, PK90, SY82, Ser84, Wel87a, Wel88a, vE83]. L^{2} [LaR82]. L^{p} [BQ89]. L_{1} [Dev83, Dev85, Obe82]. L_{2} [Nus85, Vio80]. Λ [WT80]. M [BS80a, Cam82, CP81, CW85b, CW89, Cox83, Fie82, FD82a, Hub84, JJV85, Lop89, Por84, Por85, Por86, Por91, SS87, Ser84, SCR87, Tyl87, Wel89, Wel90, Wie86, vE83]. E(X) [LS89]. L_{r} [MFPW89]. R^{m} [Sch83a]. n [Cur81, SA86]. N^{−2} [Ame84, Ame80]. N ≡ 3(mod4) [SS89a, GK82]. P [Hub84, LH82, LH83, Por84, Por85, Por91]. p^{2}/n [Por84, Por85, Por91]. p_{1} = p_{2} [SG80]. ϕ_{p} [Stu89]. ΠPS [HLS89]. q [LS88b]. R [Col83, CW89, SY82, Ser84]. R^{k} [Pes82]. R^{m} [Khm88]. S [Dav87, Lop89]. t [BS84, HS80, Mar85a, Rob76, Rob82a, Wij83]. T^{2} [Kar81b, HS80]. U [Ahm80, BGvZ86, BH89, CJV80, CV81, DT89, NP87, Spr85, WT80]. U(0, θ) [Nog88]. V [Akr86, Akr89, dWR87]. ε [CW85b]. W [LSF86]. X^{1} [LS89].

-Admissibility [BH89]. -Alternating [Bed82]. -Comparisons [BQ89]. -Contaminated [BB86]. -Contamination [CW85b]. -Contiguity [Neu82]. -Cube [Cur81, LS88b]. -Densities [Wij83]. -Differentiable [MFPW89]. -Estimated [Por86]. -Estimates [BS80a, Cam82, Dav87, Fie82, Mas83]. -Estimation [PK90, Wel90, PK89, SY82, SS87, SCR87, Wel89]. -Estimator [Tyl87]. -Estimators [Wel88a, CP81, Col83, CW85b, CW89, FD82a, Hub84, JJV85, Lop89, Por84, Wel87a, Wie86, vE83]. -Mean [Vos80]. -Means [Pol81]. -Minimax [Mie81]. -NN [BM87b]. -Norm [Obe82]. -Normalized [TP82]. -Optimal [Stu89, Con81, HM85b, LS88b, SS89a, Stu80, Stu87]. -Optimality [Con82]. -Optimum [GK80, GK82]. -Order [Ame80, Ame84]. -Processes [Wel90, NP87, Wel89]. -Sample [Gra88]. -Statistic [CV81]. -Statistics [Akr89, MP87b, Ahm80, Akr86, BGvZ86, CJV80, DT89, LM85, Ser84, Spr85]. -Sufficiency [BNJ88]. -Test [BS84, HS80, Loh89, Kar81b]. -Treatment [WSS86]. -Type [Cox83]. -Values [LH82, LH83]. -Wishart [Sam80].

Absolute [Hor88, KR81a]. Acceptable [Cas87]. Accounting [HWB82]. Accuracy [Fin85, Gra83, LS83a, Sin81a]. Accurate [Sto81]. Acknowledgement [Tie88]. Acknowledgment [Ale89a, CLS86]. Action [Lia88, Wij85, Wij93]. Adapting [Car82]. Adaptive [Abr84, AdF80, BNR83, Bic82, BZ80, HW85, HM87b, HW89, Hud85, Hud86, Kre87, Lai87, Mul89, PK89, PK90, Ruk82, Sch85a, Wei87a, Wei87b]. Addendum [GM89, Neu88, Sil80, SV80b, Wal80]. Additional [Mam86]. Additive [Bre89, BHT89a, BHT89b, CGW89, Cox89, DS83, ES89, FL87, GG89, GK89, HS89a, KA89, Sto85, Sto86, Tit89]. Additive-Probability [FL87]. Additivity [LS80]. Adequacy [NJ85]. Adjustment [Hab84, Hab86]. Adjustments [RS87]. Admissibility [BBZ82, Bro80b, BCSC83, Bro88, BH89, CK85, CM89a, Gaf87, GS81a, Joh84, Koz82, LaM82, MS87a, Mar83a, MG81, MGSV89, Rie81a, Ruk86, Zam81]. Admissible [BF85a, Bru80, But81, Cli88, FKZ89, Hwa82b, Kub87, Man84, MGV85, Pie85a, RR81, SA86, Sto81, VM84]. Admits [BLR82]. Advances [Gre89]. Affine [BNB83a, Vil82]. after [Tak87]. Against [HIP85, HP88a, MFP85, NT88, Rie81b, RW81]. Age [FR85]. Agreeing [Wak81]. Ahead [CM87]. Akaike [Kni89a]. Algebraic [Mal83]. Algorithm [Boh82, DR82a, Lee83, Wu83]. Algorithms [RW80]. Aligned [Tar81]. Allan [BG82]. Allocation [HW89, Lai87]. Almost [Dev81a, Sen80, Tar81]. Alternating [Bed82]. Alternative [BS84, MFP85, NT88, Tie87, Tie88]. Alternatives [HIP85, Mar82b, MS83, MS92, Tyl83a]. Amalgamation [GM87a, GM89]. Amemiya [Dav84, Dav85b]. Amount [HM87a]. Analogs [Bro88]. Analysis [And87, AA88, ABJ83, BS86, Bai87, Ber86a, BB86, Bri87, Bro84b, CR86, Chi81, Dah88, DGG81, DG81, Dia87, Dia89, Efr86, Fel86, Flu86, Fre86a, GMK^{+}84a, GMK^{+}84b, Gho86, Goo85b, Gra87, Hal86a, HP88a, HH87, Hin86, Hoo83, Hoo84c, Kem87a, KSV81, LM87, LD89, Mar83a, MO86, MS87b, Nel87, Ols86, OBWS89, RP86, Rob82b, Sch81a, Sha86, SD87, ST86, Sin86, Spe87a, Spe87b, Sri86, Ste89b, SM80, Tib86, Tib88, Tju87, Tob87, Tuk87b, Tuk87a, Web86, Wij84, Wu86a, Wu86b, WO86]. Ancillarity [Sko85, SM88]. Ancillary [Kar89b, Ped81]. and/or [Goo85b]. Andersen [SP82]. Annals [Hin80a]. Anova [KR82, Wes86b, Wes89]. Anscombe [Tak87]. any [GSW80]. Application [BB84, BS80a, Che87a, CG80, CG81, DF82, DT89, Dev87, Dia89, FKZ89, Kin80, Li86, MV87, Ros89, Sch86a, SD87, Tsa86a, Wij85, Wij93]. Applications [ALP81, Ber80b, BM87b, Bjø84, Che87b, Das89, Ede88, GS81a, GS81b, Hwa82a, Hwa82b, JSV84, JJV85, JDP83, KR81a, KR82, Kun83, LW82, LPQ81, Lin89a, Mam89, MG83, MLV84, MRW86, RW82b, Sie86, Swe86b, Van87, Wan85, Wei87a, Wit83]. Applied [SW80]. Applies [Rub84]. Applying [Sch81b, SEB83]. Appreciation [Ney81]. Approach [AZ85, Ass88, BP84, Bir89, CL83, DFP89, DL81, Eub81, FA86, Fu82, Gat84, GM84, Khm88, LS88a, LT83, Mau85, MLV84, Nov80, PT81, Ras80, Sha88b, Sim83, TK86, Van87, Wij84]. Approaches [Yao84]. Approximate [BW81, Bra84, Huf83, Wan85]. Approximating [DH80, Yao87]. Approximation [AZGH82, AZGH83, CV81, Cox88a, Cox88b, DJ89, Jeg88, Mar83c, Mar86, Mar91, Muk81, Por85, Por91, Rup81, Sch83b, Wei87b]. Approximations [Ber84, Bro85, Kre86]. Arbitrary [Cso86, Gle86, Gly80, Har81b]. ARC [Woo82]. Area [Bro84c]. AREs [Loh84a]. Arguments [Sha82]. Arising [RW82a]. ARMA [Han81, Pot84, HIP85, HP88a, Han75, Han80a, Han80b, Har88, Kre87, Pot83, Tan83a, TT83]. ARMA-Models [Pot84, Pot83]. Armed [Eic88, Jos75, Jos85, Kak83, Kee85b, Kel81, Lai87]. Array [Chi86]. Arrays [Che80b, HS89b]. Article [BG82]. Aspect [LH84]. Aspects [Wij84]. Assay [Ker85]. Assessing [LR89]. Assessors [Sch89a]. Associated [Akr86, Akr89]. Association [ALP81, FR83, Goo85b, HR86, JDP83]. Associations [LW89a, LW89b]. Assuming [FPV87]. Asymmetric [Col86]. Asymmetry [Goo85b, Mac86, Mac87]. Asympotic [Fei81]. Asymptotic [AS80, AR86, AJ81, ACF82, AA88, AH88, Bai83, Bai84, BB84, Bas87, BHHW83, BN83, Ber80a, BF81b, BF84, Bic84b, Bjø84, Cap88, CW87, Che80c, CM87, Col86, Cro80, Cuz85, Dah88, Dar83, Dav87, Dav85a, Dev81b, Dev88, Dos85b, DW88, FK85, FK86, Fal85, Fie82, Flu86, Fuj85, FS89, Gil83b, Gly80, GL88, HT85a, Hal81a, Iac83, JSV84, Jan86, Joc86, Kau87, Kee85a, Ker85, KN84, LaR82, Lai81, LH82, LH83, LT84, LY88, LSF86, Leu82, Leu84, Li86, Li87, Lor83, Mas81a, MFP85, Mie86, NT88, Nis84, Nus85, Par86, Pie82, Pol86, Por89, Por84, Por85, Por86, Por88, Por91, PS83, Ran82, Rao78, Rao80, Ree85, Rie81a, RP80, Rob80, Rot81, RV80, Rus76, Rus82, SW88, Sch81a, Sch81b, SP88, Sha89, SW84, Sil85a, SCR87]. Asymptotic [Sin81a, Sin81b, Sko85, Ste89a, Ste89b, Sti85, Stu84, Swe80a, Swe82, Swe86a, Tah89, Tak87, Tan83a, TL84, Tru89, Tyl81, Tyl83a, Tyl83b, VJ87, Vio80, Wan81, WJT86, Wei85, Wel82, Wen89, Wes86b, WT82, Woo85a, WT85, Woo89a, Wri81a, Wu81a, Zeh81, vE83, vdV89]. Asymptotically [Bha86, ESS82, Ham81, Kla87, Kog87, Mar85b, MP87a, SY81, Sch86b, Shi80, Ste88, Sto84, Swe86b, Wan86b, Zer80]. Asymptotics [Cox83, DF84, Mam89, Mor86, Neu87, Neu88, Sen88]. Attainable [HW84]. Attainment [MFPW89]. August [Ney81]. Autocorrelation [AZGH82, AZGH83, Sak83]. Automatic [DP84a, DL88a]. Autoregression [AZGH82, AZGH83, Cro80, Kou86, KL89]. Autoregressions [Bos88]. Autoregressive [AZGH82, AZGH83, BB84, BMMT89, Bha82, Bha86, Cam82, CW88a, FD89, HD89, Kaw80, KW84, Kni89a, Kun84b, LS83a, Pot89, Shi81, Sri87, SS89b, TT83, Tsa84, Wan86a]. Average [ACF82, BMP85, God80, Hal84b, MF83, Nai84a, Oeh85, Pol87]. Averaged [Sco85]. Averages [DR86].

Bahadur [Bas80, Bas83, Epl81, Kou84a, Pra88, Wel86]. Balance [HS83, JN83]. Balanced [CW80b, CW83, Che83, KW81, KR81b, Kun84a, SM80]. Ban [BB83]. Bandit [Eic88, Jos75, Jos85, Kee85b, Lai87]. Bandits [CB85, CW88b, Kak83, Kel81]. Bands [Nai86, Pie85a]. Bandwidth [Abr82, HM87a, HM85a, Mar85b, MP87a, MS87c, Ric84]. Based [AG79, AG80, BLR82, Bic84c, Bra84, Bro85, Chr89, Daw85a, Daw85b, DJ89, Fou80, Gra83, Hab81, Hal89a, HP88a, Han80b, Hen88, LM85, Mas81b, Mil89, OT81, Pha80, Sch85b, Sch83a, Sch80, SS81, SV80b, Tor81, Tru89, Wah81, Wel86]. Basis [Bar82, DS82a, DS82b, Fra82]. Batch [LT83]. Bayes [HH83a, Bar86a, Ber80c, Ber86c, BB86, BBC81a, BBC81b, BR82, BL89, But81, Cla86, CS84a, Cre82, DH80, Daw86, DF86c, DF86a, DF86b, DL86, Dos86, FD83, GBT82, Haf80, HH83b, Har86, Hjo86, Kar88, KP86, Le 86, Ler86, Lia88, Lin86a, Lin85, LC87, MY84, MD84, Mar83b, MM87, McK84, Nog88, Nov80, Pha80, Pos86, Rit85, Rob83, Sha82, SW80, Sim86, Sti85, Str81, Yao84, Zeh81, vH87]. Bayesian [Cha85, SV80b, Amm84, AZ85, AR89, Bar86b, Cha84, CB85, Cla85, DH81, Dos85a, Dos85b, DL81, Gen84, GS85a, GMS86, Gol81, GC87, Kra84, Kuo83, Lo84b, Lo86a, Lo87, Lo88, Ras80, Rub81, Sha88b, Tsa86a, Was89, WZ83, Wen89, Wit86]. Bayesianly [Rub84]. Be [Ruk83, Bic84a]. Bechhofer [KJ86]. Before [Fin85]. Behavior [DW88, Jur81, Kee83, Kee87a, Leu84, Por84, Por85, Por86, Por88, Por91, Sil85a, Sin81b, Wri81a]. Behaviour [Dav87, Gil83a, Kre86, Sti85]. Behrens [Rob82a, Chi81, DF82, Rob76]. Belief [Wal87]. Belsley [Atk84]. Berger [Bro80a]. Berkson [GS81a]. Bernoulli [Jos85, Jos75, Kel81, KJ86]. Berry [Jos85, Ahm80, Che81b, Fri89, Jos75]. Best [ASZ85, BK80, BK81, EO87, Fox81, FGP88, GM87b, HW84, Hsu81, Hsu84, KA83, Pet80, Tak81, Tak82, TWZ80]. Beta [WT80]. Better [CHL83a, CHL83b, Kum87]. Between [Bai84, BK80, BDD89, Dos89, Kar81c, LW89a, MM87, MG83, RW81, Swe80b, vE83, BQ89, HP88b, LW89b, Lop89].

Bias [Abr84, AS80, DS89, MYZ89, Sha88a, SS89b, Var82b, Var85]. Biased [GVW88, Har81a, Hec85a, Hec85b, Smi84]. Biases [Bic84a, Ros89]. Binary [Bed82]. Binomial [CS84b, Hwa82a, Hwa82b, SA86]. Binormality [Szk88]. Bio [Ker85]. Bio-Assay [Ker85]. Bioassay [GS81a, Kuo83]. Birnbaum [BG82]. Bivariate [Dab86, DK80, Sni81, TLC86]. Block [BY80, Che83, Con81, HM85b, HS83, KW81, MN83, MC88, Stu87, Tar81, KP81]. Blocks [Con82]. Blum [RS85]. Bona [Gaj86]. Books [Atk84, Sil85b, Wij84]. Bootstrap [AS85, Ath87, BS83a, BO88, Ber82a, Ber84, BS85a, Ber86a, BS87, Ber88, BQ89, BF81b, BF84, Bic88, Bos88, BGL88, BD84, CR86, DR88, Efr86, Efr88, Fel86, Fre86a, Gho86, GZ89, Hal86b, Hal86a, Hal86c, Hal88a, Hal88b, HDR89, Hin86, Kni89b, Kun89, LS87, Liu88, LS88c, Lo87, Lo88, Loh88, Mam89, MO86, Ols86, OBWS89, RP86, Rob88, Rom89, Rub81, Sha86, ST86, Sin81a, Sin86, Sri86, Tib86, Tib88, Vea88, Web86, Wen89, Wu86a, Wu86b, WO86]. Bootstrapped [HY87]. Bootstrapping [BMMT89, CM89b, Fre81, Fre84, GPSB84, Nav89, Sau89]. Boscovich [KB85]. Both [Hal83c, Jen88]. Bound [BK80, Fri89, Fuj85, Hu85, MFPW89, Tah89, TK86, Yat88, vE83, vdV89]. Boundary [Fer82a, Rao78, Rao80, Sie86]. Bounded [BBC81a, CS81, CM82, GSW80, Li84b]. Bounds [Ale80, Ale82, AG79, AG80, Ber80a, BMWZ87, DP84b, Dev88, FS89, Gor89, Hwa82b, JBB83, Loh84a, Loh89, MY84, Nai84a, Nai84b, RW82a, SS81]. Bowl [GSW80]. Bowl-Shape [GSW80]. Boxes [HLS89]. Bracketing [Hos89]. Breakdown [Hub84, Yoh87]. Breakdown-Point [Yoh87]. Broad [Hwa85]. Broken [DG80]. Brown [Kil81]. Brown-Mood [Kil81]. Brownian [AR89]. Buyer [Mal87].

C [Tib88]. Calculation [ES81b]. Calculations [Bra84, Rub84]. Calculus [Koz82]. Calibration [Daw85a, Daw85b, Oma88, Sch85b]. Calibration-Based [Daw85a, Daw85b, Sch85b]. can [Bic84a, Ruk83]. Cancer [BBMT81]. Canonical [Gly80, Hab81, HP88b, JB83, JBB83, LS85b, LM87, YB80]. Capacities [Bed82]. Carl [Sac84]. Carlo [HM87b, Joc86]. Carriers [Sil85a]. Case [Ath87, Atw80, Fer82b, Fox81, FD83, GBT82, Kel81, Kni89b, Lia88, Lo84a, Mie86, MLV84, MV87, SP88]. Case-Cohort [SP88]. Case-Control [MLV84, MV87]. Cases [BS89, Hwa82a, VJ87]. Categorical [Kau87]. Categories [Goo85b]. Cauchy [Ree85, RW80]. Cell [RS84]. Cells [Kog87]. Censored [Akr88, CY87, CHL83a, CHL83b, Cuz85, Dab86, FR81, GP85, HT86, JS84, KSV81, Kum87, LV85, MP87a, Mie86, MM80, Pha80, Rei81, Slu84, SV80b, SV80a, TW83, Tan83b, VJ87, VC84, Yan83]. Censoring [Akr86, Akr89, Jac89, Leu84, MGSV89, Rot85, TLC86, Wel85]. Censorship [DFP89, MR88, Sch85a, Sch86a, Tsa86a]. Central [HT82, HT93, ST81]. Centrality [SA82]. Certain [BL83, Bha82, ESS82, Pie82, Rei80, Rot85, Sch83a]. Change [AR86, Ass88, AR89, Car88, GHP86, MFP85, Pol85, Pol87, Pra88, Yao87]. Change-Point [AR86, Car88, MFP85, Pra88, Yao87]. Changes [Mou86]. Changing [MD84]. Character [But81]. Characterisations [Let81]. Characterization [BL83, Gen84, GMS86, JM80, JM82, Kos87, Sak83, Slu82, Sma83, SS89b, WS80, vD85]. Characterizations [Ebe83, Fab73, Fab81, Gaf87]. Characterizing [Wu80]. Chart [BF81a]. Chernoff [Kal82]. Chervonenkis [SY89]. Chi [Ame80, Ame84, BB83, Ber80d, BM85, Dav85a, DE85a, DE85b, Dic85, Dro89, Fie85, GM83, Goo85a, Goo85c, HT86, Leo85, Mar82a, MP85, MM80, Moo82, Pie85b, Pla85, QR85, RS84, RS87, SA82, Sun85]. Chi-Square [Ame80, Ber80d, BM85, DE85a, DE85b, Dic85, Dro89, Fie85, Goo85a, Goo85c, HT86, Leo85, MP85, MM80, Pie85b, Pla85, QR85, RS87, Sun85]. Chi-Squared [Ame84, Dav85a, GM83, RS84]. Choice [HM88, Hau88, Pet80, Ric84]. Cholesky [EO87]. Choose [MM87]. Choosing [Mas83, MG83, Wah85a]. Chosen [DS82c]. Class [Akr86, Akr89, BBC81a, Bjø81, Bro81, BF85b, BM89b, Cho87a, CRM87, Gra88, HL84, Hwa82b, Hwa85, JSV84, KS77, KS83, Klo84, Lo84b, OT81, Por89, RR81, SW84, SG81, Ton89, Tyl83b, VC84, Zam81, vdV88]. Classes [BMWZ87, BCS80, BC81, BCS89, Dro89, Ebe83, Gaf81, Loh84a, Mar82b, SY89, VD87]. Classically [BR82]. Classification [HT89, Sch81a]. Classified [Goo85b]. Clinical [SG80, Sim86, Smi84, WSS86]. Clinical-Trials [Sim86]. Closeness [SKS89]. Closest [Sch87b]. CLT [BS80a]. Clustering [Pol81]. Cochran [Wat82]. Coefficient [DG80]. Coefficients [And51, And80, Fl80, HC86, LR82, NT88, Sak83]. Coherence [Reg87]. Coherent [AS89, HS89a, LS83b, LS85a]. Cohort [SP88]. Coin [Hec85a, Hec85b, Smi84]. Coincidences [Hen88]. Collapsibility [Dav89]. Collection [Fin85]. Collinearity [Atk84]. Combating [Abr84]. Combination [BNB88]. Combinations [Hel80, Mas81a, Sch80]. Combinatoric [Mau85]. Combining [BK83, Mar82a, Mar85a]. Comment [Tak82]. Commentary [SP82]. Comments [Dav84, Dav85b]. Common [Bha80, Flu86, Kub87, Nai80, Ton89]. Compact [ST81]. Comparative [Fis83, Wak81]. Comparing [DF82, Gra88, HM85b, Leh88, MN83, Sch89a, Stu87]. Comparison [BO88, Ber88, Bic88, BGL88, DR88, Efr88, FHH83, Hal88a, Hal88b, LS88c, Loh88, Mar87, Ped81, Rob88, SWW84, Tor81, Vea88, VD87, Wah85a]. Comparisons [BQ89, BW82b, Hay84, Hsu84, Joe88, KR82, RSC87, Vos80, WSS86, Wes89]. Competing [Gra88, LPQ81]. Competitors [Ber82a]. Complete [BCS80, Bro81, BF85b, BCS89, BM89b, Cho87a, Dev83, Gaf81, MR87a, Mar82b, VJ87, Zam81]. Completeness [SM88, SG81]. Complex [Ber89]. Complexity [Kem87b, Ris86]. Component [DFP89, Flu86, Har81a, SEB83, Tyl83a, Tyl83b]. Components [AAO86, Che88a, Kar88, Mat84, Puk81, Sch80, Wes86b]. Composite [BNJ88, Lai88]. Compound [Boh82, Ede88]. Compromise [Kem88].

Computation [Cam82, Dos85a, JBB83]. Computing [Nie81, Nie83]. Concave [BS80b, Ton82]. Concavity [Hab89]. Concentrated [Riv89]. Concept [RP80]. Concerning [Jon87]. Concordance [ACF82, FA86]. Condition [Bro80b, BCSC83, GS81a, Hos89, Sri84, Tie87, Tie88]. Conditional [Bas80, Bas83, BB84, Dab89, Daw80, Fei81, HJS88, HR86, Hoo84a, KR82, Lin83a, MC87, Smy88, Stu86, Swe86a]. Conditionally [Cas87]. Conditioning [Sha82]. Conditions [AA88, AK85, BCSC83, FK84, GZ89, Sza80]. Confidence [Ale80, Ale82, BO88, BS83b, BW81, BS85a, BM86, BS87, Ber88, Ber80c, Bic88, BGL88, DR88, Efr88, Est82, GH87, Hal86b, Hal86c, Hal88a, Hal88b, Hoo82a, Hoo82b, Hoo84b, Hoo84a, Hsu81, Hsu84, HC82, HC86, KR81a, Li89, LS88c, Loh88, MV87, Nai84a, Nai84b, Nai86, Oma88, Pie85a, Rob88, SS81, SS84, Spr85, Vea88, Wit83, Woo86, ZM84, ZM88]. Configural [Mor86]. Conformity [Puk84, RW82b]. Confounding [Col84]. Conjecture [Hay84, Jon87, Jos75, Jos85, KS84, Son82]. Connected [Ber89]. Conservative [Hay84, Nai86]. Consistency [Bar86a, Ber86c, Bjø84, CY87, CH80, CT87, Cla86, CM89a, Dab89, Dav85a, Daw86, DW80, DP84a, Dev85, DF86c, DF86b, DL86, Dos86, FK85, FK86, FR81, GKP84, HL84, HJS88, Har86, HD89, Hjo86, JS84, Klo82, KG88a, Kni89a, KP86, Le 86, Li84a, Li84b, Lin86a, Lo84a, MF83, Nel80, Obe82, Pfa82, Pol81, Red81, Sil80, Sol81b, Str81, TT83, TC85, TC90, Wan85, Wan87, Won83, Por84]. Consistent [CGW83, GS85b, Kla87, RB86, SS80, Tsa86a, WW83, Wu80]. Consists [Li84c]. Constancy [NT88]. Constant [MFP85]. Constrained [Hat85, Hsu84, KS77, KS83]. Constraints [HS89b]. Construct [Hal86c]. Constructed [CRM87, Dan82]. Construction [Ano83, Ber83a, Che81a, Che83, GK82, GHT83, HLS89, Hud83, Mor83a, Wu89]. Contained [Mam86]. Contaminated [BB86]. Contamination [CW85b, Col86, Gol82]. Contaminations [SB89]. Content [KR84]. Contiguity [Neu82]. Contingency [AG82a, CG80, CG81, DLS80, Dav89, GM87a, GC87, GM89, Goo85b, Hab81, Loh89, Mey82, RS84, Sim83]. Continued [Stu80]. Continuous [Ber80b, Bro88, Epl82, Jen87, LS83b, Mas80, Sti85, Yu89]. Continuous-Time [Mas80]. Contoured [MRW86]. Contrast [Egu83]. Contributions [Kee85b, Wyn84, Bic84b]. Control [BBZ82, Ber83b, BF81a, sC79, CW80a, HM85b, LW82, Lee88, MN83, MLV84, MV87, Sha80, Sri84, Stu87, Zam81]. Controlling [Kem88]. Convariance [Puk81]. Convergence [BM87b, Boh82, BDD89, Che88a, Che88b, Dev81a, Dev83, HW84, HY87, JJV85, Jeg87, Jon87, KG88a, KL89, KI81, Lia88, Mar83b, Nog88, NP87, NC89, PV80, Rom88, Ruk83, Sch86a, Spe85, Sto80, Sto82, Tar81, TS80, Wu83, Yat85]. Convergent [Atw80]. Converse [Boo85b, Swe86c]. Convex [Leu82, Moo84, WW80]. Convolution [KP81]. Cook [Atk84]. Correct [Sch81a]. Corrected [AS85]. Correction [Akr89, AZGH83, And80, AG80, Bel91, BS87, BG80, Bos88, BCS89, Cha85, CW80a, Cuz90, Dav85b, DR87, Efr82a, Fab81, FK86, Fre88, GMK^{+}84a, Gho91, Hab86, Han81, Hin80b, HT93, Hud86, Jos85, Kar89a, LW89a, LS87, Mac87, Mar86, Mar91, MS92, OD80, PK90, Por91, Rao80, TC80, Tib88, TC90, Wel88a, Wel88b, Wel90, Wij93, WT85, Woo87, Woo89a]. Corrections [Ale82, Ame84, Bas83, Cha82, CHL83a, CW83, CG81, Haf81, HH83a, Hoo84b, Hoo84c, KS83, LH83, Pot84, Rob82a, Ros82, Rus82, Swe82]. Correlated [Che83, KW81, Sol81b]. Correlation [Ale89a, Ale89b, DR86, DG80, Goo85b, Hab81, Kar80, LM87]. Correlations [Gly80, HP88b, JB83, JBB83, MC88]. Corresponding [BNR83]. Cost [CC88, Kem87b]. Counterexample [Ale89a, Ale89b]. Counting [AG82b, Gil86, RH83, SP82, Sve89]. Covariance [AF84, Bed88, BS85a, BS87, DR86, DS85, Eri87, Gle86, Haf79, Haf80, Haf81, Har81b, Jen88, Lop89, Mar81, Per80, Sak83, Son82, Ste88, Ste89a, Sza83]. Covariances [SWW84, Sza80]. Covariate [SC85]. Covariates [Fri82, MP80]. Coverage [Est82, Est83, Est86, Hoo88]. Cox [SP82, AG82b, AH88, Bai83, Bai84, PdT88, PS83, Sen81, Tsi81]. Cox-Type [PS83]. Cramér [Blo87, BMWZ87, CC82, MFPW89, Sch80, vE83]. Criteria [AAO86, Atw80, BB83, Gaf81, Nis84, STT80, WT80]. Criterion [Bha86, Gol81, HD89, Hoo88, Kni89a, KN84, Mas83, O’R76, OD80]. Critical [CC82]. Critique [DuM83]. Cross [BD84, CGW83, Goo85b, Hal83b, Li84a, Li85, Li86, Li87, Mar87, Won83]. Cross-Classified [Goo85b]. Cross-Validated [CGW83, Li84a]. Cross-Validation [BD84, Hal83b, Li86, Li87, Mar87, Won83]. Crossing [Sie86]. Crossings [KS82]. Cube [Cur81, LS88b]. Cumulant [Kee87a]. Cumulative [Gil86, Gra88, Lev81]. Curiosum [Sti80]. Curtailed [EG80, HS80, HS87]. Curvatures [Ama82, CG86]. Curve [FR81, HT88, SV80b]. Curved [Ama82, Egu83, Kal81, MV87]. Curves [GMK^{+}84a, GMK^{+}84b, Mul84, MS87c, Tsa86a]. Curvilinear [Nai86]. Cyclic [Vos88].

D [Atk84]. D. [Atk84]. Darroch [Csi89]. Data [Akr88, Amm84, Atk84, Ber82b, Bha81, Bri88, But82, CY87, CL88, CHL83a, CHL83b, Chi86, Dab86, DK80, Dia89, EK83, FA86, Fin85, FR81, Fri82, Gaj88, Goo85b, GP85, HT86, Hal83c, Hau88, Jac89, JS84, JS83, Ken83, KG88b, KSV81, Kum87, LH84, LV85, Mie86, MM80, NC89, Pha80, RS84, RS87, Rei81, Rit85, Sch85a, SCR87, Slu84, SV80b, SM80, Sza83, TW83, Tan83b, TC85, TC90, Van83, Wah81, Woo85b, Woo87, Yan83]. Data-Adaptive [Sch85a]. Data-Based [Wah81]. Data-Smoothing [LH84]. Deal [Nai80]. Decision [Das89, Ede88, Efr82b, Gat84, GM84, HP86, Kem88, Man81, Rit85, Ruk82, WZ83, Wit86]. Decisions [FL87]. Decomposition [EO87]. Decompositions [KR82]. Deconvolution [LR82, RR83]. Decreasing [Bir87b, vH87]. Deficiencies [Swe80b]. Deficiency [Fal84, Kal82]. Defined [Loh84a, Rup81]. Definitions [JB83]. Degree [BGvZ86]. Delayed [Eic88]. Delta [SW81b]. Demographic [BRH85]. Densities [Bir87b, Gly80, Hal81b, MRW86, Mul84, Wij83]. Density [Abr84, AdF80, BM87b, Bir87a, BDD89, BL89, CGW83, CG85, Cli88, Dev83, DP84a, DP84b, Dev87, Don88, Eub81, Gaj86, Hal83b, HM87a, Hal87, HM88, Har88, Klo82, Klo84, KI81, Mar85b, MP87a, Mar87, Mie86, Rom88, Ros75, Ros82, SY81, Sch85a, Sch86a, Sch87b, STT80, Sco85, Sil80, Sil82, Sil85b, Sin81b, Sto84, SW81b, TS80, Vio80, Wah81, Wal80, Wij85, Wij93, vH87, Lo84b]. Density-Quantile [Eub81]. Depend [BL83]. Dependence [Dos89, GM83, Koy87, Moo82, RP80, Sch87a, SW81a]. Dependent [AG79, AG80, BF89, DT89, FT86, LPQ81, RSC87, Riv84b, Ton89, Tsa86a]. Derivatives [RR83, Sil80, Sin81b]. Derived [Rie80]. Deriving [Gat84]. Description [Ris83]. Design [Atw80, Bro84c, Cha84, Cha85, sC79, CW80a, Che87b, Dav85a, Gaf87, Hec85a, Hec85b, Kee84, Kun83, Kuo83, LS80, Li84c, PT83, Wyn84, Ylv87]. Designed [HS83, Woo89b]. Designs [BS86, BY80, Bru80, CW80b, Che80a, Che81a, CW83, Che83, Con81, Con82, CC88, ESS81, ESS82, Gaf81, GK80, GK82, HM85b, HLS89, Hoe81, Hoo89, JN83, KW81, KW84, Kun83, Kun84a, KM87, LS85b, Li83, Li84c, LS88b, MN83, MC88, Pes82, SY84, SS89a, SC82, Smi84, Smy88, Spr80, Stu80, Stu89, Stu87, Tar81, VD87, Vos88, Wu81b, Wu89]. Detecting [Ass88, AR89, BF81a, Mou86, Pol87]. Detection [But81, DS86, Pol85, SM82, Sin84]. Detector [Bro84a]. Determinants [Lin89b]. Determination [HD89, Mar84]. Determined [Cur81]. Deviation [Fu82, Kal83b, Kou84b, Ros75, Ros82, SD87, Wu86c]. Deviations [Kal85, KK86, Vio80]. Diagnostic [PT82]. Diagnostics [Atk84, MLV84, Pre81]. Diagonal [Sam80]. Diagrams [Smi89]. Diameter [GH87]. Dichotomous [Dav84, Dav85b, Sim86]. Difference [BK80, FHG81, GM80, Joh84, Nel80]. Different [Kem88]. Differentiability [Cla83]. Differentiable [MFPW89]. Differential [Ama82, BNJ88, Nag82, PT83]. Differentials [BS80a]. Dimension [Cso86, Mam89, Ros82]. Dimensional [Dav89, KW84, Ros75, Sau89, Sch83b]. Dimensionality [Sto86]. Dimensions [Sco85]. Dip [HH85]. Direction [ASZ85]. Directional [JS83, Ken83, KG88b, Sha80]. Directions [JS85, WW83, Wu80]. Dirichlet [CG81, Fab81, AG82a, Bha81, Chr86, CG80, DS82c, Fab73, GC87, JM80]. Discontinuous [AJ81]. Discordancy [Swe86b]. Discount [Kel81]. Discounted [Fin83, Kak83]. Discovering [Cha81, Cha82]. Discrete [Ano83, Ber83a, Bro88, FD83, GHT83, Hud83, Hwa82a, Hwa82b, Jen87, Li87, Lia88, Mas80, Mor83a, Muk81, Nov80, NC89, STT80, SCR87, Yao84]. Discrete-Time [Mas80, Yao84]. Discriminant [Chi81, Hab84, Hab86, Sch81a, Wac80]. Discrimination [DW80, Dev81b, KS82, Mar83b]. Discussion [And87, BO88, Bai87, Bar82, Bar86a, Ber86a, Ber88, Ber83a, Ber86c, Bic88, Boo85a, Bre89, BM85, Bri86, Bri87, BGL88, BS85b, CR86, CGW89, CW85a, Cla86, Cox85, Cox89, Daw86, DD87, DR88, Dia85, Dia87, Dic85, DL86, DJRS85, Dos86, Efr86, Efr88, ES89, Fel86, Fie85, FH86, Fra82, Fre86a, Fri85, GG89, GK89, Gho86, GK85, Goo85a, Goo85c, Gra87, Hal86a, HH87, Har86, HT85b, Hin86, Hjo86, Hud83, Jon85, Kem87a, Koe87, KA89, KP86, Kun86, Le 86, Leo85, Lin86a, LS88c, Loh88, Mal85, MP85, Mil85, ML86, MO86, Mor83a, Nel87, Ols86, Pie85b, Pla85, Poo86, RP86, Rob86, Rob88, Sch85b, Sha86, She85, ST86, Sin86, Sri86, Sun85, Swi85, Tib86, Tib88, Tit89, Tju87, Tob87, Tsa86b, Tuk85b, Tuk85a, Tuk87b, Tuk87a]. Discussion [Vea88, Wah85b, Web86, Weg86, Wes86a, WO86]. Disorders [BF81a]. Dispersion [DS86, Dav87, FJ84, MP87b, ST81]. Distance [BM89a, DL88a, DL88b, KD83, Kou86, LaR82, Wan86a, Yat85]. Distances [Hsu81]. Distortion [Abr84]. Distributed [Ber82b, Cur81]. Distribution [Ale80, AA88, ABJ83, Ass88, Bai83, BM86, Ber89, Bha81, Bic81, BMP85, CS84b, CK85, CLS85, CLS86, CC82, CDM85, DK80, Dar83, DED89, DW80, DP84b, Ede88, ES81b, Fab73, Fab81, Fl80, FGP88, Fuj85, FS89, Gil83b, GM83, GCG87, Gor89, GKP84, Gup86, GL88, Hal82, HH83a, HH83b, HC82, JM80, JM82, Kee83, Kem87b, KN84, KM80, LSF86, Let81, Lev81, Lo85, Lo86b, Loh84b, Loy80, MR88, Man84, MR87b, Nai80, Nog88, Par86, Pha80, Pol85, Pol87, Por89, Por86, PS83, Rao78, Rao80, Rit85, RV80, SA82, Sch83a, Sch87b, SP88, SCR87, Ste81, Tar81, Tyl83a, Tyl87, Vio80, Wit83, Woo85b, Woo87, Yao87, Yu89, Zeh81]. Distribution-Free [DW80, DP84b, GKP84, Tyl87]. Distributions [AG82a, And51, And80, And82, BL83, Bas87, Ber82a, Ber89, BS80b, CW88a, Chi81, CG80, CG81, Cur81, DS86, Eat81, Efr82a, Efr82c, Fis82, FHW81, Gat84, GVW88, Gol82, Hat85, Hwa82b, Jeg87, Jeg88, Jew82, Kal85, KRS87, LM85, Leu82, LS88a, Loh84a, MR87a, Mou86, MFPW89, Nie81, Nie83, Red81, Riv84a, Riv84b, Riv89, RW82a, Rob82b, Rus76, Rus82, Sin84, Smi87, SK86, Ste89a, TL84, Var85, WT80, WS80, Wan86b, vH87]. Disturbances [Sol81b]. Diversity [FA86]. Divisible [Bro84a]. Domain [Chi88]. Domains [Man81]. Domination [BH89, Hwa85]. Double [CS84a, CS84b]. Doubly [CY87]. Down [Ker85]. Drift [AR89, Fer82a, MY84, SYW89]. Dual [BNB83a]. Duality [DJ89]. Dynamic [Ass88, AR89, LW82]. Dynkin [SW80].

Eaton [Wij84]. ED50 [Ker85]. Edgeworth [BG80, AS85, BS89, BG78, BQ89, BGvZ86, Bos88, CJV80, Doe83, Hal83a, Hel80, JR83, Nav89, Sch89b, Sti80]. Editorial [Hin80a, Mil80]. Effect [Gho80, Gho91, GM83, Moo82, Pie82, dWR87]. Effective [Sco85]. Effects [CD86, Dah88, FKZ89, GMT86, LR89, MP87b]. Efficiencies [QR85]. Efficiency [BS86, Bas83, BHHW83, BN83, Cap88, Egu83, Epl81, Est86, GSW80, Gre80, Joc86, Kal82, Kal83a, KL87, Kee84, Lin83a, Lin85, Lor83, Neu82, Nus85, Pra88, Rot81, SP88, Tan83a, TL84, Wu81b, Yoh87, vE83]. Efficient [Bha86, BR87, Dah89, Dio81, Huf83, KI81, LS88b, Mar85b, Sch86b, Shi80, Ste86b, Ste88]. Efron [Dev87, Ped81, Sin81a, Ste86a]. Egon [Ney81]. Eigenfunctions [KM80]. Eigenvalues [ABJ83, LM87]. Eigenvectors [Tyl81]. Element [Che80d]. Elements [WZ83]. Elliptical [Ton82].

Elliptically [DS86, MRW86, Sin84]. Ellsberg [Fis83]. Embedding [MR88]. Empiric [GM83, Por86]. Empirical [Akr86, Akr89, BB86, BQ89, CK85, Cre82, CM89b, Daw85a, Daw85b, DT89, Fou80, Fl80, GVW88, GBT82, Haf80, Kar88, KL89, Lia88, Lin85, Loy80, MD84, Mil89, Nog88, Pha80, Rob83, Sch86a, Sch85b, Sch83a, Sti85, Stu86, Var85, Wac80, Wit83, Yao84, Yu89, Zeh81, vH87]. Employ [VM84]. Emptying [HLS89]. Endpoint [Hal82, Loh84b]. Entries [Goo85b]. Entropy [GY88, Joe88, Yat85]. Equal [Kem87b]. Equality [Per80]. Equation [SYW89]. Equations [Cla83, Fer82b, FHG81, Joh84, Kar81c, NC89, Ree85]. Equivalence [Bai84, Che87a, Dev83, VJ87]. Equivalent [Sil84]. Equivariant [EO87, Ebe83, Kar89b, TWZ80]. Error [AS80, AF84, BD84, Dev81b, Fab88, Fuj85, FS89, Hal84a, Kem87b, KW84, Pol86, Rit85, Ruk83, SC85, TP82, WT82, WT85, Yat88]. Errors [Ame80, Ame84, AF84, Bas87, BR87, Cha89, Ful80, Gle81, GCG87, GH87, HC86, Kel84, Mil89, Nel80, Rie81b, Sha80, Sta85, WF82b, Yaj88]. Errors-in-Variables [AF84, Ful80, GCG87, GH87, Mil89, WF82b]. Esseen [Ahm80, Che81b, Fri89]. Estimability [Mat84, SB80]. Estimable [Har81b]. Estimate [Boh82, BDD89, Dab88, Dab89, Dah88, Dev88, DS89, ES81a, Fuj85, GKP84, Kab80, KA83, MSS81, Red81, Shi81, Sil80, WJT86, Yao87]. Estimated [AF84, BN83, Ber82a, Neu87, Neu88, Por86, Ran82, RS84, Sch80]. Estimates [Abr82, Bai83, Bar86a, Ber84, Ber86c, BS80a, BD84, Cam82, CW88a, Cla86, CDM85, DK80, Dav87, DR84, Daw86, Dev81a, Dev83, DP84a, Dev85, DF86c, DF86a, DF86b, DL86, Dos85a, Dos85b, Dos86, Fie82, FT86, FD83, Fre84, Hal81b, HW84, HW85, Har86, Hjo86, KR81a, KR82, KP86, LW82, LT84, Le 86, Li84a, LH84, Li85, Lin86a, Lo84b, Mas83, Puk81, Rie80, Rom88, Ros75, Ros82, SS81, Sto84, Str81, Stu84, SM80, TT83, Vio80, Wah81, Wei85, Wes86b, Wri81a, Wri82, Wu80, Yoh87]. Estimating [AK88, And51, And80, And84, BBMT81, Bir87a, Bir87b, Car86, CS81, Cha81, Cha82, Cho87a, Cho87b, CLS85, CLS86, Dos89, DFP89, DuM83, Eva83, Fer82b, Hal82, Joh84, Kel89, Kem87b, LPQ81, Loh84b, Ruk85, SC82, Shi80, Smi87, Woo85b, Woo87, vdV88]. Estimation [AG82a, AK88, AF84, AF88, ASZ85, Ano83, AK85, BK83, Ber80a, Ber82b, BM87a, Ber80b, BBZ82, Ber83a, Bha80, BCY83, Bic81, Bic82, BR87, BF89, BRH85, BS80b, Bro80a, BF85b, Bro88, BC89, Car88, CR82, Cas80, Chi86, CY81, CM82, CGW83, CS84a, CS84b, CRM87, Dah89, Dav84, Dav85b, DG80, DD89, DW80, DP84b, Dev87, DS85, Dos85a, Dos85b, Ede88, Fab88, Fer82b, Fin85, Fis82, Fre86b, Fre88, FGP88, Fu82, FHG81, Gaf87, GSV86, GH82, Geo86, GM80, GHT83, GNS87, GY88, Gle81, GS81b, Gup86, GS85b, Hab89, Haf79, Haf80, Haf81, Hal81a, Hal83b, Hal87, HM88, Han75, Han80a, Han80b, Han81, HM85a, HT88, Har81a, Har81b, Hat85, Hoo82b, Hoo83, Hoo84c, Hoo88, HT82, HT93, HM87b, Hud83, Hud85, Hud86, Hwa85, Kal83b, Kar89b, Kar87, Kau87]. Estimation [Kaw80, Kla87, KD83, Kou86, Kre87, KI81, Kub87, LaM82, LS83a, LM87, LR82, Lo82, Lo85, Lo86b, LB89, Mar85b, Mar87, Mas80, McK86, MR87b, Mil82, Moo84, Mor83a, MS87b, Mul89, NS85, Nog88, Nov80, Par86, Pha80, PK89, PK90, Pot83, Pot84, Ras80, Rei80, Rie81a, Ris83, Rob83, Ruk86, Ruk88, SY81, SY82, SA86, ST82, Sag82, SA82, Sch85a, Sch86b, STT80, SS87, Sha88a, SW89, SW80, Shi84, Sil82, Sil85b, SCR87, SS80, Sri87, Sta85, Ste86b, Ste81, SW81b, Sve89, Sza80, Sza83, Tah89, Tak87, TW83, Tsa86a, TLC86, TP82, Tua81, Van87, Van83, Var82b, Var82a, Vil82, Wal80, WB84, Wel89, Wel90, WF82b, Woo85a, Woo89a, Wri81b, Wu81a, Yaj88, Yao84]. Estimator [Ame80, Ame84, AdF80, Bai84, Bed87, Bed88, Ber80c, Ber82c, Bir89, BMP85, BH89, CY87, Dav85a, DED89, Est83, Est86, ES81b, FK85, FK86, Fal85, Fox81, Gil81, Gil83a, God80, Gra83, HT85a, HM87a, HH83a, HH83b, Har88, Ker85, KB85, LV85, MF83, MR88, Mar84, Mau85, Nai80, Oeh85, PV80, Rao78, Rao80, RU84, Rot85, Sau89, SS81, SV80b, SV80a, Swe80a, Swe82, Tan83b, TWZ80, Tyl87, Wan86a, Wan87, Wel82, Wel85, vH87]. Estimators [Akr86, Akr89, AAO86, AA88, Ano83, Bai84, BK80, BK81, BNR83, BB83, Ber80b, Ber83a, Ber83b, BM87b, BF85a, Cas87, CL88, Che81b, Chi88, CH80, CT87, Cli88, CP81, Col83, CW85b, CW89, Cox88b, CRM87, Cro80, Dio81, DL88b, DW88, EO87, Edd80, Egu83, Fal84, FKZ89, FR81, FD82a, Ful80, Gaj86, GHT83, Gle86, Hal84a, Har84, HL84, HJS88, Hub84, Hud83, Hud85, Hud86, Hwa82a, Hwa82b, JJV85, Jeg87, Jeg88, Jur81, Kee85a, KK86, Kil81, Kla84, Kla87, Kla89, Klo82, Klo84, KG88b, LaR82, Lee88, Leu82, Lin86b, LT83, Lop89, MC87, Man84, MP87a, McK84, MGV85, MGSV89, Mie86, Mor86, Mor83a, Mul84, MS87c, Nel80, Pfa82, Pie82, Por84, Por85, Por91, RR81, RB86, Ruk83, Sch86a, Sco85, SS85, SW87, Sha89, SD87]. Estimators [Sil85a, Sin81b, Sol81b, Sri84, Ste89a, SS89b, Sto80, Tan83a, TS80, TL84, TK86, Tru89, TC85, TC90, VM84, WW83, Wan85, Wan86b, Wel86, Wel87a, Wel88a, Wie86, Yat85, dWR87, vE83]. Euclidean [BK80]. Evaluation [Bha82]. Events [Eva83]. Eventual [ESS81]. Everywhere [Dev81a]. Evidence [HM87b, Wal87]. Evoked [MTG84]. Exact [Chr89, KW84, Nie81, Sin81b, WT80]. Exactly [Hoe81]. Example [Ped81]. Examples [Bro80a, Kal83a]. Exchange [AS89]. Exchangeability [LN81]. Existence [MSS81, Puk81]. Expansion [BG78, BG80, BQ89, BGvZ86, CJV80, Doe83, Fuj85, Hal83a, KN84, Rob80]. Expansions [BS89, Fie82, FS89, Hel80, JR83, Nav89, Sch81a, Sch89b, Tak87, Wit83, WT82, WT85, Woo86, Woo89b]. Expectation [LS89]. Expectations [AG79, AG80, Gol84]. Expected [Che80c, GH87, HS80, Hoo82a, Hoo84b, Hoo88, KM80]. Experimental [Bro84c, Cha84, Cha85, PT83, Wyn84]. Experiments [Bed82, Col84, DG81, HS83, Kee84, Ker85, Kou84a, Leh88, MM87, MG83, SWW84, Swe80b, Woo89b]. Expert [GS85a]. Explicit [SM80, Sza80]. Exploratory [Hal89a]. Explosive [BMMT89, Jeg88, KL89]. Exponent [HT85a]. Exponential [Ama82, Ano83, BLR82, BL83, BLE86, BNB83a, BNB83b, Ber80b, Ber83a, Egu83, Fei81, Fri82, GSV86, GHT83, Hau88, Hu88, Hud83, Hwa82a, Hwa82b, Jew82, Kal81, Kou84a, Kou84b, LS89, Lin83c, Lin86b, MV87, Mor82, Mor83a, Mor83b, Por88, RR81, Ruk86, Ste86b, Sti85, WS80]. Exponential-Type [Ste86b]. Exponentiality [GHP86]. Extended [Edw87, Irl84]. Extended-Paulson [Edw87]. Extension [Dok87]. Extensions [Sim80]. Externally [Gen84, GMS86]. Extraction [Bel84, Bel91]. Extrapolation [Hoe81, O’R76, OD80]. Extreme [AGW86, BDD89, DR84, DD89, DED89, Gil83b]. Extreme-Value [DD89, DED89].

F [Tib88]. Factor [AA88, Kel81]. Factorial [CD86, Che80a, Col84, CC88, VD87, Vos88]. Factorials [Gho80, Gho91]. Factorization [GMY81, Hal84b]. Factors [BS86]. Failure [BMP85, Jac89, SS81, Wan86b]. Families [Ama82, Ano83, BLE86, BNB83b, Ber80b, Ber83a, Fei81, Fis82, FD82b, GHT83, HM85b, Hu88, Hud83, Hwa82a, Hwa82b, Kal81, Kou84a, Kou84b, LM85, MR87a, MV87, Mor82, Mor83a, Mor83b, MFPW89, Por88, Sch81b, Tak81]. Families-Curvatures [Ama82]. Family [Efr82a, Efr82c, Egu83, Hau88, LS89, Lin83c, Lin86b, RR81, Sti85]. Feedback [DG81]. Ferguson [HH83a, HH83b, WK81]. Fide [Gaj86]. Fiducial [Bar82, DS82a, DS82b, Fra82]. Field [Ste88, Ste89a]. Fields [DLS80]. Fieller [Kos87]. Filtering [AK85]. Finding [Boh82]. Finetti [FD82b, Reg87]. Finite [Ale89a, Ale89b, BL81, Bro81, Cam82, Chr86, FHW81, GBT82, GH87, Gor83, Gut82, Hub84, Joc86, Ken83, KM87, LT83, Lo86a, Lo88, MG81, MG83, MGV85, Ruk83, SK86]. Finitely [HS89a, Li84c]. First [Kou86, MF83, MC88, Sri87, Wan86a]. First-Order [Kou86, Sri87, Wan86a]. Fisher [Rob82a, Chi81, DF82, Gra83, Riv89, Rob76]. Fit [Bar89, Bas87, Chr89, Dro89, Fou80, GM83, GP85, HT86, Hau88, Khm88, Kou86, LM85, MM80, Moo82, Nei88, Nie83, QR85, Sch83a, Sch83b, Sve81]. Fitting [Bha82]. Fixed [Har81b, LS83a, MSS81, Rot85, Sil85a, Spr85, SG81, SS89b, ZM84, ZM88]. Fixed-Size [MSS81]. Fixed-Width [Spr85]. Flattening [Abr84]. Foliations [BNB83a]. Form [PS83]. Formal [BG78, BG80]. Forms [Bha82, GL88]. Formula [JS89, Pol86]. Formulation [Hat85]. Fourier [Fl80]. Fractional [Che80a, CC88, PR81]. Fractions [Stu80]. Frechet [Cla83]. Free [BY80, CM87, CC88, DW80, DP84b, GKP84, Swe86b, Tyl87]. Frequency [Chi88, Rub84]. Frequentist [LB89, WF82a]. Fully [TWZ80]. Function [Ale80, BBMT81, Bha86, BC89, CY87, Che88b, CK85, DW80, Dev81a, Eub81, Fre86b, Fre88, FGP88, Fuj85, GSW80, HH83a, HH83b, Har84, HL84, HM85a, Kel84, Kla87, Klo82, LS89, LR82, Lo82, Lo86b, Loy80, MR88, Mas81a, Mie86, Mil82, MV87, NJ85, Pha80, Rup81, Sch85a, Sch87b, SS81, Sil82, Sim83, SM88, Ste88, Ste89a, Stu84, SW81b, TW83, Tan83b, TLC86, Vio80, Wal87, Woo85b, Woo87, Yao84, Yu89, Zeh81]. Functional [AF88, Bar82, Cla83, DS82a, DS82b, Fra82, Sen88, Vil82, Wit83]. Functional-Model [Bar82, DS82a, DS82b, Fra82]. Functionals [Bri86, Das89, DL88a, Don88, FH86, HJS88, Kee83, Kee87a, Kun86, Li84b, MY86a, MY86b, ML86, Poo86, Rob86, TL84, Tsa86b, Weg86, Wes86a]. Functions [AK88, AdF80, BLE86, BLB89, BS83b, BNR83, BS85a, Ber86b, BS87, BS80a, CM89b, DR86, Epl82, Fab88, Fri89, Gaj86, Hal84a, Har81b, Hel80, Hwa85, JSV84, Kem88, Lev81, LT83, Mas83, McC83, Mor82, Mor83b, Neu87, Neu88, RH83, Rei81, Ruk88, WT80, Wri81b, vD85]. Further [Gaf87, Kee85b, PR81, WW83, Zam81]. Future [Fin83, HP88b, JB83, JBB83, Kak83].

Gait [OBWS89]. Galton [BN83, Swe86a]. Gambling [vD85]. Gamma [Ber80b, FKZ89, Gup86]. Gauss [BK81, Sti81]. Gaussian [Bed87, Bed88, Bro84a, Epp87, Epp88, FT86, God80, Hor88, LR82, Man84, McK84, SK86, Ste89a, Tan83a]. GCV [Wah85a]. Gemmell [Wat82]. General [AA88, Bai84, BK81, BBZ82, Bic84b, Bjø84, Car86, Col84, Das89, Dio81, Dok87, Fin85, Ham81, Hoo83, Hoo84c, JSV84, Kun89, Li83, Lin83b, Mar83a, MR87b, Mil82, MV87, PS83, PT83, Sch89a, Sen88, SW89, Spr80]. Generalised [BS83b, Loy80]. Generalization [Dia89, KP81]. Generalizations [SM88]. Generalized [ALP81, Ber80c, CP81, CKWY88, Csi89, Dro89, FK85, FK86, Gor89, Kre86, Li85, Li86, Li87, RU84, Ser84, Sto86, TK86, TC85, Tsa86a, TC90, TWZ80, Vos88, Wah85a, Pre84]. Generated [Epl82]. Generating [Col84]. Generator [Vos88]. Geoffrey [Bin85]. Geometric [Csi89, MLV84]. Geometries [BN86]. Geometry [Ama82, AK88, BNJ88, GM87a, GM89, Lin83b, Lin83c]. Gill [SP82]. Given [BK83]. Global [BMWZ87, Sto82]. GLSE [Toy86]. GML [Wah85a]. Good [Est86]. Goodness [Bar89, Dro89, Fou80, GP85, HT86, Khm88, Kou86, LM85, Nie83, QR85, Sch83a, Sch83b, Sve81]. Goodness-of-Fit [Dro89, Fou80, GP85, HT86, Khm88, Kou86, LM85, Nie83, QR85, Sve81]. Graph [FR83]. Graph-Theoretic [FR83]. Graphical [DF84, LW89a, LW89b]. Graphs [SK86]. Graybill [Nai80]. Greatest [Leu82]. Grenander [Bir89]. Gross [Rie81b, Rit85]. Grouped [Wri82]. Grouping [Wu89]. Groups [Gen84]. Growth [GMK^{+}84a, GMK^{+}84b]. Grubbs [But82]. Grubbs-Type [But82]. Guarantee [Gra83].

Hadamard [Che87a]. Hajek [Tie87, Tie88]. Harald [Blo87]. Harmonic [Bro85]. Having [BMP85, Goo85b]. Hazard [Dok87, LV85, MFP85, Sch85a, TW83, Tan83b]. Heat [SYW89]. Heavy [Wel85]. Hermite [Wal80]. Heteroscedastic [CR82]. Heteroscedasticity [CR81, Car82, Ham81, MS87b, SW87]. Heteroscedasticity-Robustness [SW87]. Hidden [Ros89]. Hierarchal [Wes86b]. High [Sau89, Yoh87]. High-Dimensional [Sau89]. Higher [KS82, Kee87a]. Higher-Order [Kee87a]. Hilbert [Man84]. Hill [HT85a]. Hinkley [Ped81]. Histogram [Kog87, LV85]. Histograms [Bir87b, Sco85]. Holland [Son82]. Holley [Ros89]. Homogeneity [CS87, CM89a, MTG84]. Honest [Li89]. Hotelling [HS80, JS89, Kar81b]. Hsu [Mil80]. Huber [Jon87]. Hypersphere [Riv84a]. Hypotheses [BCS80, BCS89, BM89b, HT89, Jen88, Lai88, Mar82b, Rom89, Wan81]. Hypothesis [BM89b, CKWY88, Dio81, Doe83, KN84, Lor83, Ruk82, Wu86c].

I.I.D. [Fox81, Liu88]. Identically [Ber82b]. Identifiability [FHW81, Ken83]. Identification [Che80d, LW82, Pos86]. Identifying [Atk84, Sch87b]. Identity [Cre82, Haf79, Haf81, SW80]. Ignorance [Tor81]. II [CG81, Por91, Atw80, CG80, Dos85b, Jeg88, KM87, Lin83c, MM80, Por85, Vil81]. Images [Dan82]. Implications [SCR87]. Implicitly [Rup81]. Impossible [Gut82]. Improper [HS89a]. Improved [Ano83, BW81, Ber83a, GHT83, Hud83, HC86, Mor83a]. Improving [Ber80b, Ber83b, FHH83, Gaj86, Hwa82a, TS80]. Imputation [SW88]. Inadmissibility [BBZ82, BC81, BH89, Epl81, Sri84, TWZ80, Yu89, Zam81]. Inadmissible [Ber80b, Ber83b, Fox81]. Incidence [BRH85, Gra88]. Including [KR82]. Incomplete [BRH85, Che83, EK83, MN83, TC85, TC90, WT80]. Incompletely [AK85]. Inconsistency [BMP85]. Inconsistent [DF86a, FD82a, FD83]. Increased [Lin85]. Increasing [BMP85, Mam89, Wan86b]. Independence [BM85, Dab86, Daw80, DE85a, DE85b, Dic85, Fie85, Goo85a, GC87, Goo85c, Hab81, JS85, Kar81c, Leo85, Loh89, Mal83, MP85, Pie85b, Pla85, Ros75, Ros82, SW84, Sun85]. Independent [Ber82b, Bro80a, DR82a, Fri89, Mar82a, Mar85a, Swe86b, Yao87]. Index [CDM85, DD89, DED89, DuM83, Li87]. Indexed [Wri81b]. Indices [Hal89a]. Indifference [BS84]. Induced [RP80, Sen81]. Inequalities [Eat82, HT89, KR81a, Por89, RSC87, Ton82, Ton89]. Inequality [Ale89a, Ale89b, DF82, Dev87, JS89, MRW86, Ros89, Ste86a]. Inference [AR86, AJ81, Amm84, BS83a, Bar82, BB84, CW87, DS82a, DS82b, DH81, Don88, Fei81, Fra82, HWB82, HS89a, Kra84, KR81b, LS83b, LN81, Lin85, Lo86a, MRW86, Reg87, RW80, RW82a, SM88, Smy88, Swe86a, Tyl81, Vil81, VC84, Yan83]. Inferior [Che80c]. Infinite [Ath87, Jan86, Kni89b, Kni89a, Sch83b]. Infinite-Dimensional [Sch83b]. Infinitely [AK88, Bro84a, Li84b]. Infinitesimal [Bic84c, Kun84b]. Infinity [Dro89, Por88]. Influence [Atk84, Bri86, FH86, Kel84, Kla87, Kun86, MY86a, MY86b, ML86, Poo86, Rei81, Rob86, Smi89, Tsa86b, Weg86, Wes86a]. Influential [Atk84]. Information [Ama82, BHHW83, Bha80, BRH85, Gra83, Hab84, Hab86, Jan86, Kni89a, LY88, Mam86, Pet80, Riv84a, Tor81, VM84, vdV89]. Inherent [HM87a]. Initial [AK85]. Inner [Vil81]. Innovation [Khm88]. Integers [Ris83]. Integral [NC89, Wri81b]. Integrals [DM83, SC82]. Integrated [AS80, Hal84a, Kaw80]. Integration [Sha88b]. Intensities [BRH85, RH83]. Intensity [BBMT81, Kar87, Lo82]. Interaction [DLS80, Hab81, MS88]. Interactions [DS83]. Interblock [Bha80, Col84]. Intergroup [FA86]. Intermediate [Kal83a]. Internal [Koy87]. Interpretation [Csi89, Was89]. Interpretations [BM85, DE85a, DE85b, Dic85, Fie85, Goo85a, Goo85c, Leo85, MP85, Pie85b, Pla85, Sun85]. Intersection [Dav89, OT81]. Interval [Akr88, But82, Hal86c, Hoo83, Hoo84c, KR81a, MC87]. Interval-Censored [Akr88]. Intervals [BO88, Ber88, Bic88, BGL88, CM87, DH81, DR88, Efr88, Est82, Hal86b, Hal88a, Hal88b, Hsu81, Hsu84, LS88c, Loh88, Rob88, Spr85, Vea88, Wit83]. Intrablock [Col84]. Intrinsic [HWB82]. Invariance [God80, Hoo82b, MT84, PR81, Sen81, Sen82]. Invariant [ASZ85, Bro88, Fox81, FGP88, Har81a, Hoo82a, Hoo84b, Hoo84a, Lai81, MP80, Mar81, Mar83a, PO80, SS84, Sma83, Szk88, Tak81, Tak82, Wes89, Yu89]. Invariantly [Ebe83]. Invariants [And82, MC86, Wij85, Wij93]. Invention [Sti81]. Inverse [Ale80, Haf79, Haf81, Jen88]. Inverting [Hal83a]. Investigation [Sko85]. Involved [But81]. Involving [Che80c, Wan81]. Irreversible [HW89]. Islands [BW82a]. Isotonic [DR82a, Kel89, Lee81, Lee83, Muk81, ST82, WW80]. Isotropic [Eat81]. Isotropy [Let81]. Iterated [Har84]. Iterative [Csi89].

J [Lin80, Pre84, Tib88, Wij84]. J. [Bic84b]. Jack [Ano84, Bro84c, Sac84, Wyn84]. Jackknife [Ber84, Ber86a, CR86, Che81b, ES81a, Efr86, Fel86, Fre86a, Gho86, Hal86a, Hin86, KR82, KR81b, Kun89, MO86, Ols86, RP86, Sha86, SW87, SW89, ST86, Sin86, Sri86, Tib86, Tib88, Web86, Wu86a, Wu86b, WO86]. Jackknifing [Sen88]. Johnson [Zab82]. Joint [Bai83, Gil83b]. Jointly [DFP89]. Judgments [GS85a]. Julius [RS85]. June [Ney81]. Justifiable [Rub84]. Justification [Bha83].

Kaplan [Dab88, Dab89, Mau85, MGSV89, RU84, Rot85, Wan87]. Karlin [KS84]. Kendall [ACF82]. Kernel [Abr82, BDD89, Cli88, CDM85, Dab89, Dev83, DP84a, Dev85, Dev88, DJ89, Edd80, Fal84, Fal85, GKP84, Hal84a, HM87a, HM88, Mar85b, MP87a, Mie86, Mul84, MS87c, RH83, Rom88, Sch85a, Sil80, Sil84, Sto84, TW83, Tan83b, TS80, Tru89, Won83]. Kernels [SY81]. Kiefer [Ano84, Bro84c, Che87a, Huf83, Sac84, Wyn84]. Known [SWW84]. Kolmogorov [MS92, FGP88, Hu85, MS83, Nie81, Wie86, Yat85]. Kramer [Hay84]. Kuh [Atk84]. Kulkarni [KJ86]. Kullback [Hal87, Jan86].

L [Wij84]. L. [Lin80]. Lack [Chr89, Nei88]. Lack-of-Fit [Chr89]. Lagrangian [Pot84, Pot83, PT83]. Large [AG82b, BC81, Che80d, Dav85a, DD89, FT86, Fu82, Gil83a, GVW88, Gle81, Gor83, Hal83b, Kal85, KK86, Kou84b, Lo87, Por84, Por85, Por91, SP82, SS81, SD87, Sim83, SV80b, SV80a, ST81, TC85, TC90, Tsi81, Wu86c]. Large-Sample [FT86]. Latent [Chi81, Gly80, HR86]. Latin [KW81]. Lattice [BS86, Muk81, Ros89]. Law [Abr82, Est83, Har84, ST81]. Lawley [WT80]. Laws [Nag82, Woo82]. Layouts [Gaf81]. Learning [Man81]. Least [AS80, BK80, BW81, BH89, CW88a, Chi88, CH80, CT87, Cox88a, Cro80, ES81b, Fre84, GCG87, Hal83b, Kab80, Kar80, Kaw80, Kin80, LW82, Lin86b, MF83, Nel80, Sol81b, Sti81, TT83, TK86, Van87, Vil82, WW83, Wei85, Wei87a, Wu80, Wu81a]. Least-Squares [CT87, Cro80, Fre84, Kab80, Lin86b, Van87, Wei85]. Lecture [Bic84b]. Leibler [Hal87, Jan86]. Length [Gru88, Ris83, Var82b]. Lengths [LPQ81, Pol87]. Levels [Vos80, Woo86]. Lévy [AJ81, CW89]. Life [GHP86, LPQ81]. Likelihood [Ame80, Ame84, AAO86, BN86, BNJ88, Ber80d, Boh82, BMP85, Bra84, Bro84a, BDD89, CL88, Cla83, DK80, Dok87, Efr82b, FK85, FK86, GH82, God80, GS81b, Gra83, Hat85, Hin79, Hin80b, HM87b, JJV85, Jen87, Kar87, Klo82, Kou84b, LT84, MSS81, McC83, MC86, MGSV89, Par86, Ped81, Per80, Pfa82, Por89, Por88, QR85, Red81, Ree85, RW81, ST82, Sag82, Sau89, STT80, Sil82, Sim80, Str81, Swe80a, Swe82, SM80, TC85, TC90, Vil82, WB84, Wan85, Was89, Won86, Yao87, vH87]. Likelihoods [Lin83b, Lin83c]. LIL [BS80a]. Limit [DR86, Est83, Gil81, Gil83a, Hec85a, HY87, HT82, Hos89, HT93, Ked84, Mas81b, Nag82, PV80, WJT86, Wel82, Wel85]. Limiting [CW88a, CC82, ES81b, GCG87, Kee83, Kee87a, Wac80, dWR87]. Line [Gil83a, LS89]. Linear [And51, And80, And84, AG79, AG80, BK80, BK81, BN83, Bre89, Bro84a, BF85a, BHT89a, BHT89b, BD84, Car82, CR82, Cha84, Cha85, Che88a, CGW89, CH80, Cox89, CRM87, Cuz85, DLS80, DD87, Dio81, Doe83, Epl81, Epl82, ES89, FK85, FK86, FPV87, FKZ89, Fin85, Fre84, Fuj85, Gaf87, GG89, GK89, Hab81, HIP85, Ham81, Har81b, Hel80, HT82, Hos89, HT93, HC86, JS84, Jeg88, Jen88, Kla87, Koe87, KA83, KA89, KD83, Kre86, KN84, LaM82, Li84b, LR82, Man84, Mas81a, Mas81b, NJ85, Neu87, Neu88, Pes82, Pie85a, PK89, PK90, PT81, Pot89, Pre84, Sau89, Sch89b, Sch80, SB80, SS87, SW87, Sha88a, Shi80, SS84, Slu84, Spr80, SWW84, Sto81, Swe80b, Tit89, Toy86, Vil82, WS80, Wel87a, Wel87b, Wel87c, Wel88a]. Linear [Wel88b, Woo89b, Wu86c]. Linear-By-Linear [Hab81]. Linearity [Sen80]. Linearization [KR81b]. Linearly [BK83]. Linera [Kun83]. Link [LD89]. Local [BN83, Hec85a, KL87, LY88, Neu87, Neu88, Rie81a, Sch86a, Tah89, Tru89, Tyl83a, Woo85a, Woo89a]. Locally [AS89, Kal81, Kar80, Kla87]. Locating [Che88b]. Location [BS80b, Col83, CW89, Dav87, DF86a, Dro89, FJ84, FD82b, Gat84, HS87, Jur81, Kla84, LM85, Leh88, Lo84a, Lop89, Mas83, Mor86, Ree85, SS85, Shi84, Tak81, Wie86]. Location-Scale [Dro89]. Location/Scale [LM85]. Log [DLS80, Jen87, Por89, Sau89]. Log-Likelihood [Jen87, Por89]. Log-Linear [DLS80, Sau89]. Logarithm [Har84]. Logistic [Pre81, SC85, TC65, TC80]. Logit [Ame80, Ame84, Dav84, Dav85a, Dav85b]. Lognormality [LH82, LH83]. Long [DT89, Yaj88]. Long-Memory [Yaj88]. Long-Range [DT89]. Look [Fis83]. Loss [Ama82, BS80b, DS85, FGP88, GSW80, GY88, Gle86, Hal87, Hwa85, Kel89, Kem87b, Kem88, LY88, Lee81, Lee88, LB89, NS85, Ruk88, Shi84, TP82]. Losses [Gup86]. Low [Est82]. Lower [Ber80a, Che88b, DP84b, Tah89, WF82a, Yat88, vE83]. LR [Kre86]. LR-Statistics [Kre86]. Lu [Mil80].

Macro [Van83]. Magnitudinal [GMT86]. Main [Gho80, Gho91]. Majorization [Joe88]. Making [Man81]. Manova [PO80, Gly80, LM87]. Many [Li84b, Li84c, Por86, AK88]. Marginal [Ton89]. Market [Fin83]. Markoff [BK81]. Markov [DLS80, SK86, Van83, VC84]. Martingale [Jac89, Nel80]. Matched [MV87, MLV84]. Matching [AGW86]. Matrices [Dav87, Eri87, LM87, Lin89b, Lin89a, Mar81, Per80, Sak83]. Matrix [AF84, BS85a, BS87, Chi81, DS85, Gle86, Haf79, Haf80, Haf81, Riv84a, Sza83]. Max [Lee83, MYZ89]. Maxima [DF82, Mit81]. Maximal [And82, Sma83, Wij85, Wij93]. Maximizing [CP81]. Maximum [Ame80, Ame84, AAO86, Ber80d, Boh82, BMP85, BDD89, CL88, Che88b, Cla83, Col86, DK80, Efr82b, FK85, FK86, GH82, God80, GS81b, Gra83, Hat85, HS89b, HM87b, JJV85, Kar87, Klo82, KR84, LT84, MSS81, MGSV89, Par86, Pfa82, Red81, ST82, Sag82, Sau89, STT80, Sil82, Str81, Swe80a, Swe82, SM80, TC85, TC90, Vil82, WB84, Wan85, Yao87, vH87]. Maximum-Likelihood [Hat85]. McCullagh [Pre84]. Mean [AS80, Ame80, Ame84, Ath87, Bed87, Ber80c, Ber82c, Bha80, Bic81, BL81, BF85a, Bro85, CS81, CY81, CM82, Cla85, DD87, Ede88, GSV86, GZ89, Gle86, GHP86, Hal81a, HC82, Joh84, Kem87b, Kni89b, Koe87, Kub87, Man84, Mar85a, MR87b, Nai80, NS85, RW80, Sri87, Ste81, SV80a, Swe86a, Sza83, Vos80, Wel87b, Wel87c, Wel88b, Wen89]. Means [BS83a, Bro80a, BF85b, Col86, GM80, GY88, Hud85, Hud86, Kel89, Lee88, LB89, MP80, Pol81, RH83, Rob76, Rob82a, Spr80, Sza80]. Measure [Bro85, DuM83, Fou80, Hoo82a, Hoo84b, Hoo88, Ros75, Ros82, SKS89, Wac80]. Measurement [LS80, Sta85, SC85]. Measurements [CW80b, CW83, Kun83, Kun84a]. Measures [ACF82, And82, DH81, FJ84, FR83, JSV84, KL87, Mac86, Mac87, Sch86a, SW81a, SB89, ST81, Tor81, Wak81]. Measuring [Koy87, RW82b]. Median [Dos85a, Dos85b, GPSB84, Kil81, MD87, TC65]. Medians [Tru89]. Meier [Dab88, Dab89, Mau85, MGSV89, RU84, Rot85, Wan87]. Memorial [BG82]. Memory [Yaj88]. Method [CG85, CP81, Cox88b, GH82, HLS89, Li82, Li85, Mas80, McK86, Pol87, Sch89a, Sil82, Sil84, TW83]. Methodologies [BQ89]. Methods [Abr84, Ber86a, Bri88, BL89, CR86, Dok87, DJ89, Efr86, Fel86, Fre86a, GK82, Gho86, Hal83c, Hal86a, Hin86, Jac89, KR81b, Loh84b, MO86, Ols86, Por88, RP86, Sha86, Sha88a, ST86, Sin86, Sri86, Tib86, Tib88, Web86, Wu86a, Wu86b, WO86]. Metric [Abr84, ST81]. Metrics [FA86]. Midranges [Gil83b]. Min [Lee83, MYZ89]. Min-Max [Lee83, MYZ89]. Minaxity [Ruk86]. Minimal [Mar82b]. Minimality [GMY81]. Minimax [Bed82, Ber82c, Bic81, BL81, Bir87a, Bjø81, BS80b, Cas80, CL83, Col83, CW85b, CW89, Gaj88, Geo86, Gle86, HC82, KW84, Kub87, Lo82, Mas83, MR87b, Mie81, SY82, Ste89b, Tah89, Wan86b, Wie86]. Minimaxity [Hoo89, Li82, Li83, Rie81a, Woo85a, Woo89a]. Minimization [Kee85a]. Minimizing [AS80, Obe82]. Minimum [Ame80, Ame84, BM89a, Ber80d, CC88, Dav85a, DL88a, DL88b, Egu83, Hab84, Hab86, Har81a, Har81b, KD83, Kou86, Ris83, Ste89a, Wan86a, Yat85].

Minimum-Variance [Har81b]. Minorant [Leu82]. Mises [CC82, FHW81, Riv89, Sch80]. Missing [DK80, Goo85b, Sza83]. Misspecified [Ste88]. Mixed [Bro84b, GC87, LT84, Rob82b, SM80, TL84, Wes86b, Wes89]. Mixed-Dirichlet [GC87]. Mixing [Ede88, FK84]. Mixture [BS89, Hat85, Lin83a, Lin83b, Lin83c, Lin86b, Pol86]. Mixtures [AG82a, CG80, CG81, FHW81, FS89, Haf79, Haf81, Jen87, Jew82, Ken83, LS88a, Lin89a, RW82a, WT80]. ML [Bai84]. MLE [GSW80, Toy86]. Modal [ST82]. Mode [BL89, Edd80, Oeh85, Rom88]. Model [AF84, AG82b, AH88, AR89, Bai83, Bai84, BK81, Bar82, Bha82, BR87, Bro84b, Che88a, CM87, CT87, CKWY88, Dav84, Dav85b, DS82a, DS82b, DD87, ES81b, Fin85, FL87, Fra82, Fuj85, Ful80, Gaj88, Gle81, GCG87, GMT86, HP88a, Ham81, Hau88, Hoo83, Hoo84c, HC86, Kar89b, Kar87, Koe87, KD83, Kun83, LT84, Lo84a, McK86, PdT88, Por86, Pot89, SY84, SP82, Sen81, Shi80, Sim86, SS84, Sto81, SM80, Tsi81, Wel87a, Wel87b, Wel87c, Wel88a, Wel88b, Wes86b, Woo82, Yaj88]. Model-Free [CM87]. Modeling [GS85a, KG88a, Ris86]. Modelling [Smi89]. Models [AK85, BL83, BNB83a, BNB88, BNJ88, BB84, BHHW83, Ber82b, BM89a, Bre89, BHT89a, BHT89b, BD84, Car82, CR82, Cha84, Cha85, CGW89, CJ89, CH80, CW85b, CKWY88, Cox89, DH80, DLS80, DS83, Dok87, Dro89, ESS81, ESS82, ES89, FK85, FK86, FKZ89, Fre81, Fre84, Fri82, GG89, GK89, GVW88, GH87, Goo85b, HP88a, Han75, Han80b, Han81, Har81a, Har81b, HF82, HR86, Jeg88, Kau87, Kil81, KA89, Kra84, Kre86, LW82, LR89, LW89a, LW89b, Li84b, Lin86b, Liu88, LC87, MS88, Nav89, Nel80, Nus85, PK89, PK90, PT81, PT82, Pos86, Pot83, Pot84, Pot89, PS83, Rei80, Rot85, Sau89, Sch86b, SB80, SS87, SW87, Sha88a, Spe85, Sto85, Sto86, Sve81, Sve89, TT83, Tit89, Tsa86a, Tsa84, Van83, Var85, Vil82, Wei85]. Models [Wei87a, Wes89, WF82b, Woo89b, vdV88, Pre84]. Moderate [Kal83b, Kal85]. Modes [Mul84]. Modified [MS83, MS92, VJ87]. Moment [CP81, DED89, Lin89b, Lin89a]. Moments [LS85b]. Monospline [Bru80]. Monotone [CT87, FPV87, HR86, Muk88, Wri81a, Wri82, vH87]. Monotonicity [BP84, RP80, Wij83]. Monro [Rup85]. Monte [HM87b, Joc86]. Mood [Kil81]. Morris [Wij84]. Most [EG80, Kal81, Szk88]. Motion [AR89]. Motivated [DR84]. Moving [Chi86, DR86, God80, Hal84b, MF83]. Muirhead [Wij84]. Multi [HT89, Kel81, Lai87, WZ83]. Multi-Armed [Kel81, Lai87]. Multi-Bayesian [WZ83]. Multi-Hypotheses [HT89]. Multinomial [Bro88, DR82b, DR87, Kal85, Lev81, RW81]. Multinormal [KR81a]. Multiparameter [Ano83, Ber83a, GHT83, Gup86, Hud83, Mor83a]. Multiple [BW82b, DM83, FHH83, Geo86, Hay84, Hsu84, Nis84, PT82, Ruk82, SW88, Sha80, Tyl83a, Vos80, Wac80]. Multiplicative [DS83, Kar87, Sve81, Sve89]. Multiplicity [Gly80]. Multiplier [Pot83, Pot84]. Multistage [Wit86]. Multivariate [ALP81, AF84, And51, And80, AAO86, ABJ83, BM86, BM89a, Ber80c, Ber82c, BZ80, BF85a, BF85b, Chi81, Cli88, CC82, DS86, Dav87, Eat82, EK83, Fie82, FJ84, FD89, FR83, Fuj85, FS89, Gle81, GMT86, Haf80, Hen88, Hoo83, Hoo84c, HC82, JM82, Kar81a, Kar89a, Kre86, KN84, LS88b, Lop89, Mar82b, Mar83a, NS85, OT81, Oma88, PO80, Per80, Rup85, Rus76, Rus82, SM82, SS85, Sin84, SS84, Ste81, SW81b, Sza80, Sza83, Tyl87, Wei87b, Wij84, Wij84]. Multiway [RS84]. Mutual [DFP89].

Naiman [JS89]. Naive [SS81]. Natural [BLE86, Mor82, Mor83b]. Nature [MP84]. Near [Chr89, Kel81]. Nearest [Abr84, Hen88, Li84a, Sch83a, Sch83b, Stu84]. Nearly [CW87, Lai88]. Necessary [Bro80b, BCSC83, GS81a, GZ89, Sri84, Sza80]. Negative [Hwa82a, Hwa82b, JDP83]. Negatively [RSC87]. Neighbor [Abr84, Hen88, Li84a, MC88, Sch83a, Sch83b, Stu84]. Neighbourhoods [Bic84c, CW89, Wie86]. Nelder [Pre84]. Nested [Cox88a]. Neurophysiology [Bri88]. Neutrality [JM80]. Newton [Rup85]. Neyman [GMY81, Jan86]. Nile [BS83b]. NN [BM87b]. no [DS89]. Noise [Che80d, HM87a, KA83]. Noisy [MTG84, NC89, Yao84]. Non [AR89, Ber82b, Ber89, Bro84a, Fis82, LR82, Liu88, SA82]. Non-Bayesian [AR89]. Non-Centrality [SA82]. Non-Gaussian [Bro84a, LR82]. Non-I.I.D. [Liu88]. Non-Identically [Ber82b]. Non-Null [Ber89]. Non-Parametric [Fis82]. Nonasymptotic [Bir87a, Bir89]. Noncentral [Mar82a, Mar85a, WT80, Wij83]. Noncentrality [Cho87a, Wij83]. Noncompact [RB86]. Nondegenerate [Nag82]. Nondeterministic [Kab80]. Nonergodic [BB84]. Nonexistence [GH87]. Nonexponential [Sch81b]. Nonfull [O’R76, OD80]. Nonidentifiable [Red81]. Nonlinear [AF88, BW81, CG86, FL87, HWB82, KG88a, Mal80, MLV84, Nei88, Obe82, RR81, RB86, SW84, SD87, WF82b, Wu81a]. Nonlinearity [HWB82]. Nonlocal [KL87]. Nonnegative [Das89, Har81a, Mat84, Puk81, TS80]. Nonnormal [Wes86b]. Nonnull [KS85, KN84]. Nonparametric [Amm84, Bar86b, BBMT81, BHHW83, BNR83, BF81a, BCY83, BF89, Bro88, BC89, But82, Car88, CY87, CL88, CB85, Cla85, CRM87, DH80, DW80, Dev81a, Dev81b, Dio81, Dos85a, Dos85b, DL81, Est83, Est86, FR81, FR85, Fre86b, Fre88, GMK^{+}84a, GMK^{+}84b, GH82, Har84, HM85a, HT88, HS87, Klo82, Klo84, Li84a, Li89, Lo84b, Mar83b, MGV85, Muk88, Mul89, Neu82, Puk84, Rei80, Ric84, Rom89, Sag82, SW81a, STT80, Sco85, Spe85, SS80, Spr85, Sto80, Sto82, Sto85, SV80b, Tsa86a, TLC86, Var82b, Var82a, VC84, Vos80, WB84, Wit83, Won83, Yan83, Yat88]. Nonparametrics [Wan85]. Nonstationarity [FD89, Pot89]. Nonstationary [BBMT81, Bas87, Bel84, Bel91, CW87, Hal84b, HF82, Kau87, Tsa84]. Nonsymmetric [Ste86a]. Nontransitive [FL87]. Norm [BK80, Fab88, Obe82, Ste89a]. Normal [And51, And80, AA88, Ber80c, Ber82c, Ber89, Bic81, Bro80a, CV81, CS81, DK80, Ede88, Efr82a, Efr82c, Est83, FS89, GM80, Gle86, GMT86, Haf80, Hat85, HC82, Kel89, Kem87b, Kub87, Let81, LB89, MC87, Mar85a, Nai80, NS85, O’R76, OD80, Per80, Por85, Por91, Ruk85, SM82, Ste81, Sto81, Swe80b, Sza80, Sza83]. Normality [AH88, BF84, Cso86, Dav85a, Ebe83, FK85, FK86, Fal85, HT85a, JSV84, LR89, LY88, Lee81, LSF86, Mas81a, Ran82, SW84, Stu84, Swe80a, Swe82, Wes86b]. Normalized [TP82]. Norms [Vio80]. Note [AH88, BS80a, Dev85, ESS82, GPSB84, GH80, Hin80b, IK80, Kak83, Kee87b, Kra84, Lo82, Mal87, Mil80, Nel80, PV80, Pha80, Puk84, Red81, Sch85a, Stu89, Tan83b, Wal80, Wan87, WK81, Zeh81, vD85]. Notes [Bic84b]. Nuisance [AK88, Fer82b, Wan81]. Null [Ber89, BM89b, CKWY88, Doe83]. Null-Hypothesis [Doe83]. Number [Con82, Dav85a, Dro89, Hal86c, HS89b, Hen88, Por88, Ree85]. Numbers [Che80b, ST81].

Observational [Ros89]. Observations [BF89, Che83, Jen87, KW81, Kun89, Mam86, Muk81, VC84, Wri82]. Observed [BN86, Gra83, GS85b, KA83]. Occurring [Bha82]. Off [Sam80]. Off-Diagonal [Sam80]. Offspring [Swe86a]. One [But83, CM87, Don88, Gho80, Gho91, HS87, JJV85, Kel81, KW84, Ler86, Mar82b, Mar85a, Pol86, RR81, Rie81b, Sti85, Sza83, Wel87a, Wel88a]. One- [JJV85]. One-and [Rie81b]. One-Dimensional [KW84]. One-Parameter [RR81, Sti85]. One-Population [Sza83]. One-Sample [HS87]. One-Sided [Don88, Mar82b, Mar85a]. One-Step [But83, Wel87a, Wel88a]. One-Step-Ahead [CM87]. Only [BL83]. Operations [Daw80]. Operators [GMS86, KM80]. Optics [Che87a]. Optimal [AZ85, BNR83, BW82a, But83, Cha84, Cha85, sC79, CW80a, Che83, Che87b, Con81, DW88, Eub81, ESS81, ESS82, Fin83, Gaf81, GH80, HP88a, Ham81, HM85a, HM85b, HT89, HP86, Hoe81, Hoo89, KS77, KS83, Kog87, KJ86, Kun83, LM85, Lai88, LS85b, LS88b, MN83, Mar83b, MP87a, Mil82, Mou86, Nai84b, Pes82, Pie85a, Pol85, Pol87, PT83, RSC87, SS89a, Shi81, Spe85, Spr80, Sto80, Sto82, Sto84, Stu80, Stu89, Stu87, Wah81, YB80, ZM84, vD85]. Optimal-Partitioning [HT89]. Optimality [Atw80, Bas80, Che80a, Che81a, Con82, Gaf81, Gaf87, Gre80, Hal83b, Irl84, JN83, Kab80, KS85, Kou84a, Kun84a, KM87, Lai81, Li86, Li87, Nai84a, Rom88, Wel82, Zeh81]. Optimization [Che87b, DW88]. Optimum [Edd80, GK80, GK82, KW81, KW84, MS88, Mul84, SY81, Zer80]. Order [Ame80, Ame84, Bha86, BQ89, Bir87a, BDD89, CV81, DuM83, DR82b, DR87, Egu83, GSW80, GS81a, GL88, HM88, Han80a, Hel80, HD89, KS82, Kee87a, Kee84, Kou86, Kre86, Lee88, MF83, Mar83c, Mar86, Mar91, Mas81a, MC88, MRW86, Pot83, Pot84, RW80, RW82a, Rus76, Rus82, Sen80, Shi80, Sko85, Sol81a, Spr80, Sri87, Tan83a, TC65, TC80, Toy86, Tsa84, Wan86a, Wen89]. Ordered [Goo85b, Wri81b]. Ordering [Cap88, Hor88, RP80, RW81, Sch87a, Sim80]. Orderings [BW82b, Loh84a, Mac86, Mac87]. Orders [CC88]. Orthogonal [AdF80, Che80b, Hal83c, HS89b, HS83, Mas80, Sch80, Wah81]. Orthogonal-Series [AdF80]. Orthogonality [CD86]. Other [Ber86a, CR86, Efr86, Fel86, Fre86a, Gho86, Hal86a, HP88a, Hin86, MO86, Ols86, PO80, RP86, Sha86, ST86, Sin86, Sri86, Sto85, TC65, Tib86, Tib88, Web86, Wu86a, Wu86b, WO86, MGSV89]. Outlier [But81, But82]. Outliers [DS86, SM82, Sin84].

P. [Pre84]. Paired [Joe88]. Pairwise [GMY81]. Pao [Mil80]. Pao-Lu [Mil80]. Paper [Dav84, Dav85b, WK81]. Parabolic [Fer82a]. Paradox [Gut82, SKS89]. Parameter [ASZ85, Bai83, BW81, Bic81, Cho87a, CG86, CRM87, Dah89, ES81b, Fer82b, FT86, Gaf87, Gat84, GBT82, HWB82, HF82, HT82, HT93, Kab80, Kaw80, LS83a, LM87, MF83, MSS81, MY84, MR87b, RR81, Rao78, Rao80, RB86, RW82b, Ruk83, SA86, SA82, SD87, Sti85, Wah85a, Wij83, ZM84, ZM88, dWR87, vdV88]. Parameters [AK88, Bai84, BL83, Ber80b, BS80b, Cam82, Chi86, Chi88, Dav87, Dio81, DR82b, DR87, FKZ89, FHG81, Gly80, God80, HW84, HW85, HP86, Kee85a, Li84b, Pie82, Por84, Por85, Por86, Por88, Por91, Ran82, Ruk85, Sch80, Shi80, Shi84, Sil85a, SS84, TT83, TP82, Tua81, Van83, Wan81]. Parametric [BS83b, BHHW83, BM89a, Bic84a, Che88a, CKWY88, DH80, Fis82, Kra84]. Parametric-Nonparametric [BHHW83]. Pareto [JM82]. Parsimonious [Sto81]. Part [CG81, Dos85a, CG80, Dos85b, Lin83c]. Partial [CKWY88, Dok87, Gaf87, LS87, Pet80, Sak83, Won86]. Partially [GS85b, Lin85, Wri81b]. Particular [Ped81]. Partitioned [SB80]. Partitioning [HT89]. Partitions [KRS87]. Partly [Che88a]. Past [HP88b, JB83, JBB83]. Pathologies [DL88b]. Pattern [Gre89]. Patterned [Sza80, Sza83]. Patterns [Sag82]. Paulson [Edw87]. PBIB [Con82]. Peak [Mul89]. Pearson [Ney81, Jan86]. Penalized [Klo82, STT80, Sil82]. Penalized-Likelihood [STT80]. Penalty [FPV87, Sim83]. Peninsulas [BW82a]. Performance [CY81, Dev88]. Periodic [PdT88]. Permutation [JR83, Lam85, Rob80, Ros89, Tar81]. Permutations [FA86]. Pfanzagl [Bic84b]. Phadia [WK81]. Phase [LR82]. Phenomenon [Bro80a]. Piecewise [Epl82, Fri82]. Pitman [JJV85, Neu82, Rot81, SKS89]. Pivotal [AS85]. Plane [Dab88]. Plans [Gho80, Gho91, KS84]. Plot [Gil86]. PLS [HD89]. Plus [Gho80, Kem87b, Gho91]. Point [AR86, But82, Car88, Dos89, DM83, Fu82, GM80, KI81, MFP85, Pra88, SY81, SS89b, Tah89, Tua81, Woo85a, Woo89a, Yao87, Yoh87]. Points [Dav85a, Hoe81, Li84c]. Pointwise [GKP84]. Poisson [DR87, Hud86, AR86, BBMT81, Boh82, BF85a, BF85b, DR82b, DM83, GY88, Hud85, Hwa82a, Hwa82b, Joh84, LT84, Lo82, Nov80, SW80, Sve81, TP82]. Policies [FR85]. Polynomial [Bru80, Fab88, Hal89a, LS85b, LS88b, Stu80, Stu89]. Polynomial-Based [Hal89a]. Polynomials [Nie81, Nie83]. Pooling [GMS86]. Population [Che80d, Cla85, Gly80, GM87b, LT83, Lo86a, Lo88, MG83, MGV85, Sza83, VM84]. Populations [BL81, Che80c, Gor83, Kub87, Nai80, Per80, RW81]. Positive [LS89, RP80, Sch87a]. Positively [Ton89]. Positivity [KR81a]. Posterior [Bha81, Gat84, SB89]. Postulate [Zab82]. Power [BLE86, Ber86b, JR83, Ler86, Pol86, Wes89]. Powerful [Bar89, EG80, Kal81, Szk88]. Prediction [Bas87, BD84, But82, CM87, FR83, Hoo84a, Hoo88, Ste88, Wei87a, Ylv87]. Predictions [LS85a]. Predictive [Hin79, Hin80b]. Predictor [Tak81]. Predictors [Tak82, Wei87a, YB80]. Preliminary [SS87]. Presence [AK88, Che80d, IK80, Leu84, MGSV89, TLC86, Var82b]. Preservation [LY88]. Preserving [BK81]. Price [DS89]. Principal [Flu86, Tyl83a, Tyl83b]. Principle [Abr84, MT84, OT81, Sto86]. Principles [Sen81, Sen82]. Prior [DS82c, MD84, Ris83, SYW89, VM84]. Priority [Ale89a, CLS86, Tie88]. Priors [BB86, HS89a, SB89]. Probabilistic [KR81a]. Probabilities [DS82c, Gol84, Man81, Ruk83, Sie86, WT82, WT85]. Probability [AdF80, BW82b, Bra84, Cha81, Cha82, Daw85a, Daw85b, Dev81b, Eat82, Fis83, FL87, Fou80, Har88, Hec85b, Hu85, Irl84, Klo82, KR84, Lai81, Lin80, MRW86, Pol86, Sch85b, Sch89a, STT80, Sil82, Smi87, Ton82, Van83, Wak81, WF82a, Wal80]. Problem [BS83b, BBZ82, Ber83b, sC79, CW80a, Che87a, Che87b, DF82, Ede88, GS81a, Huf83, Kee85b, Kel84, Kem88, Lai87, LS80, Mal87, Mar83a, Mar85b, Pet80, Pra88, Ras80, Rob76, Rob82a, Ruk86, Slu82, Sri84, Sza80, Sza83, Wah85a, WB84, Yat89, Zam81]. Problems [BP84, Bro81, BS84, BF85b, Bro88, BM89b, But81, DW88, Gut82, KS77, KS83, Lia88, MG81, Mie81, Neu82, PO80, Ruk82, Sta85, TK86, Wit86, Yat88, Yu89, vD85]. Procedure [CY81, CM82, Hay84, KJ86, Mar83c, Mar86, Mar91, Pos86, Ram83, Rup85]. Procedures [AR89, Bjø81, CS84a, FHH83, GH80, GM84, HP88a, HS80, Liu88, LC87, Mie81, Rit85, Ruk82, Sha80, Vos80, ZM84, ZM88]. Process [AR86, BBMT81, Bed87, Bed88, Bha81, Boh82, Bri88, Cam82, Chr86, DT89, Epp88, Fer82a, Gil86, God80, Han80a, HY87, KA83, KL89, MY84, RH83, Rao78, Rao80, SW80, Shi80, SYW89, Sri87, Sve89, Swe86a, Toy86, Tua81]. Processes [AJ81, Akr86, Akr89, AG82b, BB84, Bas87, BMMT89, BFY87, Bro84a, CW87, CW88a, Dah89, Dos89, DM83, Hal84b, HD89, HT82, Hos89, HT93, Kaw80, KW84, Kni89a, Kre87, Kun84b, Li82, LR82, Mas80, MFP85, McK84, Mil89, NP87, SC82, SP82, Sen81, Spr80, Ste86b, Stu86, Tan83a, Var82a, VC84, Wan86a, Wel89, Wel90, HT93]. Product [Gil81, Gil83a, HY87, PV80, WJT86, Wel82, Wel85]. Product-Limit [Gil83a, HY87]. Profile [BNJ88]. Programming [Sha89]. Progress [Hin80a]. Projection [Boo85a, BS85b, CW85a, Cox85, DF84, Dia85, DJRS85, DJ89, FJ84, Fri85, GK85, Hal89a, Hal89b, HT85b, Hub85a, Hub85b, Jon85, Jon87, Mil85, She85, Swi85, Tuk85b, Tuk85a, Wah85b]. Projection-Based [DJ89]. Projections [Eat81]. Proof [Hay84]. Proper [Wij85, Wij93]. Properties [BBC81b, But83, CW89, Cro80, Cuz85, Dos85b, FT86, Ful80, Hal84a, Jan86, Joc86, KR81a, Kee85a, Ker85, KA83, KR81b, KJ86, LaR82, LT84, LT83, MC87, Mie86, Nis84, Rob76, Rob82a, Rot81, Sha89, Smi84, Sol81a, TC65, TT83, Tru89, Wal80, WJT86, Wei85, Wu83]. Property [BL81, Col83, God80, Hoo84a, Kab80, Kar81b, RW80, SY82, Sim80, Tsa86a, Wen89]. Prophet [RSC87]. Proportional [Dok87]. Proportionality [Eri87].

Proportions [RS84]. Prove [Mas81b]. Pseudo [Che81a, GS81b, Par86]. Pseudo-Youden [Che81a]. Pseudofactor [Vos88]. Publications [Ano84]. Purely [Kab80]. Pursuit [Boo85a, BS85b, CW85a, Cox85, DF84, Dia85, DJRS85, FJ84, Fri85, GK85, Hal89a, Hal89b, HT85b, Hub85a, Hub85b, Jon85, Jon87, Mil85, She85, Swi85, Tuk85b, Tuk85a, Wah85b].

Quadratic [Bha82, Gle86, Gre80, Gup86, GL88, Kre86, Lee81, Lee88, LT83, Mat84, Mor82, Mor83b, NS85, Ros75, Ros82, Shi84, Ste89a]. Qualitative [BFY87, Hal83c, LW89a, LW89b, Rie82]. Quantal [Amm84, Bha81]. Quantile [CLS85, CLS86, DD89, Eub81, Fal85, LaR82, Sen81]. Quantiles [Fal84, Rei80, Rot85, Ruk86, Wit83]. Quantitative [Hal83c, LW89a, LW89b]. Quasi [McC83, Pfa82]. Quasi-Likelihood [McC83]. Quasirandom [Sha88b]. Quick [Pfa82]. Quotient [And82].

R [Atk84, RS85]. Random [Ahm80, Akr86, Akr89, Bri88, Cur81, DG80, Est83, ES81b, FKZ89, Fri89, GSV86, GS85b, Haf79, Haf81, Hor88, JDP83, Koy87, KS84, LR89, Leu84, MR88, MP84, NT88, Oeh85, RSC87, Sch85a, Sch86a, SC82, SW81a, Ste88, Ste89a, ST81, Ton82, Ton89, WJT86, Wri81b, Yao87]. Randomization [PR81, Rom89, Smy88, WSS86]. Randomized [Hoo89, Li83, Tar81, Wu81b]. Randomly [CHL83a, CHL83b, FR81, HT86, KSV81, Kum87, MP87a, TW83, Tan83b]. Randomness [HIP85]. Range [DT89]. Ranges [SB89]. Rank [Akr88, BN83, BNR83, Cap88, Cuz85, Cuz88, Cuz90, Dab86, Doe83, Epl81, Epl82, HIP85, HP88a, Leu84, Mas81b, Neu87, Neu88, O’R76, OD80, Pra88, Rie81b, Rie82, Rus76, Rus82, Sch89b, Sen80, SW84, Slu84, Sni81, Tar81, Wu86c]. Rank-Based [HP88a]. Ranked [Dia89]. Ranking [FA86]. Ranks [Mas81b]. Rao [BMWZ87, MFPW89, vE83]. Raphson [Rup85]. Ratcliff [Csi89]. Rate [BMP85, Che88b, JJV85, LV85, Nov80, Sch86a, SS81, SW80, Wan86b]. Rates [AS89, BRH85, Che88a, Che81b, HW84, HJS88, HY87, Lia88, Mar83b, Nog88, NP87, NC89, PV80, Ruk83, Sch81a, Spe85, Sto80, Sto82, TS80, Yan83, Yat85]. Ratio [AAO86, Bro84a, Hec85b, Irl84, Jen87, Kou84b, Lai81, MC86, Ped81, Per80, Por89, QR85, RW81, Ruk85, Wes89, Wij83]. Rational [BLB89]. Rationality [Sen88]. Ratios [Sam80, Sim80, Wac80, Wij85, Wij93]. Real [vdV88]. Reality [Ber89]. Recall [Chr86]. Recentered [Cas87]. Recovery [Bha80]. Rectangular [BS86, KR84, Ton82]. Recurrence [Joh84]. Recursion [Sol81a]. Recursive [Cam82, Han80b, Sen82]. Reduction [DS89, Kel89, Sto86]. Reference [ACF82]. Refinement [Kun83]. Regarding [Jos75, Jos85]. Region [Ber80c]. Regions [BW81, BS85a, BS87, BW82a, HWB82, Hoo84a, KR84, Li89, MV87, Oma88, Was89]. Regression [AG82b, And51, And80, AH88, Atk84, Bai83, Bai84, Bar86b, Ber86a, Bha83, BCY83, BM87b, Bic84c, BF89, BZ80, Bru80, BD84, BC89, But83, CR86, Cas80, sC79, CW80a, Cha86, Cha89, CH80, CT87, CG86, Cox88a, CRM87, Cuz88, Cuz90, Dav84, Dav85b, DW80, Dev81a, DR82a, Efr86, Fab88, FPV87, Fel86, Fre81, Fre86a, Gaf87, GMK^{+}84a, GMK^{+}84b, Gho86, Gle81, GCG87, GKP84, Hal84a, Hal86a, Hal89b, HWB82, Har84, HL84, HM85a, HT88, Hin86, Hoe81, HM87b, JS84, Jon87, Kar80, Kar81c, Kau87, Kel89, Kil81, Kin80, Klo84, KG88a, KSV81, KD83, LW82, LS85b, Lee81, Lee83, Li84a, Li84b, Li84c, Li86, Li89, LD89, LS88b, Loy80, Mam89, MYZ89, MP87b, McK86, Mil82, MO86, MLV84, Muk81, Muk88, Mul84, MS87b, MS87c, NT88, Nai86, Nav89]. Regression [NJ85, Nei88, Nel80, Nis84, Nus85, O’R76, OD80, Obe82, Ols86, Pes82, Pie85a, PdT88, Por84, Por85, Por86, Por91, Pot89, Pre81, PS83, RP86, RR83, Ric84, RB86, Riv89, SY84, ST82, Sam87, SP82, Sen81, Sha86, SD87, Sil85a, ST86, Sin86, SS84, Sol81b, Spe85, SS80, Sri86, SC85, Sto81, Sto82, Sto85, Stu80, Stu89, Stu84, Tib86, Tib88, Toy86, Tsi81, WS80, Web86, Wei85, Wei87a, Wel86, Won83, Wri81a, Wri81b, Wri82, Wu86a, Wu86b, WO86, Yaj88, Yat88, Yat89, Yoh87, Atk84, Atk84]. Regressions [Mil89]. Regressors [CT87]. Regret [CM82, Tah89]. Regular [Atw80, BB84, HT85a, HW84, HW85, Wit83]. Regularity [Tie87, Tie88]. Regularization [Cox88b, Li82]. Regularized [NC89]. Rejoinder [BHT89b, Daw85b, DE85a, DF86b, Hal88a, Hub85b, MY86b, Spe87a, Wel87b, Wu86b]. Related [BBMT81, Bas87, BF85b, GH87, KR81a, MGV85, WS80, WW80]. Relating [Por89]. Relation [Lop89, Tak81, vE83]. Relationship [AF88, BDD89]. Relationships [And84]. Relative [Fal84, MV87, PS83, Rot81, vE83]. Relevant [Rub84]. Reliabilities [DFP89]. Reliability [DL81]. Remarks [Bas80, Bas83]. Renewal [Fre86b, Fre88, Var82a]. Rényi [Nie81]. Repeated [CW80b, CW83, Hu88, KR81b, Kun83, Kun84a, LS80, MTG84, Sen81, Tak87, WT82, WT85]. Replacement [FR85, Gil81, KS84]. Replicate [VD87]. Replicates [Chr89]. Replication [KR81b, NJ85, PR81]. Reply [Ano83, DS82b]. Representation [Fei81, Lev81, Tie87, Tie88]. Representations [CL88, Gly80, Wal87, Wel86]. Reproducibility [BLE86]. Reproductive [BNB83b, BNB88]. Require [Hoe81]. Required [Hal86c]. Resampling [Ber86a, CR86, Efr86, Fel86, Fre86a, Gho86, Hal86a, Hin86, Loh84b, MO86, Ols86, RP86, Sha86, Sha88a, ST86, Sin86, Sri86, Tib86, Tib88, Web86, Wu86a, Wu86b, WO86]. Research [Bro84c]. Residual [DGG81, GHP86, KL89, TWZ80]. Residuals [Loy80, Mil89, Por86, Sen82, Wel86, Atk84]. Respect [GSW80]. Response [Amm84, Bha81]. Responses [Eic88, Sim86]. Restricted [Bic81, DR82b, DR87, Lee88, Loh84a, MR87b, MRW86, RW80, RW82a, Smy88, WSS86, ZM84, ZM88]. Restriction [DR82b, DR87]. Restrictions [And51, And80, Bir87a]. Result [KP81, Kou84b, SD87, Wu86c]. Results [BBZ82, BM89b, CS84b, DW80, Gle81, JS84, KG88a, Li83, PR81, Ruk86, SW88, SP88, WW83, Zam81]. Return [vD85]. Review [Atk84, Bic84b, Bin85, Eat82, Pre84, Wij84]. Revisited [Fis83]. Ridge [BZ80, Cas80, Li86]. Right [Jac89, KSV81, MP87a]. Right-Censored [KSV81, MP87a]. Risk [Ber80a, Bir87a, Bir87b, Gra88, Li85, MY84, Mar83b, Mar83c, Mar86, Mar91, MV87, PS83, Toy86, Zer80]. Risks [Kem88, LPQ81, TK86]. Robb [Wij84]. Robbins [Rup85]. Robust [BS83b, Ber80a, Ber82b, Ber80c, BB86, Bic84c, BF89, CR81, CR82, HT88, Kar80, Kin80, Lam85, Li84c, Mam89, MYZ89, Pes82, Qua85, Rie80, Rie81a, Rit85, SY84, Sam87, Sil85a, Wan81, Was89, Wel86, Yoh87, ZM88]. Robustification [FK84]. Robustness [Bic84a, BFY87, DL88a, GC87, HH83a, HH83b, Kar81a, Kar81b, KS85, Kar89a, KG88b, Kun84b, Ram83, Rie81b, Rie82, SW87, Wu81b]. Role [LN81]. Root [Abr82, FD89]. Roots [Chi81, Gly80, Ree85, Tyl83a]. Rotationally [LC87]. Rule [But82, Sha82, Sto84]. Rules [HP86, HW89, Lia88, Sim86, WSS86]. Run [CC88, Pol87].

S [Bin85, Ney81]. Sample [AG82b, BNR83, BL81, Bro81, Che80c, CL83, CM87, CS84a, CS84b, Dah88, DR86, DG80, Est83, Fie82, FT86, Fu82, GPSB84, Gil83a, GVW88, Gle81, Gol81, Gra88, Gut82, Hal83b, Hen88, HS80, HS87, Hub84, Joc86, Kee83, Kee87a, Kni89b, Lam85, Leu84, Lo87, Mar84, Neu82, RS87, Rie81b, SP82, SS81, Slu84, SV80b, SV80a, TC85, TC90, Tsi81, VJ87]. Samples [AG79, AG80, Est82, KR81b, MSS81, MP87a, Mit81, Rob80, Sta85, SV80a]. Sampling [Ale89a, Ale89b, Ass88, AR89, Ber82a, BF84, Chr86, GVW88, Gor83, Gra83, Hal81a, HLS89, Iac83, KS84, Lo86a, MG83, MGV85, MP84, SG81, Ste89b]. Sanford [Atk84]. Savage [Lin80, Ros89]. Savings [HS80, HS87]. Savits [KP81]. Scalar [Sak83]. Scale [Ber80b, Dro89, FKZ89, Gaj88, HT88, LM85, Sil85a, Swe86b, Tak81, Wel86]. Scale-Free [Swe86b]. Scaling [Csi89]. Scatter [Tyl87]. Scheffe [Boo85b, Swe86c]. Scheme [Wu89]. Schur [Bjø81, Ton82]. Schur-Concave [Ton82]. Score [BNR83, Lin83a, Mas81a, MFP85, Neu87, Neu88]. Score-Functions [BNR83]. Score-Statistic [MFP85]. Scores [BN83]. Search [AZ85]. Seasonal [HF82]. Second [BQ89, Egu83, GSW80, GS81a, Kee84, Mar83c, Mar86, Mar91, MC88, Sko85, Sol81a, Spr80, Tan83a, Toy86, Wen89]. Second-Order [Sko85, Toy86, Wen89]. Seemingly [Kar81c]. Seismology [Bri88]. Selected [Eat82, GM87b]. Selecting [Ber82c, Kra84]. Selection [BP84, Bha86, Bjø81, Bjø84, But81, But83, Edw87, Eub81, GH80, GM84, HM87a, HM85a, KJ86, MP87a, Mie81, Nag82, Nis84, Pot89, Shi80, Sto81, Sto84, Tsa84, Var85, Woo82]. Self [Dah89, KG88a, TC85, Tsa86a, TC90]. Self-Consistency [TC85, TC90]. Self-Consistent [Tsa86a]. Self-Modeling [KG88a]. Self-Similar [Dah89]. Semi [Hwa82b, VC84]. Semi-Markov [VC84]. Semimartingale [McK86]. Semiparametric [CKWY88, Sch86b, vdV88]. Sense [SKS89]. Sensitive [MS83, MS92]. Sensitivity [GC87]. Sequence [AGW86, Car86, Yao87]. Sequences [Atw80, DT89, SS84, SW81b]. Sequential [AZ85, BBC81a, BBC81b, BCS80, BC81, BCSC83, BCS89, Che80c, CY81, CM82, Chr86, Cla85, Dan82, DG81, Edw87, EG80, Fer82b, FR85, GSV86, GM80, GNS87, Gra83, GM84, Hal81a, Hec85a, Hec85b, Irl84, KS77, KS83, Kar88, Kee84, Kou84a, Lai81, Lai88, Mar83c, Mar84, Mar86, Mar91, Mas81b, NS85, Nov80, Qua85, Ras80, SW80, SG80, SYW89, Slu84, Smi84, Spr85, Sri87, Ste86b, Tah89, Tak87, Woo85a, Woo86, Woo89a, Zer80]. Sequentially [MY84, Woo89b]. Serial [HIP85, Kar80]. Series [AdF80, Bel84, Bel91, Bri86, Chi88, Dah88, Epp87, ESS81, ESS82, FD89, FT86, FH86, Hal81b, Hal83c, HP88a, HF82, JB83, JBB83, Kab80, Kau87, KS82, Ked84, Kee85a, Kun86, MY86a, MY86b, Mas80, ML86, Poo86, PT81, PT82, Pos86, Rob82b, Rob86, Slu82, Sol81a, Tsa86b, Wah81, Wal80, Weg86, Wei87a, Wes86a]. Set [Cas87, Hoo82b, Koy87, Li87, Moo84, RW82b, Sak83, Wri81b]. Sets [BS83b, BM86, GH87, Hoo82a, Hoo84b, HC82, HC86]. Setting [Lin83a]. Several [Per80, Rob80, Sco85, TP82]. Shape [GSW80, Ler86]. Shapes [Sch81b]. Shapiro [LSF86, VJ87]. Shapiro-Wilk [LSF86, VJ87]. Sharp [BCSC83, Sri84]. Sheffer [Nie81, Nie83]. Shift [HS87]. Shifted [Sco85]. Shortcoming [Kal81]. Shorth [Gru88]. Shrinkage [Geo86, GNS87, SS85, SS87]. Sided [Don88, Mar82b, Mar85a]. Sieve [Bed87, Bed88, CG85]. Sieves [GH82, Kar87, McK86]. Sign [FK84, IK80, MY84]. Signal [Bel84, Bel91]. Signals [MTG84]. Signed [Sen80, Wu86c]. Significance [Hu88, Sen81, Tak87, Vos80, WT82, WT85]. Signs [HP86]. Simar [Boh82]. Similar [Dah89]. Simple [BM89b, Doe83, Kil81, Mas81b, MP84, Pie85a, RS87, WB84]. Simulated [Ber86b]. Simulations [CM87, Hal86c]. Simultaneous [Ale80, Ale82, Ber80b, GY88, Gup86, HT88, Hoo83, Hoo84c, Hoo88, Hsu81, Hsu84, Hud85, Hud86, Mie81, Nai84a, Nai84b, O’R76, OD80, Shi84, TP82]. Simultaneously [Hwa85]. Sine [Woo82]. Single [BLR82, VD87, Vos88]. Single-Generator [Vos88]. Singular [Atw80]. Sinusoidal [Ked84]. Situation [Sch85a, Sch86a]. Size [Che80c, Gol81, HS80, HS87, Loh89, MSS81, Mar84, SG81, ZM84, ZM88]. Sizes [RP80]. Skewness [Mac86, Mac87]. Slippage [But81, DS86]. Slope [Bha83, BCY83, Leu82]. Slope-of-Greatest-Convex-Minorant [Leu82]. Small [BF81a, Bic84a, CM87, Con82, Dah88, Fie82, Ruk83]. Smallest [Hoo82a, Hoo84b]. Smirnov [MS92, FGP88, Hu85, MS83, Nie81]. Smooth [Hel80, Mas81a, Mul84, SYW89]. Smoothers [Bre89, BHT89a, BHT89b, CGW89, Cox89, ES89, GG89, GK89, KA89, Mal80, Tit89]. Smoothing [AK85, Cox83, HDR89, LH84, Li86, Mit81, Nus85, RH83, RR83, Sil84, Sim83, Spe85, Wah81, Wah85a]. Smoothness [KA83]. Sobolev [JS83, JS85]. Solution [Huf83, Kos87, Mar85b, Rob76, Rob82a, WB84]. Solutions [Cla83, DW88, NC89, Sza80]. Some [BBMT81, BF81b, BMWZ87, Bri88, CL88, Che80a, Chi81, Con81, ESS81, FHH83, Fis82, FS89, Ful80, Gaf81, KS85, LW89a, Let81, Li83, Mal80, MGV85, Mie86, Nag82, Rob83, Rot81, SY84, Sen81, SS85, SS84, SY89, VJ87, WT82, WT85, Wu81b, DT89, DL88b, LW89b, Liu88, Rom89, Stu89, Sve89]. Source [Chi86]. Sources [Atk84]. Space [AK85, Bic81, Gut82, Li84c, Man84, MR87b, RB86, SM88, Wij84, ZM84, ZM88]. Spaces [Bro81, CP81, ST81]. Spacing [Eub81]. Spacings [Swe86b]. Sparse [Sim83]. Spatial [MD87, Sag82]. Special [ACF82]. Species [Cha81, Cha82]. Specifications [HF82]. Specified [AK85, Fin85]. Spectra [Kee87a, Sma83]. Spectral [CG85, Dia89, Hal84b, Kee83, Mas80, Sam87, Shi81, Tua81]. Spectrum [Kab80]. Sphere [LC87]. Spheres [Bin85]. Spherical [Cha86, Cha89, Kin80, Riv89]. Spherically [BW82b, BS80b, HC86]. Sphericity [Let81]. Spline [CKWY88, Li86, Nus85, Sil84, Spe85, Wah85a]. Splines [AS80, Cox83, RR83, WW80]. Spread [JSV84, Kla84, Kla89]. Spring [JN83]. Springer [Bic84b]. SPRT [BCSC83]. Square [Abr82, AS80, Ame80, Ber80d, BM85, DE85a, DE85b, Dic85, Dro89, Fie85, Goo85a, Goo85c, HT86, Hal84a, KW81, Leo85, MP85, MM80, Pie85b, Pla85, QR85, RS87, Sun85]. Squared [Ame80, Ame84, BB83, Dav85a, GM83, Kem87b, Mar82a, Moo82, RS84, SA82, TP82]. Squares [AS80, BK80, BW81, BH89, CW88a, Chi88, CH80, CT87, Cox88a, Cro80, ES81b, Fre84, GCG87, Hal83b, Kab80, Kar80, Kaw80, Kin80, LM85, LW82, Lin86b, MF83, Nel80, Sol81b, Sti81, TT83, TK86, Van87, Vil82, WW83, Wei85, Wei87a, Wu80, Wu81a]. Stability [BDD89, McK84]. Stabilizing [Son82]. Stable [DuM83]. Stage [CW88b, Fre84, Lor83, NS85, Ram83]. Stagewise [Sha80]. Standard [Hwa82a]. Standardized [Jen87]. Starshaped [DR82b, DR87]. State [AK85]. Stationary [Car86, Epp87, Epp88, FT86, Fre84, HT82, Hos89, HT93, Jeg88, Ked84, Kre87, Kun89, Slu82, Sri87, Tua81, Yaj88]. Statistic [BLR82, BR82, BM85, CV81, Car86, CC82, Dar83, DE85a, DE85b, Dic85, Fie85, Gil86, Goo85a, Goo85c, Kar89b, Kou84b, Leo85, Mas81b, MFP85, MP85, Pie85b, Pla85, Sch80, Sun85]. Statistical [And84, ABJ83, Bic84b, BS80a, Bri88, Bro81, Chi86, Daw80, DW88, DR82b, DR87, JSV84, Lo86a, Mal83, Mam86, Mor83b, Reg87, RW80, Sha89, Sie86, Smi89, Vil81, Wal87, Wan81, Wij84]. Statistician [Rub84]. Statistics [AS85, Ahm80, ALP81, Akr86, Akr89, AG79, AG80, BK83, BL83, BNR83, BGvZ86, BDD89, CJV80, DT89, Doe83, DM83, Ebe83, GL88, Hel80, Hin80a, Hu85, Jen87, KR82, Kre86, LM85, Lin80, MT84, Mas81a, Mas81b, MC86, MP87b, Neu87, Neu88, Pie82, Por89, Ran82, RV80, Rus76, Rus82, Sch83a, Sch89b, Sen80, Ser84, Sha88b, Spr85, Ste86a, Tar81, TC65, TC80, VJ87, Wij84, Wu86c, dWR87, Bic84b, Bin85]. Stein [Dev87, DS85, Gut82, LH84, Li85, Ram83, SKS89, Ste86a]. Step [But83, CM87, JJV85, Wel87a, Wel88a, Yao84]. Steps [Wij85, Wij93]. Stepwise [MM87]. Stimuli [MTG84]. Stochastic [BM87a, BM89a, BFY87, Bro84a, BH89, CT87, DW88, FHG81, Hor88, Hwa85, Jeg88, Kel89, KA83, LW82, Li82, Muk81, Por89, Pot89, RP80, Ris86, RW81, Rup81, SC82, Sha89, Sim80, Wei85, Wei87a, Wei87b]. Stock [Fin83]. Stopping [BBC81a, BW82a, Chr86, Fin83, Kee87b, Mou86, RSC87]. Strategic [LS85a]. Strategies [KS77, KS83, SG81, Ste89b, vD85]. Stratified [BF84, KR81b, VM84]. Strict [Dav89]. Strong [CY87, CH80, CT87, Dev83, FR81, HJS88, HD89, Jeg87, Jeg88, MR88, Nel80, Pol81, Sil80, Sol81b, ST81, Wan85]. Strongly [FT86]. Structure [CJ89, Har81b, HS83, Sve89, Toy86]. Structures [Wak81]. Student [Rob82a, BS84, Rob76]. Studentized [CV81]. Studies [MLV84, MV87, Ros89, SP88]. Study [AG82b, Lo87, SP82, TC85, TC90, Tsi81]. Sub [BL83]. Sub-Vector [BL83]. Subfamily [BLR82]. Subfields [GMY81]. Subgroups [Col84]. Subjective [Man81]. Subjects [DG81]. Subseries [Car86]. Subset [Bjø81, Bjø84, But81, GH80]. Subsets [CG86]. Subspaces [Cox88a]. Substituting [Pie82, dWR87]. Successive [Gor83]. Sufficiency [BNJ88, Hoo82b, SM88]. Sufficient [BR82, BCSC83, Ebe83, GMY81, GS81a, Sri84, Sza80]. Sufficiently [Atw80]. Sufficientness [Zab82]. Suggested [Ker85]. Sum [LM85]. Sums [DF82, Wri81b]. Supercritical [Swe86a]. Supremum [Fab88]. Sure [Sen80, Tar81]. Survey [RS84, RS87, Ste89b]. Surveys [CL83]. Survival [Bai83, Bai84, CY87, FR81, Fri82, GP85, MV87, Slu84, SV80b, SV80a, Tsa86a, TLC86, Yan83]. Symbols [Che80b]. Symmetric [BW82b, BS80b, BL89, CLS85, CLS86, CG80, CG81, DS86, DM83, FD82b, HC86, Lo85, LC87, MT84, Riv84a, Riv84b, RV80, Sch87b, Sin84]. Symmetrically [MR87a]. Symmetry [JS83, Kin80, Sni81, Wu86c]. System [DFP89]. Systematic [AG79, AG80, Iac83]. Systems [Hoe81, LW82].

T. [Dav84, Dav85b]. Table [BM85, DE85a, DE85b, Dic85, Fie85, Goo85a, Goo85c, Leo85, Loh89, MP85, Pie85b, Pla85, Sun85]. Tables [AG82a, CG80, CG81, DLS80, Dav89, GM87a, GC87, GM89, Goo85b, Hab81, Mey82, RS84, Sim83]. Tail [Ale80, Cap88, CDM85, DR84, DuM83, Hu85, Hwa82b, Jur81, Loh84a, MS83, MS92]. Tail-Behavior [Jur81]. Tail-Orderings [Loh84a]. Tails [DS82c, Smi87]. Takacs [Nie83]. Tau [ACF82]. Technique [MM87]. Techniques [BBMT81, But81, Mar87]. Tendency [ST81]. Tends [Dro89, Por88]. Terms [Kre86]. Test [BLR82, BN83, BB83, BM89a, Bha83, BS84, Che80c, Cla85, FK84, Fou80, Gil86, GC87, HH85, Hec85b, HM85b, Hen88, HS80, IK80, Kar81b, Loh89, MS83, MS92, NT88, RP80, Ros75, Ros82, SEB83, SS87, Sha80, Stu87, Sve81, Tar81, vH87]. Testing [BM87a, Ber89, BM85, BM89b, CHL83a, CHL83b, CKWY88, Cso86, DE85a, DE85b, Dic85, Epp87, Epp88, Fie85, FD89, Gil81, Goo85a, Goo85c, GHP86, HP88a, HF82, HT89, Kar88, KN84, Kum87, Leo85, LSF86, MY84, MP85, Mie81, MTG84, NJ85, Nei88, Neu82, Pie85b, Pla85, PT81, Qua85, Ruk82, Sch83a, SG80, Sun85, Swe86b, Sza83, Tak87]. Tests [Akr88, BLR82, Bar89, Bas80, Bas83, Bas87, Bed82, BS85a, BS87, BBC81a, BBC81b, BCS80, BC81, BCSC83, BCS89, Cap88, CR81, Chr89, CS87, CM89a, Cuz85, Dab86, Dan82, Dav89, Dro89, DR82b, DR87, EK83, EG80, Epl81, Epl82, Fer82a, GM83, GP85, Gra88, Gre80, Hab81, HT86, HIP85, Ham81, Hec85a, HS87, Hu88, Irl84, Jan86, Joc86, JR83, JS83, JS85, Kal81, Kar80, Kar81a, Kar81c, KS85, Kar89a, Khm88, Kin80, Kou86, LM85, Lai81, Lai88, Lam85, Ler86, Leu84, Lor83, MP80, Mar81, Mar82a, Mar82b, Mar83a, Mar85a, MS88, MFP85, MM80, Moo82, OT81, PO80, Per80, Pol86, PT82, Pot83, Pot84, Pra88, QR85, RS84, RS87, Rie80, Rie81b, Rie82, RW81, Rob80, Rom89, Ros89, Sch83b, Sch80, Sen81, SW84, SYW89, Slu84, Sni81]. Tests [Sti85, Szk88, Tak87, Tyl83b, WT80, Wan81, Wes89, WT82, WT85, Zer80]. Their [CG80, CG81, FS89, Kin80, Bro88, Por89, vE83]. Theorem [Ahm80, Boo85b, Bro81, Che87a, Cho87a, Das89, Fei81, FD82b, Gen84, Hec85a, HT82, HT93, Swe86c, Tak87, Tie87, Tie88, Wel85, Zam81]. Theorems [BF85b, Hos89, Mas81b, Sen80]. Theoretic [FR83, Gat84, GM84]. Theoretical [BO88, Ber88, Bic88, BGL88, DR88, Efr88, Hal88a, Hal88b, LS88c, Loh88, Rob88, Vea88]. Theory [ACF82, AGW86, BF81b, Bjø81, Bjø84, BY80, Che87b, CM87, Dah88, Das89, DR84, DR86, Efr82a, Efr82b, Efr82c, Flu86, Fu82, GVW88, GS81b, Gre89, Hal81a, Iac83, JB83, Kal83a, Kal83b, Kau87, LPQ81, Lin83b, Mal80, Mor83b, NT88, PS83, PT83, SP88, SW89, SCR87, SV80b, SV80a, WF82a, WZ83, Won86, Wu81a, Bic84b, Wij84]. there [DS89]. Thickness [DuM83]. Third [Kre86]. Thoughts [Rob83]. Three [Dav89, Lor83, WT80]. Three-Dimensional [Dav89]. Three-Stage [Lor83]. Ties [IK80]. Time [Bel84, Bel91, Bri86, Chi88, Dah88, Epp87, ESS81, ESS82, FD89, FT86, FH86, GSV86, Gil86, HP88a, HF82, Jac89, JB83, JBB83, Kab80, Kau87, KS82, Ked84, Kee85a, Kee87b, Kun86, MY86a, MY86b, Mas80, ML86, Nov80, Poo86, PT82, Pos86, Rob82b, Rob86, Slu82, Sol81a, SV80a, Tsa86b, Weg86, Wei87a, Wes86a, Yao84]. Times [BBC81a, Mou86, PT81]. Tools [DGG81]. Topics [Eat82]. Topological [LS88a]. Total [Gil86, KR81a, Tor81, VM84]. Touchstone [MS87a]. Tradition [Hin80a]. Trajectories [Cho87b]. Transfer [Bha86, LR82]. Transform [Che87a]. Transformation [BNJ88, CJ89, Dok87, Efr82a, Efr82c]. Transformations [BK81, BW81, Son82]. Transforming [Mey82]. Transforms [Bra84]. Translation [Ebe83]. Treatment [KW84, Lai87, LS80, Lee88, WSS86]. Treatments [HM85b, MN83, Stu87]. Trend [BY80, CC88, GHP86, Puk84, RW82b]. Trend-Free [BY80, CC88]. Trial [SG80]. Trials [Sim86, Smi84, WSS86]. Trigonometric [Hal81b, LS85b]. Trimmed [But82, Col86, DD87, Koe87, Wel87b, Wel87c, Wel88b]. Triple [Hal81a]. Truncated [Bha83, BCY83, CL88, Gaj88, Woo85b, Woo87]. Truncation [WJT86]. Tubes [JS89]. Tukey [Hay84]. Turning [Gol84]. Two [Atk84, BK83, BNR83, BGvZ86, BM85, Che80c, CW88b, DE85a, DE85b, Dic85, Dos89, DR82a, Eic88, Fab73, Fab81, Fie85, Fre84, Gaf81, GM80, Goo85a, GM87a, GM89, Goo85c, Hab81, Hen88, JJV85, Jos75, Jos85, Kak83, Kar81c, Kee85b, Klo82, Kub87, Lam85, Leo85, Leu84, Lia88, MS88, MP85, Nai80, NS85, Pie85b, Pla85, Ram83, Rie81b, Rob76, Rob82a, Ros75, Ros82, Sha82, Sil85b, Slu84, Sun85, Wij83, Wij84]. Two-Action [Lia88]. Two-Armed [Eic88, Jos75, Jos85, Kak83, Kee85b]. Two-by-Two [GM87a, GM89]. Two-Dimension [Ros82]. Two-Dimensional [Ros75]. Two-Sample [Hen88, Lam85, Leu84, Rie81b, Slu84]. Two-Stage [CW88b, Fre84, NS85, Ram83]. Two-Step [JJV85]. Two-Way [BM85, DE85a, DE85b, Dic85, Fie85, Gaf81, Goo85a, Goo85c, Hab81, Leo85, MS88, MP85, Pie85b, Pla85, Sun85]. Type [AJ81, Bha83, But82, Cla83, Cox83, DF82, Fal84, Har88, Hen88, KR82, Kil81, MM80, Nie81, Pol86, PS83, Sma83, Ste86b, Yat88, Yat89, Pol86]. Types [Pie82].

UMPU [BLR82]. Unbalanced [MS88, Wes86b]. Unbiased [BK80, BK81, DS89, Fis82, Har81b, KA83, Li85, Mat84, Puk81, Tak81]. Unbiasedness [CS87, PO80, Per80]. Uncertainty [FL87]. Undominated [Lo84a]. Unidimensionality [HR86]. Unified [BP84]. Uniform [Dab89, FR81, GL88, HL84, HJS88, Hu85, Kun84a, Nog88, SS81, Sil80, Swe80a, Swe82, Wan87, vH87]. Uniformly [Bar89, EG80]. Unimodal [BL89, Lo86b, SB89]. Unimodality [HH85]. Union [Dav89, OT81]. Union-Intersection [OT81]. Uniqueness [Cla83, ESS81, MSS81, MD87]. Unit [FD89, HP88b, Riv84b]. Univariate [Hor88, Pos86]. Universal [BH89, Hwa85, Ris83]. Unknown [Gle86]. Unordered [Goo85b]. Unrelated [Kar81c]. Unstable [CW88a, Rao78, Rao80]. Up-and-Down [Ker85]. Updating [GS85a]. Upon [Hwa82a, Mil89]. Upper [Hwa82b, TK86, WF82a]. Use [Bro85, Car86, Ker85, Mas81b]. Used [CHL83a, CHL83b, Kum87]. Useful [BW82b, Cre82]. Using [AS80, And82, BS83a, But82, CP81, DH81, DFP89, Hec85b, Hoo88, Kum87, LS85b, Lin85, MM87, McK86, Muk81, Stu80, SW81b, Van83, vH87].

Validated [CGW83, Li84a]. Validation [BD84, Hal83b, Li85, Li86, Li87, Mar87, Won83]. Validity [BG78, BG80]. Value [AGW86, DR84, DD89, DED89, Gaj88, KR81a, KM80, RW80]. Valued [MSS81]. Values [Car86, CC82, LH82, LH83, RSC87]. Vapnik [SY89]. Variability [Bro85]. Variable [But81, But83, Che80b, Dev85, HR86, MS87c, Sil84, Tan83b]. Variables [AF84, BR87, Cha89, Cur81, DR82a, Fri89, Ful80, Gle81, GCG87, GH87, Hor88, JDP83, KR81a, Kel84, Koy87, LW89a, LW89b, Mil89, Nis84, RSC87, SW81a, ST81, Ton82, Ton89, WF82b, Wri81b, Yao87, YB80]. Variance [AAO86, And87, Ath87, Bai87, BLE86, BLB89, Bri87, Bro84b, Car86, CP81, CW85b, Dia87, ES81a, Gor89, Gra87, HH87, Har81a, Har81b, Hoo83, Hoo84c, KR82, Kee87b, Kem87a, Kni89b, Kni89a, MC87, Mar83a, Mat84, Mor82, Mor83b, Nai80, Nel87, Puk81, Rot85, SEB83, SW87, Sha88a, SW89, Spe87a, Spe87b, SM80, Tju87, Tob87, TWZ80, Tuk87b, Tuk87a, Wes86b, Wes89, Wie86]. Variances [CM89a, Col86]. Variates [Chi81, TC65, TC80]. Variation [Abr82, HT85a, HW84, HW85]. Varying [BM87b]. Vector [BL83, FKZ89, Gle86, MSS81, Wij84]. Vector-Valued [MSS81]. Vectors [Riv84b, Tyl83b]. Version [Rup85]. Versus [GHP86]. Very [Woo86, Woo89b]. Via [BK83, Loh84a, TC85, FA86, Kar87, Tsa86a, TC90, Wu89]. Violation [LD89]. Volume [JS89].

W [Bic84b]. W. [Zab82]. Wald [SEB83]. Walk [ES81b, NT88]. Walks [GS85b]. Watson [Bin85, Dar83, Swe86a]. Way [BM85, DE85a, DE85b, Dic85, Fie85, Gaf81, Goo85a, Goo85c, Hab81, Leo85, MS88, MP85, Pie85b, Pla85, Sun85]. Weak [BM87b, Dev83, KL89, KI81, Rom88, Sil80, Woo86, Woo89b]. Wefelmeyer [Bic84b]. Weighing [Che80a, GK80, GK82, JN83, SS89a]. Weight [Hel80, Mas83]. Weighted [Chi88, Gup86, LaR82, Sch83a]. Weisberg [Atk84]. Weiss [Huf83]. Well [Mar84]. Welsch [Atk84]. Whether [CHL83a, CHL83b, Kum87]. Which [Hoe81, Jen88, BLR82, Gaj86, LW89a, LW89b]. Whole [Gil83a]. Whose [BL83]. Width [Nai84a, Spr85]. Wiener [DM83, Fer82a, MY84, SYW89]. Wilk [LSF86, VJ87]. Wilks [KN84, WT80]. William [Wat82]. Williams [KM87]. Window [SS80, Sto84]. Wishart [Haf81, Haf79, KM80, Sam80]. Withdrawals [GSV86]. Without [Chr86, Goo85b, MS88, NJ85]. Wolfowitz [Che87a]. Work [Lin80]. Worthwhile [Bic84a]. Writings [Ano84]. Wu [Tib88].

Youden [Che81a].

Zones [BS84].

References

Anderson:1988:AND

[AA88] T. W. Anderson and Yasuo Amemiya. The asymptotic normal distribution of estimators in factor analysis under general conditions. Annals of Statistics, 16(2):759–771, June 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

           http://projecteuclid.org/euclid.aos/1176350834.

Anderson:1986:MLE

[AAO86] Blair M. Anderson, T. W. Anderson, and Ingram Olkin. Maximum likelihood estimators and likelihood ratio criteria in multivariate components of variance. Annals of Statistics, 14(2):405– 417, June 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

           8966 (electronic). URL http://projecteuclid.org/euclid.
           aos/1176349929.

Andersson:1983:DEM

[ABJ83] Steen A. Andersson, Hans K. Brons, and Soren Tolver Jensen. Distribution of eigenvalues in multivariate statistical analysis. Annals of Statistics, 11(2):392–415, June 1983. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346149.

Abramson:1982:BVK

[Abr82] Ian S. Abramson. On bandwidth variation in kernel estimates — a square root law. Annals of Statistics, 10(4):1217–1223, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345986.

Abramson:1984:ADF

[Abr84] Ian S. Abramson. Adaptive density flattening — a metric distortion principle for combating bias in nearest neighbor methods. Annals of Statistics, 12(3):880–886, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346708.

Alvo:1982:ATM

[ACF82] Mayer Alvo, Paul Cabilio, and Paul D. Feigin. Asymptotic theory for measures of concordance with special reference to average Kendall Tau. Annals of Statistics, 10(4):1269–1276, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345992.

Anderson:1980:AOS

[AdF80] G. Leigh Anderson and Rui J. P. de Figueiredo. An adaptive orthogonal-series estimator for probability density functions. Annals of Statistics, 8(2):347–376, March 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344958.

Amemiya:1984:EME

[AF84] Yasuo Amemiya and Wayne A. Fuller. Estimation for the multivariate errors-in-variables model with estimated error covariance matrix. Annals of Statistics, 12(2):497–509, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346502.

Amemiya:1988:ENF

[AF88] Yasuo Amemiya and Wayne A. Fuller. Estimation for the nonlinear functional relationship. Annals of Statistics, 16(1):147–160, March 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350696.

Arnold:1979:BEL

[AG79] Barry C. Arnold and Richard A. Groeneveld. Bounds on expectations of linear systematic statistics based on dependent samples. Annals of Statistics, 7(1):220–223, January 1979. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344567. See

[AG80].

Arnold:1980:CBE

[AG80] Barry C. Arnold and Richard A. Groeneveld. Correction to “Bounds on Expectations of Linear Systematic Statistics Based on Dependent Samples”. Annals of Statistics, 8(6):1401, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.

aos/1176345215. See correction [AG79].

Albert:1982:MDD

[AG82a] James H. Albert and Arjun K. Gupta. Mixtures of Dirichlet distributions and estimation in contingency tables. Annals of Statistics, 10(4):1261–1268, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345991.

Andersen:1982:CRM

[AG82b] P. K. Andersen and R. D. Gill. Cox’s regression model for counting processes: A large sample study. Annals of Statistics, 10(4): 1100–1120, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345976.

Arratia:1986:EVT

[AGW86] Richard Arratia, Louis Gordon, and Michael Waterman. An extreme value theory for sequence matching. Annals of Statistics, 14 (3):971–993, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350045.

Arjas:1988:NAN

[AH88] E. Arjas and P. Haara. A note on the asymptotic normality in the Cox regression model. Annals of Statistics, 16(3):1133–1140, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350950.

Ahmad:1980:BET

[Ahm80] Ibrahim A. Ahmad. On the Berry–Esseen theorem for random U-statistics. Annals of Statistics, 8(6):1395–1398, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345212.

Akritas:1981:AIL

[AJ81] Michael G. Akritas and Richard A. Johnson. Asymptotic inference in Lévy processes of the discontinuous type. Annals of Statistics, 9(3):604–614, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345464.

Ansley:1985:EFS

[AK85] Craig F. Ansley and Robert Kohn. Estimation, filtering, and smoothing in state space models with incompletely specified initial conditions. Annals of Statistics, 13(4):1286–1316, December 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349739.

Amari:1988:EPI

[AK88] Shun-Ichi Amari and Masayuki Kumon. Estimation in the presence of infinitely many nuisance parameters–geometry of estimating functions. Annals of Statistics, 16(3):1044–1068, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350947.

Akritas:1986:EPA

[Akr86] Michael G. Akritas. Empirical processes associated with V - statistics and a class of estimators under random censoring. Annals of Statistics, 14(2):619–637, June 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:/

/projecteuclid.org/euclid.aos/1176349942. See correction

[Akr89].

Akritas:1988:RTI

[Akr88] Michael G. Akritas. Rank tests with interval-censored data. Annals of Statistics, 16(4):1490–1502, December 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176351050.

Akritas:1989:CEP

[Akr89] Michael G. Akritas. Correction: “Empirical Processes Associated with V -Statistics and A Class of Estimators Under Random Censoring”. Annals of Statistics, 17(3):1417, September 1989. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347281. See

[Akr86].

Alexander:1980:SCB

[Ale80] Charles H. Alexander. Simultaneous confidence bounds for the tail of an inverse distribution function. Annals of Statistics, 8(6): 1391–1394, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345211. See correction [Ale82].

Alexander:1982:CSC

[Ale82] Charles H. Alexander. Corrections: “Simultaneous Confidence Bounds”. Annals of Statistics, 10(1):321, March 1982. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345719. See

[Ale80].

Alexander:1989:APC

[Ale89a] Kenneth S. Alexander. Acknowledgment of priority: A counterexample to a correlation inequality in finite sampling. Annals of Statistics, 17(4):1917, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347403.

Alexander:1989:CCI

[Ale89b] Kenneth S. Alexander. A counterexample to a correlation inequality in finite sampling. Annals of Statistics, 17(1):436–439, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347027.

Ahmed:1981:GAA

[ALP81] Abdul-Hadi N. Ahmed, Ramon Leon, and Frank Proschan. Generalized association, with applications in multivariate statistics. Annals of Statistics, 9(1):168–176, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345343.

Amari:1982:DGC

[Ama82] Shun-Ichi Amari. Differential geometry of curved exponential families-curvatures and information loss. Annals of Statistics, 10(2):357–385, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345779.

Amemiya:1980:OMS

[Ame80] Takeshi Amemiya. The n^{−2}-order mean squared errors of the maximum likelihood and the minimum logit chi-square estimator. Annals of Statistics, 8(3):488–505, May 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345004. See cor-

rection [Ame84] and comment [Dav85b].

Amemiya:1984:COM

[Ame84] Takeshi Amemiya. Corrections: “The N^{−2}-Order Mean Squared Errors of the Maximum Likelihood and the Minimum Logit Chi-Squared Estimator”. Annals of Statistics, 12(2):783, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346526. See

[Ame80].

Ammann:1984:BNI

[Amm84] Larry P. Ammann. Bayesian nonparametric inference for quantal response data. Annals of Statistics, 12(2):636–645, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346511.

Anderson:1951:ELR

[And51] T. W. Anderson. Estimating linear restrictions on regression coefficients for multivariate normal distributions. Annals of Mathematical Statistics, 22(3):327–351, September 1951. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177729580. See [And80].

Anderson:1980:CEL

[And80] T. W. Anderson. Correction to “Estimating Linear Restrictions on Regression Coefficients for Multivariate Normal Distributions”. Annals of Statistics, 8(6):1400, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:/

/projecteuclid.org/euclid.aos/1176345214. See correction

[And51].

Andersson:1982:DMI

[And82] Steen Andersson. Distributions of maximal invariants using quotient measures. Annals of Statistics, 10(3):955–961, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345885.

Anderson:1984:ELS

[And84] T. W. Anderson. Estimating linear statistical relationships. Annals of Statistics, 12(1):1–45, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346390.

Anderson:1987:DWA

[And87] T. W. Anderson. Discussion: What is an analysis of variance? Annals of Statistics, 15(3):911–913, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350473.

Anonymous:1983:RCI

[Ano83] Anonymous. Reply: “Construction of Improved Estimators in Multiparameter Estimation for Discrete Exponential Families”. Annals of Statistics, 11(2):375–376, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346147. See [GHT83].

Anonymous:1984:PWJ

[Ano84] Anonymous. The publications and writings of Jack Kiefer. Annals of Statistics, 12(2):424–430, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346497.

Akman:1986:AIC

[AR86] V. E. Akman and A. E. Raftery. Asymptotic inference for a changepoint Poisson process. Annals of Statistics, 14(4):1583–1590, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350178.

Assaf:1989:DSP

[AR89] David Assaf and Ya’acov Ritov. Dynamic sampling procedures for detecting a change in the drift of Brownian motion: A non-Bayesian model. Annals of Statistics, 17(2):793–800, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347143.

Agarwal:1980:AIM

[AS80] Girdhar G. Agarwal and W. J. Studden. Asymptotic integrated mean square error using least squares and bias minimizing splines. Annals of Statistics, 8(6):1307–1325, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345203.

Abramovitch:1985:ECP

[AS85] Lavy Abramovitch and Kesar Singh. Edgeworth corrected pivotal statistics and the bootstrap. Annals of Statistics, 13(1):116–132, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346580.

Armstrong:1989:LCR

[AS89] Thomas E. Armstrong and William D. Sudderth. Locally coherent rates of exchange. Annals of Statistics, 17(3):1394–1408, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347278.

Assaf:1988:DSA

[Ass88] David Assaf. A dynamic sampling approach for detecting a change in distribution. Annals of Statistics, 16(1):236–253, March 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350702.

Anderson:1985:BIE

[ASZ85] T. W. Anderson, Charles Stein, and Asad Zaman. Best invariant estimation of a direction parameter. Annals of Statistics, 13(2):526–533, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349536.

Athreya:1987:BMI

[Ath87] K. B. Athreya. Bootstrap of the mean in the infinite variance case. Annals of Statistics, 15(2):724–731, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350371.

Atkinson:1984:RTB

[Atk84] A. C. Atkinson. Review: Two books on regression diagnostics: D. A. Belsley, E. Kuh, R. E. Welsch, Regression Diagnostics: Identifying Influential Data and Sources of Collinearity; R. D. Cook, Sanford Weisberg, Residuals and Influence in Regression. Annals of Statistics, 12(1):392–401, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346418.

Atwood:1980:CDS

[Atw80] Corwin L. Atwood. Convergent design sequences, for sufficiently regular optimality criteria, II: Singular case. Annals of Statistics, 8(4):894–912, July 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345082.

Assaf:1985:OSS

[AZ85] David Assaf and Shmuel Zamir. Optimal sequential search: A Bayesian approach. Annals of Statistics, 13(3):1213–1221, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349665.

An:1982:AAA

[AZGH82] Hong-Zhi An, Chen Zhao-Guo, and E. J. Hannan. Autocorrelation, autoregression and autoregressive approximation. Annals of Statistics, 10(3):926–936, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:/

/projecteuclid.org/euclid.aos/1176345882. See correction

[AZGH83].

An:1983:CAA

[AZGH83] Hong-Zhi An, Chen Zhao-Guo, and E. J. Hannan. Correction: “Autocorrelation, Autoregression and Autoregressive Approximation”. Annals of Statistics, 11(3):1018, September 1983. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346271. See

[AZGH82].

Bailey:1983:AJD

[Bai83] Kent R. Bailey. The asymptotic joint distribution of regression and survival parameter estimates in the Cox regression model. Annals of Statistics, 11(1):39–48, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346054.

Bailey:1984:AEB

[Bai84] Kent R. Bailey. Asymptotic equivalence between the Cox estimator and the general ML estimators of regression and survival parameters in the Cox model. Annals of Statistics, 12(2):730–736, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346518.

Bailey:1987:DWA

[Bai87] R. A. Bailey. Discussion: What is an analysis of variance? Annals of Statistics, 15(3):913–916, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350474.

Barnard:1982:DFM

[Bar82] G. A. Barnard. Discussion: The functional-model basis of fiducial inference. Annals of Statistics, 10(4):1068–1069, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345971.

Barron:1986:DCB

[Bar86a] Andrew R. Barron. Discussion: On the consistency of Bayes estimates. Annals of Statistics, 14(1):26–30, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349831.

Barry:1986:NBR

[Bar86b] Daniel Barry. Nonparametric Bayesian regression. Annals of Statistics, 14(3):934–953, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350043.

Barron:1989:UPG

[Bar89] Andrew R. Barron. Uniformly powerful goodness of fit tests. Annals of Statistics, 17(1):107–124, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347005.

Basawa:1980:RBO

[Bas80] I. V. Basawa. Remarks on Bahadur optimality of conditional tests. Annals of Statistics, 8(6):1382–1387, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345209. See cor-

rection [Bas83].

Basawa:1983:CRB

[Bas83] I. V. Basawa. Corrections: Remarks on Bahadur efficiency of conditional tests. Annals of Statistics, 11(1):347, March 1983. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346088. See

[Bas80].

Basawa:1987:ADP

[Bas87] I. V. Basawa. Asymptotic distributions of prediction errors and related tests of fit for nonstationary processes. Annals of Statistics, 15(1):46–58, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350252.

Bemis:1983:BEC

[BB83] Kerry G. Bemis and Vasant P. Bhapkar. On ban estimators for chi squared test criteria. Annals of Statistics, 11(1):183–196, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346068.

Basawa:1984:ACI

[BB84] I. V. Basawa and P. J. Brockwell. Asymptotic conditional inference for regular nonergodic models with an application to autoregressive processes. Annals of Statistics, 12(1):161–171, March 1984. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346399.

Berger:1986:RBE

[BB86] James Berger and L. Mark Berliner. Robust Bayes and empirical Bayes analysis with _{ϵ}-contaminated priors. Annals of Statistics, 14(2):461–486, June 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349933.

Berk:1981:BST

[BBC81a] R. H. Berk, L. D. Brown, and Arthur Cohen. Bounded stopping times for a class of sequential Bayes tests. Annals of Statistics, 9(4):834–845, July 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345523.

Berk:1981:PBS

[BBC81b] R. H. Berk, L. D. Brown, and Arthur Cohen. Properties of Bayes sequential tests. Annals of Statistics, 9(3):678–682, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345473.

Bartoszynski:1981:SNT

[BBMT81] Robert Bartoszyński, Barry W. Brown, Charles M. McBride, and James R. Thompson. Some nonparametric techniques for estimating the intensity function of a cancer related nonstationary Poisson process. Annals of Statistics, 9(5):1050–1060, September 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345584.

Berger:1982:GAI

[BBZ82] James O. Berger, L. Mark Berliner, and Asad Zaman. General admissibility and inadmissibility results for estimation in a control problem. Annals of Statistics, 10(3):838–856, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345875.

Brown:1981:ILC

[BC81] L. D. Brown and Arthur Cohen. Inadmissibility of large classes of sequential tests. Annals of Statistics, 9(6):1239–1247, November 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345640.

Burman:1989:NER

[BC89] Prabir Burman and Keh-Wei Chen. Nonparametric estimation of a regression function. Annals of Statistics, 17(4):1567–1596, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347382.

Brown:1980:CCS

[BCS80] L. D. Brown, Arthur Cohen, and W. E. Strawderman. Complete classes for sequential tests of hypotheses. Annals of Statistics, 8(2):377–398, March 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176344959. See correction [BCS89].

Brown:1989:CCC

[BCS89] L. D. Brown, Arthur Cohen, and W. E. Strawderman. Correction: “Complete Classes for Sequential Tests of Hypotheses”. Annals of Statistics, 17(3):1414–1416, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347280. See [BCS80].

Brown:1983:SNC

[BCSC83] L. D. Brown, Arthur Cohen, and E. Samuel-Cahn. A sharp necessary condition for admissibility of sequential tests– necessary and sufficient conditions for admissibility of SPRT’s. Annals of Statistics, 11(2):640–653, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346169.

Bhattacharya:1983:NES

[BCY83] P. K. Bhattacharya, Herman Chernoff, and S. S. Yang. Nonparametric estimation of the slope of a truncated regression. Annals of Statistics, 11(2):505–514, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346157.

Bunke:1984:BCV

[BD84] Olaf Bunke and Bernd Droge. Bootstrap and cross-validation estimates of the prediction error for linear regression models. Annals of Statistics, 12(4):1400–1424, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346800.

Broniatowski:1989:RBS

[BDD89] Michel Broniatowski, Paul Deheuvels, and Luc Devroye. On the relationship between stability of extreme order statistics and convergence of the maximum likelihood kernel density estimate. Annals of Statistics, 17(3):1070–1086, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347256.

Bednarski:1982:BEM

[Bed82] Tadeusz Bednarski. Binary experiments, minimax tests and 2-alternating capacities. Annals of Statistics, 10(1):226–232, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345705.

Beder:1987:SEM

[Bed87] Jay H. Beder. A sieve estimator for the mean of a Gaussian process. Annals of Statistics, 15(1):59–78, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350253.

Beder:1988:SEC

[Bed88] Jay H. Beder. A sieve estimator for the covariance of a Gaussian process. Annals of Statistics, 16(2):648–660, June 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350825.

Bell:1984:SEN

[Bel84] William Bell. Signal extraction for nonstationary time series. Annals of Statistics, 12(2):646–664, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:/

/projecteuclid.org/euclid.aos/1176346512. See correction

[Bel91].

Bell:1991:CSE

[Bel91] William Bell. Correction: “Signal Extraction for Nonstationary Time Series”. Annals of Statistics, 19(4):2280, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176348401. See

[Bel84].

Beran:1980:ALB

[Ber80a] Rudolf Beran. Asymptotic lower bounds for risk in robust estimation. Annals of Statistics, 8(6):1252–1264, November 1980. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345198.

Berger:1980:IIE

[Ber80b] James Berger. Improving on inadmissible estimators in continuous exponential families with applications to simultaneous estimation of gamma scale parameters. Annals of Statistics, 8 (3):545–571, May 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345008.

Berger:1980:RGB

[Ber80c] James Berger. A robust generalized Bayes estimator and confidence region for a multivariate normal mean. Annals of Statistics, 8(4):716–761, July 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345068.

Berkson:1980:MCS

[Ber80d] Joseph Berkson. Minimum chi-square, not maximum likelihood! Annals of Statistics, 8(3):457–487, May 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345003.

Beran:1982:ESD

[Ber82a] Rudolf Beran. Estimated sampling distributions: The bootstrap and competitors. Annals of Statistics, 10(1):212–225, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345704.

Beran:1982:REM

[Ber82b] Rudolf Beran. Robust estimation in models for independent nonidentically distributed data. Annals of Statistics, 10(2):415–428, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345783.

Berger:1982:SME

[Ber82c] James O. Berger. Selecting a minimax estimator of a multivariate normal mean. Annals of Statistics, 10(1):81–92, March 1982.

CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345691.

Berger:1983:DCI

[Ber83a] James O. Berger. Discussion: Construction of improved estimators in multiparameter estimation for discrete exponential families. Annals of Statistics, 11(2):368–369, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346144. See [GHT83].

Berliner:1983:IIE

[Ber83b] L. Mark Berliner. Improving on inadmissible estimators in the control problem. Annals of Statistics, 11(3):814–826, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346248.

Beran:1984:JAB

[Ber84] Rudolf Beran. Jackknife approximations to bootstrap estimates. Annals of Statistics, 12(1):101–118, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346395.

Beran:1986:DJB

[Ber86a] Rudolf Beran. Discussion: Jackknife, bootstrap and other resampling methods in regression analysis. Annals of Statistics, 14(4): 1295–1298, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350143.

Beran:1986:SPF

[Ber86b] Rudolf Beran. Simulated power functions. Annals of Statistics, 14(1):151–173, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349847.

Berger:1986:DCB

[Ber86c] James Berger. Discussion: On the consistency of Bayes estimates. Annals of Statistics, 14(1):30–37, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349832.

Beran:1988:DTC

[Ber88] Rudolf Beran. Discussion: Theoretical comparison of bootstrap confidence intervals. Annals of Statistics, 16(3):956–959, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350935.

Bertelsen:1989:NND

[Ber89] Aksel Bertelsen. On non-null distributions connected with testing reality of a complex normal distribution. Annals of Statistics, 17(2):929–936, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347152.

Bhattacharya:1981:NCC

[BF81a] P. K. Bhattacharya and Dargan Frierson. A nonparametric control chart for detecting small disorders. Annals of Statistics, 9(3):544–554, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345458.

Bickel:1981:SAT

[BF81b] Peter J. Bickel and David A. Freedman. Some asymptotic theory for the bootstrap. Annals of Statistics, 9(6):1196–1217, November 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345637.

Bickel:1984:ANB

[BF84] P. J. Bickel and D. A. Freedman. Asymptotic normality and the bootstrap in stratified sampling. Annals of Statistics, 12 (2):470–482, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346500.

Brown:1985:AAL

[BF85a] L. D. Brown and R. H. Farrell. All admissible linear estimators of a multivariate Poisson mean. Annals of Statistics, 13 (1):282–294, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346593.

Brown:1985:CCT

[BF85b] L. D. Brown and R. H. Farrell. Complete class theorems for estimation of multivariate Poisson means and related problems. Annals of Statistics, 13(2):706–726, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349549.

Boente:1989:RNR

[BF89] Graciela Boente and Ricardo Fraiman. Robust nonparametric regression estimation for dependent observations. Annals of Statistics, 17(3):1242–1256, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347266.

Boente:1987:QRS

[BFY87] Graciela Boente, Ricardo Fraiman, and Victor J. Yohai. Qualitative robustness for stochastic processes. Annals of Statistics, 15(3): 1293–1312, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350506.

Bhattacharya:1978:VFE

[BG78] R. N. Bhattacharya and J. K. Ghosh. On the validity of the formal Edgeworth expansion. Annals of Statistics, 6(2):434–451, March 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.

aos/1176344134. See correction [BG80].

Bhattacharya:1980:CVF

[BG80] R. N. Bhattacharya and J. K. Ghosh. Correction to “On the Validity of the Formal Edgeworth Expansion”. Annals of Statistics, 8(6):1399, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345213. See [BG78].

Barnard:1982:MAA

[BG82] G. A. Barnard and V. P. Godambe. Memorial article: Allan Birnbaum 1923–1976. Annals of Statistics, 10(4):1033–1039, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345968.

Buckland:1988:DTC

[BGL88] Stephen T. Buckland, Paul H. Garthwaite, and Howard G. Lovell. Discussion: Theoretical comparison of bootstrap confidence intervals. Annals of Statistics, 16(3):962–965, September 1988. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350937.

Bickel:1986:EES

[BGvZ86] P. J. Bickel, F. Gotze, and W. R. van Zwet. The Edgeworth expansion for U-statistics of degree two. Annals of Statistics, 14(4): 1463–1484, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350170.

Brown:1989:UDS

[BH89] Lawrence D. Brown and Jiunn T. Hwang. Universal domination and stochastic domination: U-admissibility and U- inadmissibility of the least squares estimator. Annals of Statistics, 17 (1):252–267, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347014.

Bhattacharya:1980:ECM

[Bha80] C. G. Bhattacharya. Estimation of a common mean and recovery of interblock information. Annals of Statistics, 8(1):205–211, January 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344903.

Bhattacharya:1981:PDD

[Bha81] P. K. Bhattacharya. Posterior distribution of a Dirichlet process from quantal response data. Annals of Statistics, 9(4):803–811, July 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345520.

Bhansali:1982:ECQ

[Bha82] R. J. Bhansali. The evaluation of certain quadratic forms occurring in autoregressive model fitting. Annals of Statistics, 10 (1):121–131, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345695.

Bhattacharya:1983:JTT

[Bha83] P. K. Bhattacharya. Justification for a K − S type test for the slope of a truncated regression. Annals of Statistics, 11(2):697– 701, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346174.

Bhansali:1986:AES

[Bha86] R. J. Bhansali. Asymptotically efficient selection of the order by the criterion autoregressive transfer function. Annals of Statistics, 14(1):315–325, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349858.

Begun:1983:IAE

[BHHW83] Janet M. Begun, W. J. Hall, Wei-Min Huang, and Jon A. Wellner. Information and asymptotic efficiency in parametricnonparametric models. Annals of Statistics, 11(2):432–452, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346151.

Buja:1989:LSA

[BHT89a] Andreas Buja, Trevor Hastie, and Robert Tibshirani. Linear smoothers and additive models. Annals of Statistics, 17(2):453– 510, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347115.

Buja:1989:RLS

[BHT89b] Andreas Buja, Trevor Hastie, and Robert Tibshirani. Rejoinder: Linear smoothers and additive models. Annals of Statistics, 17(2):543–555, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347124.

Bickel:1981:MEM

[Bic81] P. J. Bickel. Minimax estimation of the mean of a normal distribution when the parameter space is restricted. Annals of Statistics, 9(6):1301–1309, November 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345646.

Bickel:1982:AE

[Bic82] P. J. Bickel. On adaptive estimation. Annals of Statistics, 10(3): 647–671, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345863.

Bickel:1984:PRS

[Bic84a] P. J. Bickel. Parametric robustness: Small biases can be worthwhile. Annals of Statistics, 12(3):864–879, September 1984. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346707.

Bickel:1984:RJP

[Bic84b] P. J. Bickel. Review: J. Pfanzagl, W. Wefelmeyer, Contributions to a General Asymptotic Statistical Theory. Springer Lecture Notes in Statistics. Annals of Statistics, 12(2):786–791, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346530.

Bickel:1984:RRB

[Bic84c] P. J. Bickel. Robust regression based on infinitesimal neighbourhoods. Annals of Statistics, 12(4):1349–1368, December 1984. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346796.

Bickel:1988:DTC

[Bic88] Peter J. Bickel. Discussion: Theoretical comparison of bootstrap confidence intervals. Annals of Statistics, 16(3):959–961, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350936.

Bingham:1985:RGW

[Bin85] Christopher Bingham. Review: Geoffrey S. Watson, Statistics on Spheres. Annals of Statistics, 13(2):838–844, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349566.

Birge:1987:EDU

[Bir87a] Lucien Birge. Estimating a density under order restrictions: Nonasymptotic minimax risk. Annals of Statistics, 15(3):995–1012, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350488.

Birge:1987:RHE

[Bir87b] Lucien Birge. On the risk of histograms for estimating decreasing densities. Annals of Statistics, 15(3):1013–1022, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350489.

Birge:1989:GEN

[Bir89] Lucien Birge. The Grenander estimator: A nonasymptotic approach. Annals of Statistics, 17(4):1532–1549, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347380.

Bjornstad:1981:CSP

[Bjø81] Jan F. Bjørnstad. A class of Schur procedures and minimax theory for subset selection. Annals of Statistics, 9(4):777–791, July 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345518.

Bjornstad:1984:GTA

[Bjø84] Jan F. Bjørnstad. A general theory of asymptotic consistency for subset selection with applications. Annals of Statistics, 12(3): 1058–1070, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346721.

Baksalary:1980:NBE

[BK80] J. K. Baksalary and R. Kala. A new bound for the Euclidean norm of the difference between the least squares and the best linear unbiased estimators. Annals of Statistics, 8(3):679–681, May 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345018.

Baksalary:1981:LTP

[BK81] J. K. Baksalary and R. Kala. Linear transformations preserving best linear unbiased estimators in a general Gauss–Markoff model. Annals of Statistics, 9(4):913–916, July 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345533.

Baksalary:1983:ELC

[BK83] J. K. Baksalary and R. Kala. Estimation via linearly combining two given statistics. Annals of Statistics, 11(2):691–696, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346173.

Bickel:1981:MPS

[BL81] P. J. Bickel and E. L. Lehmann. A minimax property of the sample mean in finite populations. Annals of Statistics, 9(5): 1119–1122, September 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345592.

Bar-Lev:1983:CCS

[BL83] Shaul K. Bar-Lev. A characterization of certain statistics in exponential models whose distributions depend on a sub-vector of parameters only. Annals of Statistics, 11(3):746–752, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346242.

Brunner:1989:BMS

[BL89] Lawrence J. Brunner and Albert Y. Lo. Bayes methods for a symmetric unimodal density and its mode. Annals of Statistics, 17(4): 1550–1566, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347381.

Bar-Lev:1989:RVF

[BLB89] Shaul K. Bar-Lev and Daoud Bshouty. Rational variance functions. Annals of Statistics, 17(2):741–748, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347139.

Bar-Lev:1986:RNE

[BLE86] Shaul K. Bar-Lev and Peter Enis. Reproducibility and natural exponential families with power variance functions. Annals of Statistics, 14(4):1507–1522, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350173.

Blom:1987:HC

[Blo87] Gunnar Blom. Harald Cramér 1893–1985. Annals of Statistics, 15 (4):1335–1350, December 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350596.

Bar-Lev:1982:ESW

[BLR82] Shaul K. Bar-Lev and Benjamin Reiser. An exponential subfamily which admits UMPU tests based on a single test statistic. Annals of Statistics, 10(3):979–989, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345888.

Breslow:1985:DTI

[BM85] Norman Breslow and Dirk Moore. Discussion: Testing for independence in a two-way table: New interpretations of the chi-square statistic. Annals of Statistics, 13(3):875–877, September 1985. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349635.

Beran:1986:CSM

[BM86] R. Beran and P. W. Millar. Confidence sets for a multivariate distribution. Annals of Statistics, 14(2):431–443, June 1986. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349931.

Beran:1987:SET

[BM87a] R. Beran and P. W. Millar. Stochastic estimation and testing. Annals of Statistics, 15(3):1131–1154, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350497.

Bhattacharya:1987:WCN

[BM87b] P. K. Bhattacharya and Y. P. Mack. Weak convergence of k-NN density and regression estimators with varying k and applications. Annals of Statistics, 15(3):976–994, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350487.

Beran:1989:SMD

[BM89a] R. Beran and P. W. Millar. A stochastic minimum distance test for multivariate parametric models. Annals of Statistics, 17 (1):125–140, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347006.

Brown:1989:CCR

[BM89b] Lawrence D. Brown and John I. Marden. Complete class results for hypothesis testing problems with simple null hypotheses. Annals of Statistics, 17(1):209–235, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347012.

Basawa:1989:BEA

[BMMT89] I. V. Basawa, A. K. Mallik, W. P. McCormick, and R. L. Taylor. Bootstrapping explosive autoregressive processes. Annals of Statistics, 17(4):1479–1486, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347376.

Boyles:1985:IML

[BMP85] R. A. Boyles, A. W. Marshall, and F. Proschan. Inconsistency of the maximum likelihood estimator of a distribution having increasing failure rate average. Annals of Statistics, 13(1):413–417, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346602.

Bobrovsky:1987:SCG

[BMWZ87] B. Z. Bobrovsky, E. Mayer-Wolf, and M. Zakai. Some classes of global Cramér–Rao bounds. Annals of Statistics, 15(4):1421–1438, December 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350602.

Behnen:1983:GTL

[BN83] Konrad Behnen and Georg Neuhaus. Galton’s test as a linear rank test with estimated scores and its local asymptotic efficiency. Annals of Statistics, 11(2):588–599, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346164.

Barndorff-Nielsen:1986:LOG

[BN86] O. E. Barndorff-Nielsen. Likelihood and observed geometries. Annals of Statistics, 14(3):856–873, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350038.

Barndorff-Nielsen:1983:EMA

[BNB83a] O. Barndorff-Nielsen and P. Blaesild. Exponential models with affine dual foliations. Annals of Statistics, 11(3):753–769, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346243.

Barndorff-Nielsen:1983:REF

[BNB83b] O. Barndorff-Nielsen and P. Blaesild. Reproductive exponential families. Annals of Statistics, 11(3):770–782, September 1983. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346244.

Barndorff-Nielsen:1988:CRM

[BNB88] O. E. Barndorff-Nielsen and P. Blaesild. Combination of reproductive models. Annals of Statistics, 16(1):323–341, March 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350708.

Barndorff-Nielsen:1988:DGP

[BNJ88] O. E. Barndorff-Nielsen and P. E. Jupp. Differential geometry, profile likelihood, L-sufficiency and composite transformation models. Annals of Statistics, 16(3):1009–1043, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350946.

Behnen:1983:TSR

[BNR83] Konrad Behnen, Georg Neuhaus, and Frits Ruymgaart. Two sample rank estimators of optimal nonparametric score-functions and corresponding adaptive rank statistics. Annals of Statistics, 11(4): 1175–1189, December 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346330.

Bai:1988:DTC

[BO88] Chongen Bai and Richard A. Olshen. Discussion: Theoretical comparison of bootstrap confidence intervals. Annals of Statistics, 16 (3):953–956, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350934.

Bohning:1982:CSA

[Boh82] Dankmar Bohning. Convergence of Simar’s algorithm for finding the maximum likelihood estimate of a compound Poisson process. Annals of Statistics, 10(3):1006–1008, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345890.

Bookstein:1985:DPP

[Boo85a] Fred L. Bookstein. Discussion: Projection pursuit. Annals of Statistics, 13(2):481–484, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349521.

Boos:1985:CST

[Boo85b] Dennis D. Boos. A converse to Scheffe’s theorem. Annals of Statistics, 13(1):423–427, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346604.

Bose:1988:ECB

[Bos88] Arup Bose. Edgeworth correction by bootstrap in autoregressions. Annals of Statistics, 16(4):1709–1722, December 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176351063.

Berger:1984:MSP

[BP84] Roger L. Berger and Frank Proschan. Monotonicity in selection problems: A unified approach. Annals of Statistics, 12 (1):387–391, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346417.

Bhattacharya:1989:SOC

[BQ89] Rabi Bhattacharya and Maher Qumsiyeh. Second order and L^{p}-comparisons between the bootstrap and empirical Edgeworth expansion methodologies. Annals of Statistics, 17(1):160–169, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347008.

Blackwell:1982:BCS

[BR82] D. Blackwell and R. V. Ramamoorthi. A Bayes but not classically sufficient statistic. Annals of Statistics, 10(3):1025–1026, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345895.

Bickel:1987:EEE

[BR87] P. J. Bickel and Y. Ritov. Efficient estimation in the errors in variables model. Annals of Statistics, 15(2):513–540, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350358.

Brant:1984:ALP

[Bra84] Rollin Brant. Approximate likelihood and probability calculations based on transforms. Annals of Statistics, 12(3):989–1005, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346716.

Breiman:1989:DLS

[Bre89] Leo Breiman. Discussion: Linear smoothers and additive models. Annals of Statistics, 17(2):510–515, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347116.

Borgan:1985:DIR

[BRH85] Ornulf Borgan and Henrik Ramlau-Hansen. Demographic incidence rates and estimation of intensities with incomplete information. Annals of Statistics, 13(2):564–582, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349539.

Brillinger:1986:DIF

[Bri86] David R. Brillinger. Discussion: Influence functionals for time series. Annals of Statistics, 14(3):819–822, September 1986. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350028.

Brillinger:1987:DWA

[Bri87] David R. Brillinger. Discussion: What is an analysis of variance? Annals of Statistics, 15(3):916–917, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350475.

Brillinger:1988:SSM

[Bri88] David R. Brillinger. Some statistical methods for random process data from seismology and neurophysiology. Annals of Statistics, 16(1):1–54, March 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350689.

Brown:1980:EBP

[Bro80a] L. Brown. Examples of Berger’s phenomenon in the estimation of independent normal means. Annals of Statistics, 8(3):572–585, May 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345009.

Brown:1980:NCA

[Bro80b] Lawrence D. Brown. A necessary condition for admissibility. Annals of Statistics, 8(3):540–544, May 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345007.

Brown:1981:CCT

[Bro81] Lawrence D. Brown. A complete class theorem for statistical problems with finite sample spaces. Annals of Statistics, 9(6): 1289–1300, November 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345645.

Brockett:1984:LRD

[Bro84a] Patrick L. Brockett. The likelihood ratio detector for non-Gaussian infinitely divisible, and linear stochastic processes. Annals of Statistics, 12(2):737–744, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346519.

Brown:1984:AVM

[Bro84b] K. G. Brown. On analysis of variance in the mixed model. Annals of Statistics, 12(4):1488–1499, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346805.

Brown:1984:RJK

[Bro84c] Lawrence D. Brown. The research of Jack Kiefer outside the area of experimental design. Annals of Statistics, 12(2):406–415, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346495.

Brown:1985:MVB

[Bro85] Mark Brown. A measure of variability based on the harmonic mean, and its use in approximations. Annals of Statistics, 13(3): 1239–1243, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349668.

Brown:1988:ADC

[Bro88] Lawrence D. Brown. Admissibility in discrete and continuous invariant nonparametric estimation problems and in their multinomial analogs. Annals of Statistics, 16(4):1567–1593, December 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176351054.

Bruvold:1980:ADP

[Bru80] Norman T. Bruvold. Admissible designs for polynomial monospline regression. Annals of Statistics, 8(4):913–921, July 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345083.

Boos:1980:NDC

[BS80a] Dennis D. Boos and R. J. Serfling. A note on differentials and the CLT and LIL for statistical functions, with application to Mestimates. Annals of Statistics, 8(3):618–624, May 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345012.

Brandwein:1980:MEL

[BS80b] Ann Cohen Brandwein and William E. Strawderman. Minimax estimation of location parameters for spherically symmetric distributions with concave loss. Annals of Statistics, 8(2):279–284, March 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344953.

Babu:1983:IMU

[BS83a] G. Jogesh Babu and Kesar Singh. Inference on means using the bootstrap. Annals of Statistics, 11(3):999–1003, September 1983.

CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346267.

Barnard:1983:GPN

[BS83b] G. A. Barnard and D. A. Sprott. The generalised problem of the Nile: Robust confidence sets for parametric functions. Annals of Statistics, 11(1):104–113, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346061.

Brown:1984:AST

[BS84] Lawrence D. Brown and Harold Sackrowitz. An alternative to Student’s t-test for problems with indifference zones. Annals of Statistics, 12(2):451–469, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346499.

Beran:1985:BTC

[BS85a] Rudolf Beran and Muni S. Srivastava. Bootstrap tests and confidence regions for functions of a covariance matrix. Annals of Statistics, 13(1):95–115, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346579. See correction [BS87].

Buja:1985:DPP

[BS85b] Andreas Buja and Werner Stuetzle. Discussion: Projection pursuit. Annals of Statistics, 13(2):484–490, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349522.

Bailey:1986:RLD

[BS86] R. A. Bailey and T. P. Speed. Rectangular lattice designs: Efficiency factors and analysis. Annals of Statistics, 14(3):874–895, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350039.

Beran:1987:CBT

[BS87] Rudolf Beran and Muni S. Srivastava. Correction: “Bootstrap Tests and Confidence Regions for Functions of a Covariance Matrix”. Annals of Statistics, 15(1):470–471, March 1987. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350284. See

[BS85a].

Babu:1989:EEM

[BS89] G. J. Babu and K. Singh. On Edgeworth expansions in the mixture cases. Annals of Statistics, 17(1):443–447, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347029.

Butler:1981:ABC

[But81] Ronald W. Butler. The admissible Bayes character of subset selection techniques involved in variable selection, outlier detection, and slippage problems. Annals of Statistics, 9(5):960–973, September 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345576.

Butler:1982:NIP

[But82] Ronald W. Butler. Nonparametric interval and point prediction using data trimmed by a Grubbs-type outlier rule. Annals of Statistics, 10(1):197–204, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345702.

Butler:1983:OPO

[But83] Ronald W. Butler. Optimal properties of one-step variable selection in regression. Annals of Statistics, 11(1):219–224, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346072.

Bates:1981:PTI

[BW81] Douglas M. Bates and Donald G. Watts. Parameter transformations for improved approximate confidence regions in nonlinear least squares. Annals of Statistics, 9(6):1152–1167, November 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345633.

Berry:1982:OSR

[BW82a] Donald A. Berry and PeCheng Wang. Optimal stopping regions with islands and peninsulas. Annals of Statistics, 10(2):634–636, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345806.

Bohrer:1982:SSP

[BW82b] Robert Bohrer and Henry P. Wynn. Spherically symmetric probability orderings useful in multiple comparisons. Annals of Statistics, 10(4):1253–1260, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345990.

Bradley:1980:TFB

[BY80] Ralph A. Bradley and Ching-Ming Yeh. Trend-free block designs: Theory. Annals of Statistics, 8(4):883–893, July 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345081.

Brown:1980:AMR

[BZ80] P. J. Brown and J. V. Zidek. Adaptive multivariate ridge regression. Annals of Statistics, 8(1):64–74, January 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344891.

Campbell:1982:RCE

[Cam82] Katherine Campbell. Recursive computation of M-estimates for the parameters of a finite autoregressive process. Annals of Statistics, 10(2):442–453, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345785.

Caperaa:1988:TOA

[Cap88] Philippe Caperaa. Tail ordering and asymptotic efficiency of rank tests. Annals of Statistics, 16(1):470–478, March 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350715.

Carroll:1982:AHL

[Car82] Raymond J. Carroll. Adapting for heteroscedasticity in linear models. Annals of Statistics, 10(4):1224–1233, December 1982. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345987.

Carlstein:1986:USV

[Car86] Edward Carlstein. The use of subseries values for estimating the variance of a general statistic from a stationary sequence. Annals of Statistics, 14(3):1171–1179, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350057.

Carlstein:1988:NCP

[Car88] E. Carlstein. Nonparametric change-point estimation. Annals of Statistics, 16(1):188–197, March 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350699.

Casella:1980:MRR

[Cas80] George Casella. Minimax ridge regression estimation. Annals of Statistics, 8(5):1036–1056, September 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345141.

Casella:1987:CAR

[Cas87] George Casella. Conditionally acceptable recentered set estimators. Annals of Statistics, 15(4):1363–1371, December 1987. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350598.

Clayton:1985:BNB

[CB85] Murray K. Clayton and Donald A. Berry. Bayesian nonparametric bandits. Annals of Statistics, 13(4):1523–1534, December 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349753.

Cotterill:1982:LDC

[CC82] Derek S. Cotterill and Miklós Csörgő. On the limiting distribution of and critical values for the multivariate Cramér–von Mises statistic. Annals of Statistics, 10(1):233–244, March 1982. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345706.

Coster:1988:MCT

[CC88] Daniel C. Coster and Ching-Shui Cheng. Minimum cost trendfree run orders of fractional factorial designs. Annals of Statistics, 16(3):1188–1205, September 1988. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350955.

Chauhan:1986:OFE

[CD86] Chand K. Chauhan and A. M. Dean. Orthogonality of factorial effects. Annals of Statistics, 14(2):743–752, June 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349951.

Csorgo:1985:KET

[CDM85] Sandor Csorgo, Paul Deheuvels, and David Mason. Kernel estimates of the tail index of a distribution. Annals of Statistics, 13(3): 1050–1077, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349656.

Crook:1980:ASD

[CG80] J. F. Crook and I. J. Good. On the application of symmetric Dirichlet distributions and their mixtures to contingency tables, Part II. Annals of Statistics, 8(6):1198–1218, November 1980. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345194. See

[CG81].

Crook:1981:CAS

[CG81] J. F. Crook and I. J. Good. Corrections to “On the Application of Symmetric Dirichlet Distributions and Their Mixtures to Contingency Tables, Part II”. Annals of Statistics, 9(5):1133, September 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.

aos/1176345597. See correction [CG80].

Chow:1985:SMS

[CG85] Yun-Shyong Chow and Ulf Grenander. A sieve method for the spectral density. Annals of Statistics, 13(3):998–1010, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349652.

Cook:1986:CPS

[CG86] R. Dennis Cook and Miriam L. Goldberg. Curvatures for parameter subsets in nonlinear regression. Annals of Statistics, 14(4): 1399–1418, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350166.

Chow:1983:CCV

[CGW83] Y.-S. Chow, S. Geman, and L.-D. Wu. Consistent cross-validated density estimation. Annals of Statistics, 11(1):25–38, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346053.

Chen:1989:DLS

[CGW89] Zehua Chen, Chong Gu, and Grace Wahba. Discussion: Linear smoothers and additive models. Annals of Statistics, 17 (2):515–522, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347117.

Christopeit:1980:SCL

[CH80] N. Christopeit and K. Helmes. Strong consistency of least squares estimators in linear regression models. Annals of Statistics, 8 (4):778–788, July 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345070.

Chao:1981:EPD

[Cha81] Anne Chao. On estimating the probability of discovering a new species. Annals of Statistics, 9(6):1339–1342, November 1981. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345651. See

corrections [Cha82].

Chao:1982:CEP

[Cha82] Anne Chao. Corrections: “On Estimating the Probability of Discovering a New Species”. Annals of Statistics, 10(4):1311, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345999. See [Cha81].

Chaloner:1984:OBE

[Cha84] Kathryn Chaloner. Optimal Bayesian experimental design for linear models. Annals of Statistics, 12(1):283–300, March 1984. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346407. See

correction [Cha85].

Chaloner:1985:COB

[Cha85] Kathryn Chaloner. Correction: “Optimal Bayesian Experimental Design for Linear Models”. Annals of Statistics, 13(2):836, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349564. See [Cha84].

Chang:1986:SR

[Cha86] Ted Chang. Spherical regression. Annals of Statistics, 14(3): 907–924, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350041.

Chang:1989:SRE

[Cha89] Ted Chang. Spherical regression with errors in variables. Annals of Statistics, 17(1):293–306, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347017.

Cheng:1980:OSW

[Che80a] Ching-Shui Cheng. Optimality of some weighing and 2^{n} fractional factorial designs. Annals of Statistics, 8(2):436–446, March 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344963.

Cheng:1980:OAV

[Che80b] Ching-Shui Cheng. Orthogonal arrays with variable numbers of symbols. Annals of Statistics, 8(2):447–453, March 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344964.

Cheng:1980:AEI

[Che80c] H. H. Peter Cheng. Asymptotic expected inferior sample size of a sequential test involving two populations. Annals of Statistics, 8(4):845–850, July 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345077.

Chernoff:1980:IEL

[Che80d] Herman Chernoff. The identification of an element of a large population in the presence of noise. Annals of Statistics, 8(6): 1179–1197, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345193.

Cheng:1981:OCP

[Che81a] Ching-Shui Cheng. Optimality and construction of pseudo-Youden designs. Annals of Statistics, 9(1):201–205, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345348.

Cheng:1981:BER

[Che81b] K. F. Cheng. On Berry–Esseen rates for jackknife estimators. Annals of Statistics, 9(3):694–696, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345477.

Cheng:1983:COB

[Che83] Ching-Shui Cheng. Construction of optimal balanced incomplete block designs for correlated observations. Annals of Statistics, 11 (1):240–246, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346074.

Cheng:1987:AKW

[Che87a] Ching-Shui Cheng. An application of the Kiefer–Wolfowitz equivalence theorem to a problem in Hadamard transform optics. Annals of Statistics, 15(4):1593–1603, December 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350612.

Cheng:1987:OPA

[Che87b] Ching-Shui Cheng. An optimization problem with applications to optimal design theory. Annals of Statistics, 15(2):712–723, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350370.

Chen:1988:CRP

[Che88a] Hung Chen. Convergence rates for parametric components in a partly linear model. Annals of Statistics, 16(1):136–146, March 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350695.

Chen:1988:LRC

[Che88b] Hung Chen. Lower rate of convergence for locating a maximum of a function. Annals of Statistics, 16(3):1330–1334, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350965.

Chikuse:1981:DSM

[Chi81] Yasuko Chikuse. Distributions of some matrix variates and latent roots in multivariate Behrens–Fisher discriminant analysis. Annals of Statistics, 9(2):401–407, March 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345405.

Chiu:1986:SEP

[Chi86] Shean-Tsong Chiu. Statistical estimation of the parameters of a moving source from array data. Annals of Statistics, 14(2):559– 578, June 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349938.

Chiu:1988:WLS

[Chi88] Shean-Tsong Chiu. Weighted least squares estimators on the frequency domain for the parameters of a time series. Annals of Statistics, 16(3):1315–1326, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350963.

Chen:1983:CTW

[CHL83a] Yuan Yan Chen, Myles Hollander, and Naftali A. Langberg. Corrections: “Testing Whether New is Better than Used with Randomly Censored Data”. Annals of Statistics, 11(4):1267, December 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346341. See [CHL83b].

Chen:1983:TWN

[CHL83b] Yuan Yan Chen, Myles Hollander, and Naftali A. Langberg. Testing whether new is better than used with randomly censored data. Annals of Statistics, 11(1):267–274, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346077. See cor-

rection [CHL83a].

Chow:1987:CCT

[Cho87a] Mo Suk Chow. A complete class theorem for estimating a noncentrality parameter. Annals of Statistics, 15(2):800–804, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350375.

Chow:1987:ET

[Cho87b] Yunshyong Chow. Estimating trajectories. Annals of Statistics, 15(2):552–567, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350360.

Christensen:1986:FSS

[Chr86] Ronald Christensen. Finite stopping in sequential sampling without recall from a Dirichlet process. Annals of Statistics, 14 (1):275–282, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349855.

Christensen:1989:LFT

[Chr89] Ronald Christensen. Lack-of-fit tests based on near or exact replicates. Annals of Statistics, 17(2):673–683, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347133.

Christian:1989:STM

[CJ89] Niels Christian and Bang Jespersen. On the structure of transformation models. Annals of Statistics, 17(1):195–208, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347011.

Callaert:1980:EES

[CJV80] H. Callaert, P. Janssen, and N. Veraverbeke. An Edgeworth expansion for U-statistics. Annals of Statistics, 8(2):299–312, March 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344955.

Cohen:1985:AED

[CK85] Michael P. Cohen and Lynn Kuo. The admissibility of the empirical distribution function. Annals of Statistics, 13(1):262–271, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346591.

Cox:1988:TPN

[CKWY88] Dennis Cox, Eunmee Koh, Grace Wahba, and Brian S. Yandell. Testing the (parametric) null model hypothesis in (semiparametric) partial and generalized spline models. Annals of Statistics, 16(1):113–119, March 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350693.

Cheng:1983:MAS

[CL83] Ching-Shui Cheng and Ker-Chau Li. A minimax approach to sample surveys. Annals of Statistics, 11(2):552–563, June 1983. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346160.

Chao:1988:SRN

[CL88] Min-Te Chao and Shaw-Hwa Lo. Some representations of the nonparametric maximum likelihood estimators with truncated data. Annals of Statistics, 16(2):661–668, June 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350826.

Clarke:1983:UFD

[Cla83] Brenton R. Clarke. Uniqueness and Frechet differentiability of functional solutions to maximum likelihood type equations. Annals of Statistics, 11(4):1196–1205, December 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346332.

Clayton:1985:BNS

[Cla85] Murray K. Clayton. A Bayesian nonparametric sequential test for the mean of a population. Annals of Statistics, 13(3):1129–1139, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349660.

Clayton:1986:DCB

[Cla86] Murray K. Clayton. Discussion: On the consistency of Bayes estimates. Annals of Statistics, 14(1):37–40, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349833.

Cline:1988:AKE

[Cli88] Daren B. H. Cline. Admissible kernel estimators of a multivariate density. Annals of Statistics, 16(4):1421–1427, December 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176351046.

Cohen:1985:EQS

[CLS85] Arthur Cohen, Shaw-Hwa Lo, and Kesar Singh. Estimating a quantile of a symmetric distribution. Annals of Statistics, 13(3): 1114–1128, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349659.

Cohen:1986:APE

[CLS86] Arthur Cohen, Shaw-Hwa Lo, and Kesar Singh. Acknowledgment of priority: Estimating a quantile of a symmetric distribution. Annals of Statistics, 14(4):1642, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350186.

Chow:1982:BRS

[CM82] Y. S. Chow and A. T. Martinsek. Bounded regret of a sequential procedure for estimation of the mean. Annals of Statistics, 10(3): 909–914, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345880.

Cho:1987:MFO

[CM87] Sinsup Cho and Robert B. Miller. Model-free one-step-ahead prediction intervals: Asymptotic theory and small sample simulations. Annals of Statistics, 15(3):1064–1078, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350493.

Cohen:1989:ACT

[CM89a] Arthur Cohen and John I. Marden. On the admissibility and consistency of tests for homogeneity of variances. Annals of Statistics, 17(1):236–251, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347013.

Csorgo:1989:BEF

[CM89b] Sandor Csorgo and David M. Mason. Bootstrapping empirical functions. Annals of Statistics, 17(4):1447–1471, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347374.

Collins:1983:MPE

[Col83] John R. Collins. On the minimax property for R-estimators of location. Annals of Statistics, 11(4):1190–1195, December 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346331.

Collings:1984:GII

[Col84] Bruce Jay Collings. Generating the intrablock and interblock subgroups for confounding in general factorial experiments. Annals of Statistics, 12(4):1500–1509, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346806.

Collins:1986:MAV

[Col86] John R. Collins. Maximum asymptotic variances of trimmed means under asymmetric contamination. Annals of Statistics, 14 (1):348–354, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349861.

Constantine:1981:SOB

[Con81] Gregory M. Constantine. Some E-optimal block designs. Annals of Statistics, 9(4):886–892, July 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345529.

Constantine:1982:OPD

[Con82] Gregory M. Constantine. On the E-optimality of PBIB designs with a small number of blocks. Annals of Statistics, 10(3): 1027–1031, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345896.

Cox:1983:ATS

[Cox83] Dennis D. Cox. Asymptotics for M-type smoothing splines. Annals of Statistics, 11(2):530–551, June 1983. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346159.

Cox:1985:DPP

[Cox85] D. R. Cox. Discussion: Projection pursuit. Annals of Statistics, 13(2):493–494, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349524.

Cox:1988:ALS

[Cox88a] Dennis D. Cox. Approximation of least squares regression on nested subspaces. Annals of Statistics, 16(2):713–732, June 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350830.

Cox:1988:AMR

[Cox88b] Dennis D. Cox. Approximation of method of regularization estimators. Annals of Statistics, 16(2):694–712, June 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350829.

Cox:1989:DLS

[Cox89] Dennis D. Cox. Discussion: Linear smoothers and additive models. Annals of Statistics, 17(2):522–525, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347118.

Collins:1981:MVE

[CP81] John R. Collins and Stephen L. Portnoy. Maximizing the variance of M-estimators using the generalized method of moment spaces. Annals of Statistics, 9(3):567–577, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345460.

Carroll:1981:RTH

[CR81] Raymond J. Carroll and David Ruppert. On robust tests for heteroscedasticity. Annals of Statistics, 9(1):206–210, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345349.

Carroll:1982:REH

[CR82] Raymond J. Carroll and David Ruppert. Robust estimation in heteroscedastic linear models. Annals of Statistics, 10(2):429–441, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345784.

Carroll:1986:DJB

[CR86] Raymond J. Carroll and David Ruppert. Discussion: Jackknife, bootstrap and other resampling methods in regression analysis. Annals of Statistics, 14(4):1298–1301, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350144.

Cressie:1982:UEB

[Cre82] Noel Cressie. A useful empirical Bayes identity. Annals of Statistics, 10(2):625–629, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345804.

Cristobal:1987:CLR

[CRM87] J. A. Cristobal Cristobal, P. Faraldo Roca, and W. Gonzalez Manteiga. A class of linear regression parameter estimators constructed by nonparametric estimation. Annals of Statistics, 15 (2):603–609, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350363.

Crowder:1980:APL

[Cro80] Martin J. Crowder. On the asymptotic properties of least-squares estimators in autoregression. Annals of Statistics, 8(1):132–146, January 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344896.

Casella:1981:EBN

[CS81] George Casella and William E. Strawderman. Estimating a bounded normal mean. Annals of Statistics, 9(4):870–878, July 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345527.

Cohen:1984:BDS

[CS84a] Arthur Cohen and Harold B. Sackrowitz. Bayes double sample estimation procedures. Annals of Statistics, 12(3):1035–1049, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346719.

Cohen:1984:RDS

[CS84b] Arthur Cohen and Harold B. Sackrowitz. Results on double sample estimation for the binomial distribution. Annals of Statistics, 12(3):1109–1116, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346728.

Cohen:1987:UTH

[CS87] Arthur Cohen and Harold B. Sackrowitz. Unbiasedness of tests for homogeneity. Annals of Statistics, 15(2):805–816, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350376.

Csiszar:1989:GID

[Csi89] Imre Csiszar. A geometric interpretation of Darroch and Ratcliff’s generalized iterative scaling. Annals of Statistics, 17(3): 1409–1413, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347279.

Csorgo:1986:TNA

[Cso86] Sandor Csorgo. Testing for normality in arbitrary dimension. Annals of Statistics, 14(2):708–723, June 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349948.

Christopeit:1987:SCL

[CT87] N. Christopeit and G. Tosstorff. Strong consistency of least-squares estimators in the monotone regression model with stochastic regressors. Annals of Statistics, 15(2):568–586, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350361.

Currie:1981:DDR

[Cur81] Iain D. Currie. On distributions determined by random variables distributed over the n-cube. Annals of Statistics, 9(4):822–833, July 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345522.

Cuzick:1985:APC

[Cuz85] Jack Cuzick. Asymptotic properties of censored linear rank tests. Annals of Statistics, 13(1):133–141, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346581.

Cuzick:1988:RR

[Cuz88] Jack Cuzick. Rank regression. Annals of Statistics, 16(4): 1369–1389, December 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176351044. See correction [Cuz90].

Cuzick:1990:CRR

[Cuz90] Jack Cuzick. Correction: “Rank Regression”. Annals of Statistics, 18(1):469, March 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347513. See [Cuz88].

Callaert:1981:ONA

[CV81] Herman Callaert and Noel Veraverbeke. The order of the normal approximation for a studentized U-statistic. Annals of Statistics, 9(1):194–200, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345347.

Chang:1980:CDO

[CW80a] Der-Shin Chang and Chi Song Wong. Correction to “Design of Optimal Control for a Regression Problem”. Annals of Statistics, 8(6):1402, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345216. See [sC79].

Cheng:1980:BRM

[CW80b] Ching-Shui Cheng and Chien-Fu Wu. Balanced repeated measurements designs. Annals of Statistics, 8(6):1272–1283, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.

aos/1176345200. See corrections [CW83].

Cheng:1983:CBR

[CW83] Ching-Shui Cheng and Chien-Fu Wu. Corrections: “Balanced Repeated Measurements Designs”. Annals of Statistics, 11(1):349, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346090. See [CW80b].

Cheng:1985:DPP

[CW85a] Ping Cheng and C. F. J. Wu. Discussion: Projection pursuit. Annals of Statistics, 13(2):490–493, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349523.

Collins:1985:MVM

[CW85b] John R. Collins and Douglas P. Wiens. Minimax variance Mestimators in ε-contamination models. Annals of Statistics, 13(3): 1078–1096, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349657.

Chan:1987:AIN

[CW87] N. H. Chan and C. Z. Wei. Asymptotic inference for nearly nonstationary AR(1) processes. Annals of Statistics, 15(3):1050–1063, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350492.

Chan:1988:LDL

[CW88a] N. H. Chan and C. Z. Wei. Limiting distributions of least squares estimates of unstable autoregressive processes. Annals of Statistics, 16(1):367–401, March 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350711.

Clayton:1988:TSB

[CW88b] Murray K. Clayton and Jeffrey A. Witmer. Two-stage bandits. Annals of Statistics, 16(2):887–894, June 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350841.

Collins:1989:MPE

[CW89] John Collins and Douglas Wiens. Minimax properties of M-, R- and L-estimators of location in Lévy neighbourhoods. Annals of Statistics, 17(1):327–336, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347020.

Chow:1981:PSP

[CY81] Y. S. Chow and K. F. Yu. The performance of a sequential procedure for the estimation of the mean. Annals of Statistics, 9 (1):184–189, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345345.

Chang:1987:SCN

[CY87] Myron N. Chang and Grace L. Yang. Strong consistency of a nonparametric estimator of the survival function with doubly censored data. Annals of Statistics, 15(4):1536–1547, December 1987. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350608.

Dabrowska:1986:RTI

[Dab86] Dorota M. Dabrowska. Rank tests for independence for bivariate censored data. Annals of Statistics, 14(1):250–264, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349853.

Dabrowska:1988:KME

[Dab88] Dorota M. Dabrowska. Kaplan–Meier estimate on the plane. Annals of Statistics, 16(4):1475–1489, December 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176351049.

Dabrowska:1989:UCK

[Dab89] Dorota M. Dabrowska. Uniform consistency of the kernel conditional Kaplan–Meier estimate. Annals of Statistics, 17(3):1157– 1167, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347261.

Dahlhaus:1988:SSE

[Dah88] Rainer Dahlhaus. Small sample effects in time series analysis: A new asymptotic theory and a new estimate. Annals of Statistics, 16(2):808–841, June 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350838.

Dahlhaus:1989:EPE

[Dah89] Rainer Dahlhaus. Efficient parameter estimation for self-similar processes. Annals of Statistics, 17(4):1749–1766, December 1989.

CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347393.

Daniels:1982:STC

[Dan82] H. E. Daniels. Sequential tests constructed from images. Annals of Statistics, 10(2):394–400, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345781.

Darling:1983:ADW

[Dar83] Donald A. Darling. On the asymptotic distribution of Watson’s statistic. Annals of Statistics, 11(4):1263–1266, December 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346340.

DasGupta:1989:GTD

[Das89] Anirban DasGupta. A general theorem on decision theory for nonnegative functionals: with applications. Annals of Statistics, 17(3): 1360–1374, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347275.

Davis:1984:CPA

[Dav84] Linda Davis. Comments on a paper by T. Amemiya on estimation in a dichotomous logit regression model. Annals of Statistics, 12(2):778–782, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346525.

Davis:1985:CAN

[Dav85a] Linda June Davis. Consistency and asymptotic normality of the minimum logit chi-squared estimator when the number of design points is large. Annals of Statistics, 13(3):947–957, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349648.

Davis:1985:CCP

[Dav85b] Linda June Davis. Correction: Comments on a paper by T. Amemiya on estimation in a dichotomous logit regression model. Annals of Statistics, 13(4):1629, December 1985. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349760. See

[Ame80].

Davies:1987:ABE

[Dav87] P. L. Davies. Asymptotic behaviour of S-estimates of multivariate location parameters and dispersion matrices. Annals of Statistics, 15(3):1269–1292, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350505.

Davis:1989:IUT

[Dav89] Linda June Davis. Intersection union tests for strict collapsibility in three-dimensional contingency tables. Annals of Statistics, 17 (4):1693–1708, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347388.

Dawid:1980:CIS

[Daw80] A. Philip Dawid. Conditional independence for statistical operations. Annals of Statistics, 8(3):598–617, May 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345011.

Dawid:1985:CBE

[Daw85a] A. P. Dawid. Calibration-based empirical probability. Annals of Statistics, 13(4):1251–1274, December 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349736.

Dawid:1985:RCB

[Daw85b] A. P. Dawid. Rejoinder: Calibration-based empirical probability. Annals of Statistics, 13(4):1282–1285, December 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349738.

Dawid:1986:DCB

[Daw86] A. P. Dawid. Discussion: On the consistency of Bayes estimates. Annals of Statistics, 14(1):40–41, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349834.

DeJongh:1987:DTM

[DD87] P. J. De Jongh and T. De Wet. Discussion: The trimmed mean in the linear model. Annals of Statistics, 15(1):36–39, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350249.

Dekkers:1989:EEV

[DD89] Arnold L. M. Dekkers and Laurens De Haan. On the estimation of the extreme-value index and large quantile estimation. Annals of Statistics, 17(4):1795–1832, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347396.

Diaconis:1985:RTI

[DE85a] P. Diaconis and B. Efron. Rejoinder: Testing for independence in a two-way table: New interpretations of the chi-square statistic. Annals of Statistics, 13(3):905–913, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349645.

Diaconis:1985:TIT

[DE85b] Persi Diaconis and Bradley Efron. Testing for independence in a two-way table: New interpretations of the chi-square statistic. Annals of Statistics, 13(3):845–874, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349634.

Dekkers:1989:MEI

[DED89] A. L. M. Dekkers, J. H. J. Einmahl, and L. De Haan. A moment estimator for the index of an extreme-value distribution. Annals of Statistics, 17(4):1833–1855, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347397.

Devroye:1981:AEC

[Dev81a] Luc Devroye. On the almost everywhere convergence of nonparametric regression function estimates. Annals of Statistics, 9(6): 1310–1319, November 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345647.

Devroye:1981:APE

[Dev81b] Luc Devroye. On the asymptotic probability of error in nonparametric discrimination. Annals of Statistics, 9(6):1320–1327, November 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345648.

Devroye:1983:EWS

[Dev83] Luc Devroye. The equivalence of weak, strong and complete convergence in L_{1} for kernel density estimates. Annals of Statistics, 11 (3):896–904, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346255.

Devroye:1985:NCV

[Dev85] Luc Devroye. A note on the L_{1} consistency of variable kernel estimates. Annals of Statistics, 13(3):1041–1049, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349655.

Devroye:1987:AES

[Dev87] Luc Devroye. An application of the Efron–Stein inequality in density estimation. Annals of Statistics, 15(3):1317–1320, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350508.

Devroye:1988:APB

[Dev88] Luc Devroye. Asymptotic performance bounds for the kernel estimate. Annals of Statistics, 16(3):1162–1179, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350953.

Dalal:1982:ICS

[DF82] Siddhartha R. Dalal and Peter Fortini. An inequality comparing sums and maxima with application to Behrens–Fisher type problem. Annals of Statistics, 10(1):297–301, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345712.

Diaconis:1984:AGP

[DF84] Persi Diaconis and David Freedman. Asymptotics of graphical projection pursuit. Annals of Statistics, 12(3):793–815, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346703.

Diaconis:1986:IBE

[DF86a] P. Diaconis and D. Freedman. On inconsistent Bayes estimates of location. Annals of Statistics, 14(1):68–87, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349843.

Diaconis:1986:RCB

[DF86b] P. Diaconis and D. Freedman. Rejoinder: On the consistency of Bayes estimates. Annals of Statistics, 14(1):63–67, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349842.

Diaconis:1986:CBE

[DF86c] Persi Diaconis and David Freedman. On the consistency of Bayes estimates. Annals of Statistics, 14(1):1–26, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349830.

Doss:1989:EJS

[DFP89] Hani Doss, Steven Freitag, and Frank Proschan. Estimating jointly system and component reliabilities using a mutual censorship approach. Annals of Statistics, 17(2):764–782, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347141.

DeGroot:1980:ECC

[DG80] Morris H. DeGroot and Prem K. Goel. Estimation of the correlation coefficient from a broken random sample. Annals of Statistics, 8(2):264–278, March 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176344952.

Diaconis:1981:ASE

[DG81] Persi Diaconis and Ronald Graham. The analysis of sequential experiments with feedback to subjects. Annals of Statistics, 9 (1):3–23, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345329.

Dempster:1981:NTR

[DGG81] A. P. Dempster and M. Gasko-Green. New tools for residual analysis. Annals of Statistics, 9(5):945–959, September 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345575.

Dalal:1980:APB

[DH80] S. R. Dalal and Gaineford J. Hall. On approximating parametric Bayes models by nonparametric Bayes models. Annals of Statistics, 8(3):664–672, May 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345016.

DeRoberts:1981:BIU

[DH81] Lorraine DeRoberts and J. A. Hartigan. Bayesian inference using intervals of measures. Annals of Statistics, 9(2):235–244, March 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345391.

Diaconis:1985:DPP

[Dia85] Persi Diaconis. Discussion: Projection pursuit. Annals of Statistics, 13(2):494–496, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349525.

Diaconis:1987:DWA

[Dia87] Persi Diaconis. Discussion: What is an analysis of variance? Annals of Statistics, 15(3):917–921, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350476.

Diaconis:1989:GSA

[Dia89] Persi Diaconis. A generalization of spectral analysis with application to ranked data. Annals of Statistics, 17(3):949–979, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347251.

Dickey:1985:DTI

[Dic85] James M. Dickey. Discussion: Testing for independence in a twoway table: New interpretations of the chi-square statistic. Annals of Statistics, 13(3):877–882, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349636.

Dionne:1981:ENE

[Dio81] Louise Dionne. Efficient nonparametric estimators of parameters in the general linear hypothesis. Annals of Statistics, 9 (2):457–460, March 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345414.

Donoho:1989:PBA

[DJ89] David L. Donoho and Iain M. Johnstone. Projection-based approximation and a duality with kernel methods. Annals of Statistics, 17(1):58–106, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347004.

Donoho:1985:DPP

[DJRS85] David Donoho, Iain Johnstone, Peter Rousseeuw, and Werner Stahel. Discussion: Projection pursuit. Annals of Statistics, 13(2):496–500, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349526.

Dahiya:1980:MLE

[DK80] Ram C. Dahiya and Ramesh M. Korwar. Maximum likelihood estimates for a bivariate normal distribution with missing data. Annals of Statistics, 8(3):687–692, May 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345020.

Dykstra:1981:BNA

[DL81] R. L. Dykstra and Purushottam Laud. A Bayesian nonparametric approach to reliability. Annals of Statistics, 9(2):356–367, March 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345401.

Doksum:1986:DCB

[DL86] Kjell A. Doksum and Albert Y. Lo. Discussion: On the consistency of Bayes estimates. Annals of Statistics, 14(1):42–45, March 1986.

CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349835.

Donoho:1988:ARM

[DL88a] David L. Donoho and Richard C. Liu. The “automatic” robustness of minimum distance functionals. Annals of Statistics, 16(2):552–586, June 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350820.

Donoho:1988:PSM

[DL88b] David L. Donoho and Richard C. Liu. Pathologies of some minimum distance estimators. Annals of Statistics, 16(2):587–608, June 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350821.

Darroch:1980:MFL

[DLS80] J. N. Darroch, S. L. Lauritzen, and T. P. Speed. Markov fields and log-linear interaction models for contingency tables. Annals of Statistics, 8(3):522–539, May 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345006.

Dynkin:1983:SSP

[DM83] E. B. Dynkin and A. Mandelbaum. Symmetric statistics, Poisson point processes, and multiple Wiener integrals. Annals of Statistics, 11(3):739–745, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346241.

Does:1983:EES

[Doe83] Ronald J. M. M. Does. An Edgeworth expansion for simple linear rank statistics under the null-hypothesis. Annals of Statistics, 11(2):607–624, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346166.

Doksum:1987:EPL

[Dok87] Kjell A. Doksum. An extension of partial likelihood methods for proportional hazard models to general transformation models. Annals of Statistics, 15(1):325–345, March 1987. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350269.

Donoho:1988:OSI

[Don88] David L. Donoho. One-sided inference about functionals of a density. Annals of Statistics, 16(4):1390–1420, December 1988. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176351045.

Doss:1985:BNEa

[Dos85a] Hani Doss. Bayesian nonparametric estimation of the median; part i: Computation of the estimates. Annals of Statistics, 13(4): 1432–1444, December 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349746.

Doss:1985:BNEb

[Dos85b] Hani Doss. Bayesian nonparametric estimation of the median; Part II: Asymptotic properties of the estimates. Annals of Statistics, 13 (4):1445–1464, December 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349747.

Doss:1986:DCB

[Dos86] Hani Doss. Discussion: On the consistency of Bayes estimates. Annals of Statistics, 14(1):45–47, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349836.

Doss:1989:EDB

[Dos89] Hani Doss. On estimating the dependence between two point processes. Annals of Statistics, 17(2):749–763, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347140.

Devroye:1984:CAK

[DP84a] Luc Devroye and Clark S. Penrod. The consistency of automatic kernel density estimates. Annals of Statistics, 12(4):1231–1249, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346789.

Devroye:1984:DFL

[DP84b] Luc Devroye and Clark S. Penrod. Distribution-free lower bounds in density estimation. Annals of Statistics, 12(4):1250–1262, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346790.

Dykstra:1982:AIR

[DR82a] Richard L. Dykstra and Tim Robertson. An algorithm for isotonic regression for two or more independent variables. Annals of Statistics, 10(3):708–716, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345866.

Dykstra:1982:ORS

[DR82b] Richard L. Dykstra and Tim Robertson. Order restricted statistical tests on multinomial and Poisson parameters: The starshaped restriction. Annals of Statistics, 10(4):1246–1252, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345989. See

correction [DR87].

Davis:1984:TEM

[DR84] Richard Davis and Sidney Resnick. Tail estimates motivated by extreme value theory. Annals of Statistics, 12(4):1467–1487, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346804.

Davis:1986:LTS

[DR86] Richard Davis and Sidney Resnick. Limit theory for the sample covariance and correlation functions of moving averages. Annals of Statistics, 14(2):533–558, June 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349937.

Dykstra:1987:COR

[DR87] Richard L. Dykstra and Tim Robertson. Correction: “Order Restricted Statistical Tests on Multinomial and Poisson Parameters: The Starshaped Restriction”. Annals of Statistics, 15(1):469, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350283. See [DR82b].

DiCiccio:1988:DTC

[DR88] Thomas J. DiCiccio and Joseph P. Romano. Discussion: Theoretical comparison of bootstrap confidence intervals. Annals of Statistics, 16(3):965–969, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350938.

Drost:1989:GCS

[Dro89] F. C. Drost. Generalized chi-square goodness-of-fit tests for location-scale models when the number of classes tends to infinity. Annals of Statistics, 17(3):1285–1300, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347269.

Dawid:1982:FMB

[DS82a] A. P. Dawid and M. Stone. The functional-model basis of fiducial inference. Annals of Statistics, 10(4):1054–1067, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345970. See

reply [DS82b].

Dawid:1982:RFM

[DS82b] Professor Dawid and Professor Stone. Reply: The functionalmodel basis of fiducial inference. Annals of Statistics, 10(4):1074, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345973. See [DS82a].

Doss:1982:TPC

[DS82c] Hani Doss and Thomas Sellke. The tails of probabilities chosen from a Dirichlet prior. Annals of Statistics, 10(4):1302–1305, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345996.

Darroch:1983:AMM

[DS83] J. N. Darroch and T. P. Speed. Additive and multiplicative models and interactions. Annals of Statistics, 11(3):724–738, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346240.

Dey:1985:ECM

[DS85] Dipak K. Dey and C. Srinivasan. Estimation of a covariance matrix under Stein’s loss. Annals of Statistics, 13(4):1581–1591, December 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349756.

Das:1986:DMO

[DS86] Rita Das and Bimal K. Sinha. Detection of multivariate outliers with dispersion slippage in elliptically symmetric distributions. Annals of Statistics, 14(4):1619–1624, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350183.

Doss:1989:PBR

[DS89] Hani Doss and Jayaram Sethuraman. The price of bias reduction when there is no unbiased estimate. Annals of Statistics, 17(1):440–442, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347028.

Dehling:1989:EPS

[DT89] Herold Dehling and Murad S. Taqqu. The empirical process of some long-range dependent sequences with an application to Ustatistics. Annals of Statistics, 17(4):1767–1783, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347394.

DuMouchel:1983:ESI

[DuM83] William H. DuMouchel. Estimating the stable index α in order to measure tail thickness: A critique. Annals of Statistics, 11(4): 1019–1031, December 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346318.

Devroye:1980:DFC

[DW80] Luc P. Devroye and T. J. Wagner. Distribution-free consistency results in nonparametric discrimination and regression function estimation. Annals of Statistics, 8(2):231–239, March 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344949.

Dupacova:1988:ABS

[DW88] Jitka Dupacova and Roger Wets. Asymptotic behavior of statistical estimators and of optimal solutions of stochastic optimization problems. Annals of Statistics, 16(4):1517–1549, December 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176351052.

deWet:1987:ESP

[dWR87] Tertius de Wet and Ronald H. Randles. On the effect of substituting parameter estimators in limiting χ^{2}U and V statistics. Annals of Statistics, 15(1):398–412, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350274.

Eaton:1981:PID

[Eat81] Morris L. Eaton. On the projections of isotropic distributions. Annals of Statistics, 9(2):391–400, March 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345404.

Eaton:1982:RST

[Eat82] Morris L. Eaton. A review of selected topics in multivariate probability inequalities. Annals of Statistics, 10(1):11–43, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345688.

Eberl:1983:ISE

[Ebe83] W. Eberl. Invariantly sufficient equivariant statistics and characterizations of normality in translation classes. Annals of Statistics, 11(1):330–336, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346084.

Eddy:1980:OKE

[Edd80] William F. Eddy. Optimum kernel estimators of the mode. Annals of Statistics, 8(4):870–882, July 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345080.

Edelman:1988:EMD

[Ede88] David Edelman. Estimation of the mixing distribution for a normal mean with applications to the compound decision problem.

Annals of Statistics, 16(4):1609–1622, December 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176351056.

Edwards:1987:EPS

[Edw87] Don Edwards. Extended-Paulson sequential selection. Annals of Statistics, 15(1):449–455, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350280.

Efron:1982:CTT

[Efr82a] Bradley Efron. Correction: “Transformation Theory: How Normal is a Family of Distributions”. Annals of Statistics, 10(3):1032, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345897. See [Efr82c].

Efron:1982:MLD

[Efr82b] Bradley Efron. Maximum likelihood and decision theory. Annals of Statistics, 10(2):340–356, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345778.

Efron:1982:TTH

[Efr82c] Bradley Efron. Transformation theory: How normal is a family of distributions? Annals of Statistics, 10(2):323–339, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345777. See

correction [Efr82a].

Efron:1986:DJB

[Efr86] B. Efron. Discussion: Jackknife, bootstrap and other resampling methods in regression analysis. Annals of Statistics, 14(4): 1301–1304, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350145.

Efron:1988:DTC

[Efr88] B. Efron. Discussion: Theoretical comparison of bootstrap confidence intervals. Annals of Statistics, 16(3):969–972, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350939.

Eisenberg:1980:CUM

[EG80] Bennett Eisenberg and B. K. Ghosh. Curtailed and uniformly most powerful sequential tests. Annals of Statistics, 8(5):1123– 1131, September 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345149.

Eguchi:1983:SOE

[Egu83] Shinto Eguchi. Second order efficiency of minimum contrast estimators in a curved exponential family. Annals of Statistics, 11(3): 793–803, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346246.

Eick:1988:TAB

[Eic88] Stephen G. Eick. The two-armed bandit with delayed responses. Annals of Statistics, 16(1):254–264, March 1988. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350703.

Eaton:1983:MTI

[EK83] Morris Eaton and Takeaki Kariya. Multivariate tests with incomplete data. Annals of Statistics, 11(2):654–665, June 1983. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346170.

Eaton:1987:BEE

[EO87] Morris L. Eaton and Ingram Olkin. Best equivariant estimators of a Cholesky decomposition. Annals of Statistics, 15(4): 1639–1650, December 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350615.

Eplett:1981:ILR

[Epl81] W. J. R. Eplett. The inadmissibility of linear rank tests under Bahadur efficiency. Annals of Statistics, 9(5):1079–1086, September 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345587.

Eplett:1982:RTG

[Epl82] W. J. R. Eplett. Rank tests generated by continuous piecewise linear functions. Annals of Statistics, 10(2):569–574, June 1982.

CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345797.

Epps:1987:TST

[Epp87] T. W. Epps. Testing that a stationary time series is Gaussian. Annals of Statistics, 15(4):1683–1698, December 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350618.

Epps:1988:TGP

[Epp88] T. W. Epps. Testing that a Gaussian process is stationary. Annals of Statistics, 16(4):1667–1683, December 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176351060.

Eriksen:1987:PCM

[Eri87] P. Svante Eriksen. Proportionality of covariance matrices. Annals of Statistics, 15(2):732–748, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350372.

Efron:1981:JEV

[ES81a] B. Efron and C. Stein. The jackknife estimate of variance. Annals of Statistics, 9(3):586–596, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345462.

Evans:1981:CLD

[ES81b] G. B. A. Evans and N. E. Savin. The calculation of the limiting distribution of the least squares estimator of the parameter in a random walk model. Annals of Statistics, 9(5):1114–1118, September 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345591.

Eubank:1989:DLS

[ES89] R. L. Eubank and P. Speckman. Discussion: Linear smoothers and additive models. Annals of Statistics, 17(2):525–529, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347119.

Eubank:1981:UEU

[ESS81] R. L. Eubank, Patricia L. Smith, and Philip W. Smith. Uniqueness and eventual uniqueness of optimal designs in some time series models. Annals of Statistics, 9(3):486–493, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345453.

Eubank:1982:NOA

[ESS82] R. L. Eubank, P. L. Smith, and P. W. Smith. A note on optimal and asymptotically optimal designs for certain time series models. Annals of Statistics, 10(4):1295–1301, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345995.

Esty:1982:CIC

[Est82] Warren W. Esty. Confidence intervals for the coverage of low coverage samples. Annals of Statistics, 10(1):190–196, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345701.

Esty:1983:NLL

[Est83] Warren W. Esty. A normal limit law for a nonparametric estimator of the coverage of a random sample. Annals of Statistics, 11(3): 905–912, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346256.

Esty:1986:EGN

[Est86] Warren W. Esty. The efficiency of Good’s nonparametric coverage estimator. Annals of Statistics, 14(3):1257–1260, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350066.

Eubank:1981:DQF

[Eub81] R. L. Eubank. A density-quantile function approach to optimal spacing selection. Annals of Statistics, 9(3):494–500, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345454.

Evans:1983:EE

[Eva83] Michael Evans. Estimating events. Annals of Statistics, 11(4): 1218–1224, December 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346334.

Feigin:1986:IDC

[FA86] Paul D. Feigin and Mayer Alvo. Intergroup diversity and concordance for ranking data: An approach via metrics for permutations. Annals of Statistics, 14(2):691–707, June 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349947.

Fabius:1973:TCD

[Fab73] J. Fabius. Two characterizations of the Dirichlet distribution. Annals of Statistics, 1(3):583–587, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:/

/projecteuclid.org/euclid.aos/1176342429. See correction

[Fab81].

Fabius:1981:CTC

[Fab81] J. Fabius. Correction to “Two Characterizations of the Dirichlet Distribution”. Annals of Statistics, 9(1):234, January 1981. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345356. See

[Fab73].

Fabian:1988:PER

[Fab88] Vaclav Fabian. Polynomial estimation of regression functions with the supremum norm error. Annals of Statistics, 16(4):1345–1368, December 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176351043.

Falk:1984:RDK

[Fal84] Michael Falk. Relative deficiency of kernel type estimators of quantiles. Annals of Statistics, 12(1):261–268, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346405.

Falk:1985:ANK

[Fal85] Michael Falk. Asymptotic normality of the kernel quantile estimator. Annals of Statistics, 13(1):428–433, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346605.

Freedman:1982:IE

[FD82a] D. A. Freedman and P. Diaconis. On inconsistent M-estimators. Annals of Statistics, 10(2):454–461, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345786.

Freedman:1982:FTS

[FD82b] David Freedman and Persi Diaconis. De Finetti’s theorem for symmetric location families. Annals of Statistics, 10(1):184–189, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345700.

Freedman:1983:IBE

[FD83] David Freedman and Persi Diaconis. On inconsistent Bayes estimates in the discrete case. Annals of Statistics, 11(4):1109–1118, December 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346325.

Fountis:1989:TUR

[FD89] Nicolaos G. Fountis and David A. Dickey. Testing for a unit root nonstationarity in multivariate autoregressive time series. Annals of Statistics, 17(1):419–428, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347025.

Feigin:1981:CEF

[Fei81] Paul D. Feigin. Conditional exponential families and a representation theorem for asympotic inference. Annals of Statistics, 9(3):597–603, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345463.

Felsenstein:1986:DJB

[Fel86] Joseph Felsenstein. Discussion: Jackknife, bootstrap and other resampling methods in regression analysis. Annals of Statistics, 14 (4):1304–1305, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350146.

Ferebee:1982:TPB

[Fer82a] Brooks Ferebee. Tests with parabolic boundary for the drift of a Wiener process. Annals of Statistics, 10(3):882–894, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345878.

Ferreira:1982:SET

[Fer82b] Pedro E. Ferreira. Sequential estimation through estimating equations in the nuisance parameter case. Annals of Statistics, 10 (1):167–173, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345698.

Friedman:1988:BIE

[FGP88] Yaakov Friedman, Alexander Gelman, and Eswar Phadia. Best invariant estimation of a distribution function under the Kolmogorov–Smirnov loss function. Annals of Statistics, 16(3): 1254–1261, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350959.

Franke:1986:DIF

[FH86] J. Franke and E. J. Hannan. Discussion: Influence functionals for time series. Annals of Statistics, 14(3):822–824, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350029.

Fuller:1981:EPS

[FHG81] Wayne A. Fuller, David P. Hasza, and J. Jeffery Goebel. Estimation of the parameters of stochastic difference equations. Annals of Statistics, 9(3):531–543, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345457.

Felzenbaum:1983:ISM

[FHH83] Alexander Felzenbaum, Sergiu Hart, and Yosef Hochberg. Improving some multiple comparison procedures. Annals of Statistics, 11(1):121–128, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346063.

Fraser:1981:IFM

[FHW81] Martin D. Fraser, Yu-Sheng Hsu, and Joseph J. Walker. Identifiability of finite mixtures of von Mises distributions. Annals of Statistics, 9(5):1130–1131, September 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345595.

Field:1982:SSA

[Fie82] Christopher Field. Small sample asymptotic expansions for multivariate M-estimates. Annals of Statistics, 10(3):672–689, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345864.

Fienberg:1985:DTI

[Fie85] Stephen E. Fienberg. Discussion: Testing for independence in a two-way table: New interpretations of the chi-square statistic. Annals of Statistics, 13(3):882–884, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349637.

Finster:1983:OSS

[Fin83] Mark Finster. Optimal stopping in the stock market when the future is discounted. Annals of Statistics, 11(2):564–568, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346161.

Finster:1985:EGL

[Fin85] Mark Finster. Estimation in the general linear model when the accuracy is specified before data collection. Annals of Statistics, 13(2):663–675, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349546.

Fisher:1982:UES

[Fis82] N. I. Fisher. Unbiased estimation for some non-parametric families of distributions. Annals of Statistics, 10(2):603–615, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345801.

Fishburn:1983:ERN

[Fis83] Peter C. Fishburn. Ellsberg revisited: A new look at comparative probability. Annals of Statistics, 11(4):1047–1059, December 1983.

CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346320.

Fill:1984:PPM

[FJ84] James Allen Fill and Iain Johnstone. On projection pursuit measures of multivariate location and dispersion. Annals of Statistics, 12(1):127–141, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346397.

Falk:1984:RST

[FK84] M. Falk and W. Kohne. A robustification of the sign test under mixing conditions. Annals of Statistics, 12(2):716–729, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346517.

Fahrmeir:1985:CAN

[FK85] Ludwig Fahrmeir and Heinz Kaufmann. Consistency and asymptotic normality of the maximum likelihood estimator in generalized linear models. Annals of Statistics, 13(1):342–368, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346597. See

correction [FK86].

Fahrmeir:1986:CCA

[FK86] Ludwig Fahrmeir and Heinz Kaufmann. Correction: “Consistency and Asymptotic Normality of the Maximum Likelihood Estimator in Generalized Linear Models”. Annals of Statistics, 14 (4):1643, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350187. See [FK85].

Farrell:1989:AAL

[FKZ89] Roger H. Farrell, Witold Klonecki, and Stefan Zontek. All admissible linear estimators of the vector of gamma scale parameters with application to random effects models. Annals of Statistics, 17(1):268–281, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347015.

Freedman:1980:EDF

[Fl80] David Freedman and David lane. The empirical distribution of Fourier coefficients. Annals of Statistics, 8(6):1244–1251, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345197.

Fishburn:1987:NNA

[FL87] Peter C. Fishburn and Irving H. LaValle. A nonlinear, nontransitive and additive-probability model for decisions under uncertainty. Annals of Statistics, 15(2):830–844, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350378.

Flury:1986:ATC

[Flu86] Bernard N. Flury. Asymptotic theory for common principal component analysis. Annals of Statistics, 14(2):418–430, June 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349930.

Foutz:1980:TGF

[Fou80] Robert V. Foutz. A test for goodness-of-fit based on an empirical probability measure. Annals of Statistics, 8(5):989–1001, September 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345137.

Fox:1981:IBI

[Fox81] Martin Fox. An inadmissible best invariant estimator: The I.I.D. case. Annals of Statistics, 9(5):1127–1129, September 1981. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345594.

Fairley:1987:PAM

[FPV87] David Fairley, Dennis K. Pearl, and Joseph S. Verducci. The penalty for assuming that a monotone regression is linear. Annals of Statistics, 15(1):443–448, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350279.

Foldes:1981:SUC

[FR81] Antonia Foldes and Lidia Rejto. Strong uniform consistency for nonparametric survival curve estimators from randomly censored data. Annals of Statistics, 9(1):122–129, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345337.

Friedman:1983:GTM

[FR83] Jerome H. Friedman and Lawrence C. Rafsky. Graph-theoretic measures of multivariate association and prediction. Annals of Statistics, 11(2):377–391, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346148.

Frees:1985:SNA

[FR85] Edward W. Frees and David Ruppert. Sequential nonparametric age replacement policies. Annals of Statistics, 13(2):650–662, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349545.

Fraser:1982:DFM

[Fra82] D. A. S. Fraser. Discussion: The functional-model basis of fiducial inference. Annals of Statistics, 10(4):1070–1073, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345972.

Freedman:1981:BRM

[Fre81] D. A. Freedman. Bootstrapping regression models. Annals of Statistics, 9(6):1218–1228, November 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345638.

Freedman:1984:BTS

[Fre84] D. Freedman. On bootstrapping two-stage least-squares estimates in stationary linear models. Annals of Statistics, 12(3): 827–842, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346705.

Freedman:1986:DJB

[Fre86a] D. A. Freedman. Discussion: Jackknife, bootstrap and other resampling methods in regression analysis. Annals of Statistics, 14 (4):1305–1308, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350147.

Frees:1986:NRF

[Fre86b] Edward W. Frees. Nonparametric renewal function estimation. Annals of Statistics, 14(4):1366–1378, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350163. See cor-

rection [Fre88].

Frees:1988:CNR

[Fre88] Edward W. Frees. Correction: “Nonparametric Renewal Function Estimation”. Annals of Statistics, 16(4):1741, December 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176351066. See

[Fre86b].

Friedman:1982:PEM

[Fri82] Michael Friedman. Piecewise exponential models for survival data with covariates. Annals of Statistics, 10(1):101–113, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345693.

Friedman:1985:DPP

[Fri85] Jerome H. Friedman. Discussion: Projection pursuit. Annals of Statistics, 13(2):475–481, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349520.

Friedrich:1989:BEB

[Fri89] Karl O. Friedrich. A Berry–Esseen bound for functions of independent random variables. Annals of Statistics, 17(1):170–183, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347009.

Fujikoshi:1989:AES

[FS89] Yasunori Fujikoshi and Ryoichi Shimizu. Asymptotic expansions of some mixtures of the multivariate normal distribution and their error bounds. Annals of Statistics, 17(3):1124–1132, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347259.

Fox:1986:LSP

[FT86] Robert Fox and Murad S. Taqqu. Large-sample properties of parameter estimates for strongly dependent stationary Gaussian time series. Annals of Statistics, 14(2):517–532, June 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349936.

Fu:1982:LSP

[Fu82] James C. Fu. Large sample point estimation: A large deviation theory approach. Annals of Statistics, 10(3):762–771, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345869.

Fujikoshi:1985:EBA

[Fuj85] Yasunori Fujikoshi. An error bound for an asymptotic expansion of the distribution function of an estimate in a multivariate linear model. Annals of Statistics, 13(2):827–831, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349562.

Fuller:1980:PSE

[Ful80] Wayne A. Fuller. Properties of some estimators for the errors-in-variables model. Annals of Statistics, 8(2):407–422, March 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344961.

Gaffke:1981:SCO

[Gaf81] N. Gaffke. Some classes of optimality criteria and optimal designs for complete two-way layouts. Annals of Statistics, 9(4):893–898, July 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345530.

Gaffke:1987:FCD

[Gaf87] Norbert Gaffke. Further characterizations of design optimality and admissibility for partial parameter estimation in linear regression. Annals of Statistics, 15(3):942–957, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350485.

Gajek:1986:IDE

[Gaj86] Leslaw Gajek. On improving density estimators which are not bona fide functions. Annals of Statistics, 14(4):1612–1618, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350182.

Gajek:1988:MVS

[Gaj88] Leslaw Gajek. On the minimax value in the scale model with truncated data. Annals of Statistics, 16(2):669–677, June 1988.

CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350827.

Gatsonis:1984:DPD

[Gat84] Constantine A. Gatsonis. Deriving posterior distributions for a location parameter: A decision theoretic approach. Annals of Statistics, 12(3):958–970, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346714.

Gilliland:1982:BEB

[GBT82] Dennis C. Gilliland, John E. Boyer, and How Jan Tsao. Bayes empirical Bayes: Finite parameter case. Annals of Statistics, 10(4): 1277–1282, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345993.

Good:1987:RSM

[GC87] I. J. Good and J. F. Crook. The robustness and sensitivity of the mixed-Dirichlet Bayesian test for “independence” in contingency tables. Annals of Statistics, 15(2):670–693, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350368.

Gleser:1987:LDL

[GCG87] Leon Jay Gleser, Raymond J. Carroll, and Paul P. Gallo. The limiting distribution of least squares in an errors-in-variables regression model. Annals of Statistics, 15(1):220–233, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350262.

Genest:1984:CTE

[Gen84] Christian Genest. A characterization theorem for externally Bayesian groups. Annals of Statistics, 12(3):1100–1105, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346726.

George:1986:MMS

[Geo86] Edward I. George. Minimax multiple shrinkage estimation. Annals of Statistics, 14(1):188–205, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349849.

Gander:1989:DLS

[GG89] Walter Gander and Gene H. Golub. Discussion: Linear smoothers and additive models. Annals of Statistics, 17(2):529–532, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176347120;   http://www.jstor.org/stable/2241565.

Gupta:1980:NOS

[GH80] Shanti S. Gupta and Deng-Yuan Huang. A note on optimal subset selection procedures. Annals of Statistics, 8(5):1164–1167, September 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345154.

Geman:1982:NML

[GH82] Stuart Geman and Chii-Ruey Hwang. Nonparametric maximum likelihood estimation by the method of sieves. Annals of Statistics, 10(2):401–414, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345782.

Gleser:1987:NCS

[GH87] Leon Jay Gleser and Jiunn T. Hwang. The nonexistence of 100(1 − α)% confidence sets of finite expected diameter in errorsin-variables and related models. Annals of Statistics, 15(4): 1351–1362, December 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350597.

Ghosh:1980:MEP

[Gho80] Subir Ghosh. On main effect plus one plans for 2^{m} factorials. Annals of Statistics, 8(4):922–930, July 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:/

/projecteuclid.org/euclid.aos/1176345084. See correction

[Gho91].

Ghosh:1986:DJB

[Gho86] Malay Ghosh. Discussion: Jackknife, bootstrap and other resampling methods in regression analysis. Annals of Statistics, 14(4): 1308–1310, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350148.

Ghosh:1991:CME

[Gho91] Subir Ghosh. Correction: “On Main Effect plus One Plans for 2^{m} Factorials”. Annals of Statistics, 19(4):2281, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176348402. See

[Gho80].

Guess:1986:TEV

[GHP86] Frank Guess, Myles Hollander, and Frank Proschan. Testing exponentiality versus a trend change in mean residual life. Annals of Statistics, 14(4):1388–1398, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350165.

Ghosh:1983:CIE

[GHT83] Malay Ghosh, Jiunn Tzon Hwang, and Kam-Wah Tsui. Construction of improved estimators in multiparameter estimation for discrete exponential families. Annals of Statistics, 11(2):351–367, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.

aos/1176346143. See discussion [Ber83a, Hud83, Mor83a, Ano83].

Gill:1981:TRP

[Gil81] R. D. Gill. Testing with replacement and the product limit estimator. Annals of Statistics, 9(4):853–860, July 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345525.

Gill:1983:LSB

[Gil83a] Richard Gill. Large sample behaviour of the product-limit estimator on the whole line. Annals of Statistics, 11(1):49–58, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346055.

Gilstein:1983:JAD

[Gil83b] C. Zachary Gilstein. On the joint asymptotic distribution of extreme midranges. Annals of Statistics, 11(3):913–920, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346257.

Gill:1986:TTT

[Gil86] Richard D. Gill. The total time on test plot and the cumulative total time on test statistic for a counting process. Annals of Statistics, 14(3):1234–1239, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350062.

Galil:1980:OWD

[GK80] Z. Galil and J. Kiefer. D-optimum weighing designs. Annals of Statistics, 8(6):1293–1306, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345202.

Galil:1982:CMO

[GK82] Z. Galil and J. Kiefer. Construction methods for D-optimum weighing designs when n ≡ 3(mod4). Annals of Statistics, 10 (2):502–510, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345791.

Gnanadesikan:1985:DPP

[GK85] R. Gnanadesikan and J. R. Kettenring. Discussion: Projection pursuit. Annals of Statistics, 13(2):500–502, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349527.

Gasser:1989:DLS

[GK89] Theo Gasser and Alois Kneip. Discussion: Linear smoothers and additive models. Annals of Statistics, 17(2):532–535, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347121.

Greblicki:1984:DFP

[GKP84] Włodzimierz Greblicki, Adam Krzyzak, and Miroslaw Pawlak. Distribution-free pointwise consistency of kernel regression estimate. Annals of Statistics, 12(4):1570–1575, December 1984. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346815.

Guttorp:1988:ADQ

[GL88] Peter Guttorp and Richard A. Lockhart. On the asymptotic distribution of quadratic forms in uniform order statistics. Annals of Statistics, 16(1):433–449, March 1988. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350713.

Gleser:1981:EME

[Gle81] Leon Jay Gleser. Estimation in a multivariate “Errors in Variables” regression model: Large sample results. Annals of Statistics, 9(1):24–44, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345330.

Gleser:1986:MEN

[Gle86] Leon Jay Gleser. Minimax estimators of a normal mean vector for arbitrary quadratic loss and unknown covariance matrix. Annals of Statistics, 14(4):1625–1633, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350184.

Glynn:1980:ARD

[Gly80] William J. Glynn. Asymptotic representations of the densities of canonical correlations and latent roots in MANOVA when the population parameters have arbitrary multiplicity. Annals of Statistics, 8(5):958–976, September 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345135.

Ghosh:1980:SPE

[GM80] Malay Ghosh and Nitis Mukhopadhyay. Sequential point estimation of the difference of two normal means. Annals of Statistics, 8 (1):221–225, January 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176344906.

Gleser:1983:EDC

[GM83] Leon J. Gleser and David S. Moore. The effect of dependence on chi-squared and empiric distribution tests of fit. Annals of Statistics, 11(4):1100–1108, December 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346324.

Gupta:1984:SSP

[GM84] Shanti S. Gupta and Klaus J. Miescke. Sequential selection procedures — a decision theoretic approach. Annals of Statistics, 12 (1):336–350, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346411.

Good:1987:AGT

[GM87a] I. J. Good and Y. Mittal. The amalgamation and geometry of two-by-two contingency tables. Annals of Statistics, 15(2):694– 711, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350369.

Gutmann:1987:SPB

[GM87b] Sam Gutmann and Zakhar Maymin. Is the selected population the best? Annals of Statistics, 15(1):456–461, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350281.

Good:1989:AAG

[GM89] I. J. Good and Y. Mittal. Addendum: The amalgamation and geometry of two-by-two contingency tables. Annals of Statistics, 17(2):947, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347155.

Gasser:1984:CNR

[GMK^{+}84a] Theo Gasser, Hans-Georg Muller, Walter Kohler, Luciano Molinari, and Andrea Prader. Correction: “Nonparametric Regression Analysis of Growth Curves”. Annals of Statistics, 12(4):1588, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346818. See [GMK^{+}84b].

Gasser:1984:NRA

[GMK^{+}84b] Theo Gasser, Hans-Georg Muller, Walter Kohler, Luciano Molinari, and Andrea Prader. Nonparametric regression analysis of growth curves. Annals of Statistics, 12(1):210–229, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346402. See

correction [GMK^{+}84a].

Genest:1986:CEB

[GMS86] Christian Genest, Kevin J. McConway, and Mark J. Schervish. Characterization of externally Bayesian pooling operators. Annals of Statistics, 14(2):487–501, June 1986. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349934.

Guttman:1986:MEN

[GMT86] Irwin Guttman, U. Menzefricke, and David Tyler. Magnitudinal effects in the normal multivariate model. Annals of Statistics, 14(4): 1555–1571, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350176.

Ghosh:1981:NFM

[GMY81] J. K. Ghosh, H. Morimoto, and S. Yamada. Neyman factorization and minimality of pairwise sufficient subfields. Annals of Statistics, 9(3):514–530, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345456.

Ghosh:1987:SSE

[GNS87] Malay Ghosh, David M. Nickerson, and Pranab K. Sen. Sequential shrinkage estimation. Annals of Statistics, 15(2):817–829, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350377.

Godolphin:1980:IPM

[God80] E. J. Godolphin. An invariance property for the maximum likelihood estimator of the parameters of a Gaussian moving average process. Annals of Statistics, 8(5):1093–1099, September 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345146.

Goldstein:1981:BCS

[Gol81] Michael Goldstein. A Bayesian criterion for sample size. Annals of Statistics, 9(3):670–672, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345471.

Goldstein:1982:CD

[Gol82] Michael Goldstein. Contamination distributions. Annals of Statistics, 10(1):174–183, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345699.

Goldstein:1984:TPE

[Gol84] Michael Goldstein. Turning probabilities into expectations. Annals of Statistics, 12(4):1551–1557, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346812.

Good:1985:DTI

[Goo85a] I. J. Good. Discussion: Testing for independence in a two-way table: New interpretations of the chi-square statistic. Annals of Statistics, 13(3):885–886, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349638.

Goodman:1985:ACC

[Goo85b] Leo A. Goodman. The analysis of cross-classified data having ordered and/or unordered categories: Association models, correlation models, and asymmetry models for contingency tables with or without missing entries. Annals of Statistics, 13(1):10–69, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346576.

Goodman:1985:DTI

[Goo85c] Leo A. Goodman. Discussion: Testing for independence in a twoway table: New interpretations of the chi-square statistic. Annals of Statistics, 13(3):887–893, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349639.

Gordon:1983:SSL

[Gor83] Louis Gordon. Successive sampling in large finite populations. Annals of Statistics, 11(2):702–706, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346175.

Gordon:1989:BDG

[Gor89] Louis Gordon. Bounds for the distribution of the generalized variance. Annals of Statistics, 17(4):1684–1692, December 1989. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347387.

Gray:1985:GFT

[GP85] Robert J. Gray and Donald A. Pierce. Goodness-of-fit tests for censored survival data. Annals of Statistics, 13(2):552–563, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349538.

Ghosh:1984:NBS

[GPSB84] Malay Ghosh, William C. Parr, Kesar Singh, and G. Jogesh Babu. A note on bootstrapping the sample median. Annals of Statistics, 12(3):1130–1135, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346731.

Grambsch:1983:SSB

[Gra83] Patricia Grambsch. Sequential sampling based on the observed Fisher information to guarantee the accuracy of the maximum likelihood estimator. Annals of Statistics, 11(1):68–77, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346057.

Graybill:1987:DWA

[Gra87] Franklin A. Graybill. Discussion: What is an analysis of variance? Annals of Statistics, 15(3):921–923, September 1987. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350477.

Gray:1988:CST

[Gra88] Robert J. Gray. A class of K-sample tests for comparing the cumulative incidence of a competing risk. Annals of Statistics, 16(3): 1141–1154, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350951.

Gregory:1980:EOQ

[Gre80] Gavin G. Gregory. On efficiency and optimality of quadratic tests. Annals of Statistics, 8(1):116–131, January 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344895.

Grenander:1989:APT

[Gre89] Ulf Grenander. Advances in pattern theory. Annals of Statistics, 17(1):1–30, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347002.

Grubel:1988:LS

[Gru88] R. Grubel. The length of the shorth. Annals of Statistics, 16 (2):619–628, June 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350823.

Ghosh:1981:NSC

[GS81a] J. K. Ghosh and Bimal K. Sinha. A necessary and sufficient condition for second order admissibility with applications to Berkson’s bioassay problem. Annals of Statistics, 9(6):1334–1338, November 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345650.

Gong:1981:PML

[GS81b] Gail Gong and Francisco J. Samaniego. Pseudo maximum likelihood estimation: Theory and applications. Annals of Statistics, 9(4):861–869, July 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345526.

Genest:1985:MEJ

[GS85a] Christian Genest and Mark J. Schervish. Modeling expert judgments for Bayesian updating. Annals of Statistics, 13(3):1198– 1212, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349664.

Guttorp:1985:CEP

[GS85b] Peter Guttorp and Andrew F. Siegel. Consistent estimation in partially observed random walks. Annals of Statistics, 13(3): 958–969, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349649.

Gardiner:1986:TSE

[GSV86] Joseph C. Gardiner, V. Susarla, and John Van Ryzin. Time sequential estimation of the exponential mean under random withdrawals. Annals of Statistics, 14(2):607–618, June 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349941.

Ghosh:1980:SOE

[GSW80] J. K. Ghosh, B. K. Sinha, and H. S. Wieand. Second order efficiency of the MLE with respect to any bounded bowl-shape loss function. Annals of Statistics, 8(3):506–521, May 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345005.

Gupta:1986:SEM

[Gup86] Anirban Das Gupta. Simultaneous estimation in the multiparameter gamma distribution under weighted quadratic losses. Annals of Statistics, 14(1):206–219, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349850.

Gutmann:1982:SPI

[Gut82] Sam Gutmann. Stein’s paradox is impossible in problems with finite sample space. Annals of Statistics, 10(3):1017–1020, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345893.

Gill:1988:LST

[GVW88] Richard D. Gill, Yehuda Vardi, and Jon A. Wellner. Large sample theory of empirical distributions in biased sampling models. Annals of Statistics, 16(3):1069–1112, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350948.

Ghosh:1988:SEP

[GY88] Malay Ghosh and Ming-Chung Yang. Simultaneous estimation of Poisson means under entropy loss. Annals of Statistics, 16 (1):278–291, March 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350705.

Gine:1989:NCB

[GZ89] Evarist Gine and Joel Zinn. Necessary conditions for the bootstrap of the mean. Annals of Statistics, 17(2):684–691, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347134.

Haberman:1981:TIT

[Hab81] Shelby J. Haberman. Tests for independence in two-way contingency tables based on canonical correlation and on linear-by-linear interaction. Annals of Statistics, 9(6):1178–1186, November 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345635.

Haberman:1984:AMD

[Hab84] Shelby J. Haberman. Adjustment by minimum discriminant information. Annals of Statistics, 12(3):971–988, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346715. See

correction [Hab86].

Haberman:1986:CAM

[Hab86] Shelby J. Haberman. Correction: “Adjustment by Minimum Discriminant Information”. Annals of Statistics, 14(1):358, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349863. See [Hab84].

Haberman:1989:CE

[Hab89] Shelby J. Haberman. Concavity and estimation. Annals of Statistics, 17(4):1631–1661, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347385.

Haff:1979:EIC

[Haf79] L. R. Haff. Estimation of the inverse covariance matrix: Random mixtures of the inverse Wishart matrix and the identity. Annals of Statistics, 7(6):1264–1276, November 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344845. See cor-

rection [Haf81].

Haff:1980:EBE

[Haf80] L. R. Haff. Empirical Bayes estimation of the multivariate normal covariance matrix. Annals of Statistics, 8(3):586–597, May 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345010.

Haff:1981:CEI

[Haf81] L. R. Haff. Corrections to “Estimation of the Inverse Covariance Matrix: Random Mixtures of the Inverse Wishart Matrix and the Identity”. Annals of Statistics, 9(5):1132, September 1981. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345596. See

[Haf79].

Hall:1981:ATT

[Hal81a] Peter Hall. Asymptotic theory of triple sampling for sequential estimation of a mean. Annals of Statistics, 9(6):1229–1238, November 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345639.

Hall:1981:TSE

[Hal81b] Peter Hall. On trigonometric series estimates of densities. Annals of Statistics, 9(3):683–685, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345474.

Hall:1982:EED

[Hal82] Peter Hall. On estimating the endpoint of a distribution. Annals of Statistics, 10(2):556–568, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345796.

Hall:1983:IEE

[Hal83a] Peter Hall. Inverting an Edgeworth expansion. Annals of Statistics, 11(2):569–576, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346162.

Hall:1983:LSO

[Hal83b] Peter Hall. Large sample optimality of least squares crossvalidation in density estimation. Annals of Statistics, 11(4): 1156–1174, December 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346329.

Hall:1983:OSM

[Hal83c] Peter Hall. Orthogonal series methods for both qualitative and quantitative data. Annals of Statistics, 11(3):1004–1007, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346268.

Hall:1984:ISE

[Hal84a] Peter Hall. Integrated square error properties of kernel estimators of regression functions. Annals of Statistics, 12(1):241–260, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346404.

Hallin:1984:SFN

[Hal84b] Marc Hallin. Spectral factorization of nonstationary moving average processes. Annals of Statistics, 12(1):172–192, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346400.

Hall:1986:DJB

[Hal86a] Peter Hall. Discussion: Jackknife, bootstrap and other resampling methods in regression analysis. Annals of Statistics, 14(4): 1311–1312, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350149.

Hall:1986:BCI

[Hal86b] Peter Hall. On the bootstrap and confidence intervals. Annals of Statistics, 14(4):1431–1452, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350168.

Hall:1986:NBS

[Hal86c] Peter Hall. On the number of bootstrap simulations required to construct a confidence interval. Annals of Statistics, 14(4): 1453–1462, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350169.

Hall:1987:KLL

[Hal87] Peter Hall. On Kullback–Leibler loss and density estimation. Annals of Statistics, 15(4):1491–1519, December 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350606.

Hall:1988:RTC

[Hal88a] Peter Hall. Rejoinder: Theoretical comparison of bootstrap confidence intervals. Annals of Statistics, 16(3):981–985, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350944.

Hall:1988:TCB

[Hal88b] Peter Hall. Theoretical comparison of bootstrap confidence intervals. Annals of Statistics, 16(3):927–953, September 1988. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350933.

Hall:1989:PBP

[Hal89a] Peter Hall. On polynomial-based projection indices for exploratory projection pursuit. Annals of Statistics, 17(2):589–605, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347127.

Hall:1989:PPR

[Hal89b] Peter Hall. On projection pursuit regression. Annals of Statistics, 17(2):573–588, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347126.

Hammerstrom:1981:AOT

[Ham81] Thomas Hammerstrom. Asymptotically optimal tests for heteroscedasticity in the general linear model. Annals of Statistics, 9(2):368–380, March 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345402.

Hannan:1975:EAM

[Han75] E. J. Hannan. The estimation of ARMA models. Annals of Statistics, 3(4):975–981, July 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343200. See correction [Han81].

Hannan:1980:EOA

[Han80a] E. J. Hannan. The estimation of the order of an ARMA process. Annals of Statistics, 8(5):1071–1081, September 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345144.

Hannan:1980:REB

[Han80b] E. J. Hannan. Recursive estimation based on ARMA models. Annals of Statistics, 8(4):762–777, July 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345069.

Hannan:1981:CEA

[Han81] E. J. Hannan. Correction to “The Estimation of ARMA Models”. Annals of Statistics, 9(1):233, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345355. See [Han75].

Hartung:1981:NMB

[Har81a] Joachim Hartung. Nonnegative minimum biased invariant estimation in variance component models. Annals of Statistics, 9 (2):278–292, March 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345394.

Harville:1981:UMV

[Har81b] David A. Harville. Unbiased and minimum-variance unbiased estimation of estimable functions for fixed linear models with arbitrary covariance structure. Annals of Statistics, 9(3):633–637, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345467.

Hardle:1984:LIL

[Har84] Wolfgang Hardle. A law of the iterated logarithm for nonparametric regression function estimators. Annals of Statistics, 12 (2):624–635, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346510.

Hartigan:1986:DCB

[Har86] J. A. Hartigan. Discussion: On the consistency of Bayes estimates. Annals of Statistics, 14(1):48, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349837.

Hart:1988:ATP

[Har88] Jeffrey D. Hart. An ARMA type probability density estimator. Annals of Statistics, 16(2):842–855, June 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350839.

Hathaway:1985:CFM

[Hat85] Richard J. Hathaway. A constrained formulation of maximumlikelihood estimation for normal mixture distributions. Annals of Statistics, 13(2):795–800, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349557.

Haughton:1988:CMF

[Hau88] Dominique M. A. Haughton. On the choice of a model to fit data from an exponential family. Annals of Statistics, 16(1):342–355, March 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350709.

Hayter:1984:PCT

[Hay84] Anthony J. Hayter. Proof of the conjecture that the Tukey– Kramer multiple comparisons procedure is conservative. Annals of Statistics, 12(1):61–75, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346392;   http://www.
jstor.org/stable/2241034.

Hwang:1982:MCS

[HC82] Jiunn Tzon Hwang and George Casella. Minimax confidence sets for the mean of a multivariate normal distribution. Annals of Statistics, 10(3):868–881, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345877.

Hwang:1986:ICS

[HC86] Jiunn Tzon Hwang and Jeesen Chen. Improved confidence sets for the coefficients of a linear model with spherically symmetric errors. Annals of Statistics, 14(2):444–460, June 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349932.

Hemerly:1989:SCP

[HD89] E. M. Hemerly and M. H. A. Davis. Strong consistency of the PLS criterion for order determination of autoregressive processes. Annals of Statistics, 17(2):941–946, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347154.

Hall:1989:SB

[HDR89] Peter Hall, Thomas J. DiCiccio, and Joseph P. Romano. On smoothing and the bootstrap. Annals of Statistics, 17(2):692–704, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347135.

Heckman:1985:LLT

[Hec85a] Nancy E. Heckman. A local limit theorem for a biased coin design for sequential tests. Annals of Statistics, 13(2):785–788, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349555.

Heckman:1985:SPR

[Hec85b] Nancy E. Heckman. A sequential probability ratio test using a biased coin design. Annals of Statistics, 13(2):789–794, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349556.

Helmers:1980:EEL

[Hel80] R. Helmers. Edgeworth expansions for linear combinations of order statistics with smooth weight functions. Annals of Statistics, 8(6): 1361–1374, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345207.

Henze:1988:MTS

[Hen88] Norbert Henze. A multivariate two-sample test based on the number of nearest neighbor type coincidences. Annals of Statistics, 16(2):772–783, June 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350835.

Hasza:1982:TNP

[HF82] David P. Hasza and Wayne A. Fuller. Testing for nonstationary parameter specifications in seasonal time series models. Annals of Statistics, 10(4):1209–1216, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345985.

Hannum:1983:CRF

[HH83a] Robert Hannum and Myles Hollander. Corrections: “Robustness of Ferguson’s Bayes Estimator of a Distribution Function”. Annals of Statistics, 11(4):1267, December 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346342. See [HH83b].

Hannum:1983:RFB

[HH83b] Robert Hannum and Myles Hollander. Robustness of Ferguson’s Bayes estimator of a distribution function. Annals of Statistics, 11(2):632–639, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346168. See correction [HH83a].

Hartigan:1985:DTU

[HH85] J. A. Hartigan and P. M. Hartigan. The dip test of unimodality. Annals of Statistics, 13(1):70–84, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346577.

Hannan:1987:DWA

[HH87] E. J. Hannan and C. H. Hesse. Discussion: What is an analysis of variance? Annals of Statistics, 15(3):923–924, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350478.

Hinkley:1979:PL

[Hin79] David Hinkley. Predictive likelihood. Annals of Statistics, 7 (4):718–728, July 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176344723. See correction [Hin80b].

Hinkley:1980:ETP

[Hin80a] David V. Hinkley. Editorial: Tradition, progress and the annals of statistics. Annals of Statistics, 8(1):??, January 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344888.

Hinkley:1980:NCP

[Hin80b] David V. Hinkley. Note: Correction to “Predictive Likelihood”. Annals of Statistics, 8(3):694, May 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345022. See [Hin79].

Hinkley:1986:DJB

[Hin86] David Hinkley. Discussion: Jackknife, bootstrap and other resampling methods in regression analysis. Annals of Statistics, 14(4): 1312–1316, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350150.

Hallin:1985:LSR

[HIP85] Marc Hallin, Jean-François Ingenbleek, and Madan L. Puri. Linear serial rank tests for randomness against ARMA alternatives. Annals of Statistics, 13(3):1156–1181, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349662.

Hjort:1986:DCB

[Hjo86] Nils L. Hjort. Discussion: On the consistency of Bayes estimates. Annals of Statistics, 14(1):49–55, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349838.

Hardle:1988:SUC

[HJS88] W. Hardle, P. Janssen, and R. Serfling. Strong uniform consistency rates for estimators of conditional functionals. Annals of Statistics, 16(4):1428–1449, December 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176351047.

Hardle:1984:UCC

[HL84] W. Hardle and S. Luckhaus. Uniform consistency of a class of regression function estimators. Annals of Statistics, 12(2):612–623, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346509.

Hedayat:1989:CSD

[HLS89] A. Hedayat, Bing-Ying Lin, and J. Stufken. The construction of ΠPS sampling designs through a method of emptying boxes.

Annals of Statistics, 17(4):1886–1905, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347400.

Hardle:1985:OBS

[HM85a] Wolfgang Hardle and James Stephen Marron. Optimal bandwidth selection in nonparametric regression function estimation. Annals of Statistics, 13(4):1465–1481, December 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349748.

Hedayat:1985:FOB

[HM85b] A. S. Hedayat and Dibyen Majumdar. Families of A-optimal block designs for comparing test treatments with a control. Annals of Statistics, 13(2):757–767, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349552.

Hall:1987:ANI

[HM87a] Peter Hall and J. S. Marron. On the amount of noise inherent in bandwidth selection for a kernel density estimator. Annals of Statistics, 15(1):163–181, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350259.

Hsieh:1987:MCE

[HM87b] David A. Hsieh and Charles F. Manski. Monte Carlo evidence on adaptive maximum likelihood estimation of a regression. Annals of Statistics, 15(2):541–551, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350359.

Hall:1988:CKO

[HM88] Peter Hall and J. S. Marron. Choice of kernel order in density estimation. Annals of Statistics, 16(1):161–173, March 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350697.

Hoel:1981:RSW

[Hoe81] Paul G. Hoel. Regression systems for which optimal extrapolation designs require exactly k + 1 points. Annals of Statistics, 9(4):909–912, July 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345532.

Hooper:1982:ICS

[Hoo82a] Peter M. Hooper. Invariant confidence sets with smallest expected measure. Annals of Statistics, 10(4):1283–1294, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345994. See

correction [Hoo84b].

Hooper:1982:SIC

[Hoo82b] Peter M. Hooper. Sufficiency and invariance in confidence set estimation. Annals of Statistics, 10(2):549–555, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345795.

Hooper:1983:SIE

[Hoo83] Peter M. Hooper. Simultaneous interval estimation in the general multivariate analysis of variance model. Annals of Statistics, 11(2):666–673, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346171. See correction [Hoo84c].

Hooper:1984:CPI

[Hoo84a] Peter M. Hooper. A conditional property of invariant confidence and prediction regions. Annals of Statistics, 12(2):745–750, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346520.

Hooper:1984:CIC

[Hoo84b] Peter M. Hooper. Corrections: “Invariant Confidence Sets with Smallest Expected Measure”. Annals of Statistics, 12(2):784, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346527. See [Hoo82a].

Hooper:1984:CSI

[Hoo84c] Peter M. Hooper. Corrections: “Simultaneous Interval Estimation in the General Multivariate Analysis of Variance Model”. Annals of Statistics, 12(2):785, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346528. See [Hoo83].

Hooper:1988:SEP

[Hoo88] Peter M. Hooper. Simultaneous estimation and prediction using the expected coverage measure criterion. Annals of Statistics, 16 (1):265–277, March 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350704.

Hooper:1989:MRO

[Hoo89] Peter M. Hooper. Minimaxity of randomized optimal designs. Annals of Statistics, 17(3):1315–1324, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347271.

Horn:1988:SOA

[Hor88] Paul S. Horn. On the stochastic ordering of absolute univariate Gaussian random variables. Annals of Statistics, 16(3):1327–1329, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350964.

Hosoya:1989:BCL

[Hos89] Yuzo Hosoya. The bracketing condition for limit theorems on stationary linear processes. Annals of Statistics, 17(1):401–418, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347024.

Hochberg:1986:ODR

[HP86] Yosef Hochberg and Marc E. Posner. On optimal decision rules for signs of parameters. Annals of Statistics, 14(2):733–742, June 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349950.

Hallin:1988:ORB

[HP88a] Marc Hallin and Madan L. Puri. Optimal rank-based procedures for time series analysis: Testing an ARMA model against other ARMA models. Annals of Statistics, 16(1):402–432, March 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350712.

Hannan:1988:UCC

[HP88b] E. J. Hannan and D. S. Poskitt. Unit canonical correlations between future and past. Annals of Statistics, 16(2):784–790, June 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350836.

Holland:1986:CAU

[HR86] Paul W. Holland and Paul R. Rosenbaum. Conditional association and unidimensionality in monotone latent variable models. Annals of Statistics, 14(4):1523–1543, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350174.

Herrmann:1980:ESS

[HS80] Nira Herrmann and Ted H. Szatrowski. Expected sample size savings from curtailed procedures for the t-test and Hotelling’s T^{2}. Annals of Statistics, 8(3):682–686, May 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345019.

Houtman:1983:BDE

[HS83] A. M. Houtman and T. P. Speed. Balance in designed experiments with orthogonal block structure. Annals of Statistics, 11(4): 1069–1085, December 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346322.

Herrmann:1987:SSS

[HS87] Nira Herrmann and Ted H. Szatrowski. Sample size savings for curtailed one-sample nonparametric tests for location shift. Annals of Statistics, 15(1):296–313, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350267.

Heath:1989:CII

[HS89a] David Heath and William Sudderth. Coherent inference from improper priors and from finitely additive priors. Annals of Statistics, 17(2):907–919, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347150.

Hedayat:1989:MNC

[HS89b] A. Hedayat and J. Stufken. On the maximum number of constraints in orthogonal arrays. Annals of Statistics, 17(1):448–451, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347030.

Hsu:1981:SCI

[Hsu81] Jason C. Hsu. Simultaneous confidence intervals for all distances from the “best”. Annals of Statistics, 9(5):1026–1034, September 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345582.

Hsu:1984:CSC

[Hsu84] Jason C. Hsu. Constrained simultaneous confidence intervals for multiple comparisons with the best. Annals of Statistics, 12(3): 1136–1144, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346732.

Hosoya:1982:CLT

[HT82] Yuzo Hosoya and Masanobu Taniguchi. A central limit theorem for stationary processes and the parameter estimation of linear processes. Annals of Statistics, 10(1):132–153, March 1982. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345696. See

correction [HT93].

Haeusler:1985:ANH

[HT85a] E. Haeusler and J. L. Teugels. On asymptotic normality of Hill’s estimator for the exponent of regular variation. Annals of Statistics, 13(2):743–756, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349551.

Hastie:1985:DPP

[HT85b] Trevor Hastie and Robert Tibshirani. Discussion: Projection pursuit. Annals of Statistics, 13(2):502–508, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349528.

Habib:1986:CSG

[HT86] M. G. Habib and D. R. Thomas. Chi-square goodness-of-fit tests for randomly censored data. Annals of Statistics, 14(2):759–765, June 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349953.

Hardle:1988:RNR

[HT88] W. Hardle and A. B. Tsybakov. Robust nonparametric regression with simultaneous scale curve estimation. Annals of Statistics, 16(1):120–135, March 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350694.

Hill:1989:OPI

[HT89] Theodore P. Hill and Y. L. Tong. Optimal-partitioning inequalities in classification and multi-hypotheses testing. Annals of Statistics, 17(3):1325–1334, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347272.

Hosoya:1993:CCL

[HT93] Yuzo Hosoya and Masanobu Taniguchi. Correction: “A Central Limit Theorem for Stationary Processes and the Parameter Estimation of Linear Processes. Annals of Statistics, 21 (2):1115–1117, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349167. See [HT82].

Hu:1985:UBT

[Hu85] Inchi Hu. A uniform bound for the tail probability of Kolmogorov– Smirnov statistics. Annals of Statistics, 13(2):821–826, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349561.

Hu:1988:RST

[Hu88] Inchi Hu. Repeated significance tests for exponential families. Annals of Statistics, 16(4):1643–1666, December 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176351059.

Huber:1984:FSB

[Hub84] Peter J. Huber. Finite sample breakdown of M- and P - estimators. Annals of Statistics, 12(1):119–126, March 1984. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346396.

Huber:1985:PP

[Hub85a] Peter J. Huber. Projection pursuit. Annals of Statistics, 13 (2):435–475, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349519.

Huber:1985:RPP

[Hub85b] Peter J. Huber. Rejoinder: Projection pursuit. Annals of Statistics, 13(2):522–525, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349535.

Hudson:1983:DCI

[Hud83] H. Malcolm Hudson. Discussion: “Construction of Improved Estimators in Multiparameter Estimation for Discrete Exponential Families”. Annals of Statistics, 11(2):370–371, June 1983. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346145. See

[GHT83].

Hudson:1985:AES

[Hud85] H. M. Hudson. Adaptive estimators for simultaneous estimation of Poisson means. Annals of Statistics, 13(1):246–261, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346590. See

correction [Hud86].

Hudson:1986:CAE

[Hud86] H. M. Hudson. Correction: “Adaptive Estimators for Simultaneous Estimation of Poisson Means”. Annals of Statistics, 14(1):360, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349865. See [Hud85].

Huffman:1983:EAS

[Huf83] Michael D. Huffman. An efficient approximate solution to the Kiefer–Weiss problem. Annals of Statistics, 11(1):306–316, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346081.

Hall:1984:BAR

[HW84] Peter Hall and A. H. Welsh. Best attainable rates of convergence for estimates of parameters of regular variation. Annals of Statistics, 12(3):1079–1084, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346723.

Hall:1985:AEP

[HW85] Peter Hall and A. H. Welsh. Adaptive estimates of parameters of regular variation. Annals of Statistics, 13(1):331–341, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346596.

Hu:1989:IAA

[HW89] Inchi Hu and C. Z. Wei. Irreversible adaptive allocation rules. Annals of Statistics, 17(2):801–823, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347144.

Hwang:1982:IUS

[Hwa82a] Jiunn Tzon Hwang. Improving upon standard estimators in discrete exponential families with applications to Poisson and negative binomial cases. Annals of Statistics, 10(3):857–867, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345876.

Hwang:1982:STU

[Hwa82b] Jiunn Tzon Hwang. Semi tail upper bounds on the class of admissible estimators in discrete exponential families with applications to Poisson and negative binomial distributions. Annals of Statistics, 10(4):1137–1147, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345979.

Hwang:1985:UDS

[Hwa85] Jiunn Tzon Hwang. Universal domination and stochastic domination: Estimation simultaneously under a broad class of loss functions. Annals of Statistics, 13(1):295–314, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346594.

Hamilton:1982:AIN

[HWB82] David C. Hamilton, Donald G. Watts, and Douglas M. Bates. Accounting for intrinsic nonlinearity in nonlinear regression parameter inference regions. Annals of Statistics, 10(2):386–393, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345780.

Horvath:1987:CRB

[HY87] Lajos Horváth and Brian S. Yandell. Convergence rates for the bootstrapped product-limit process. Annals of Statistics, 15(3): 1155–1173, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350498.

Iachan:1983:ATS

[Iac83] Ronaldo Iachan. Asymptotic theory of systematic sampling. Annals of Statistics, 11(3):959–969, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346261.

Irle:1980:NST

[IK80] Albrecht Irle and Karl-Heinz Klosener. Note on the sign test in the presence of ties. Annals of Statistics, 8(5):1168–1170, September 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345155.

Irle:1984:EOS

[Irl84] Albrecht Irle. Extended optimality of sequential probability ratio tests. Annals of Statistics, 12(1):380–386, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346416.

Jacobsen:1989:RCM

[Jac89] Martin Jacobsen. Right censoring and martingale methods for failure time data. Annals of Statistics, 17(3):1133–1156, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347260.

Janssen:1986:APN

[Jan86] Arnold Janssen. Asymptotic properties of Neyman–Pearson tests for infinite Kullback–Leibler information. Annals of Statistics, 14(3):1068–1079, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350050.

Jewell:1983:CCPa

[JB83] Nicholas P. Jewell and Peter Bloomfield. Canonical correlations of past and future for time series: Definitions and theory. Annals of Statistics, 11(3):837–847, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346250.

Jewell:1983:CCPb

[JBB83] Nicholas P. Jewell, Peter Bloomfield, and Flavio C. Bartmann. Canonical correlations of past and future for time series: Bounds and computation. Annals of Statistics, 11(3):848–855, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346251.

Joag-Dev:1983:NAR

[JDP83] Kumar Joag-Dev and Frank Proschan. Negative association of random variables with applications. Annals of Statistics, 11 (1):286–295, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346079.

Jeganathan:1987:SCD

[Jeg87] P. Jeganathan. Strong convergence of distributions of estimators. Annals of Statistics, 15(4):1699–1708, December 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350619.

Jeganathan:1988:SAD

[Jeg88] P. Jeganathan. On the strong approximation of the distributions of estimators in linear stochastic models, I and II: Stationary and explosive AR models. Annals of Statistics, 16(3): 1283–1314, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350962.

Jensen:1987:SLL

[Jen87] J. L. Jensen. Standardized log-likelihood ratio statistics for mixtures of discrete and continuous observations. Annals of Statistics, 15(1):314–324, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350268.

Jensen:1988:CHW

[Jen88] Soren Tolver Jensen. Covariance hypotheses which are linear in both the covariance and the inverse covariance. Annals of Statistics, 16(1):302–322, March 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350707.

Jewell:1982:MED

[Jew82] Nicholas P. Jewell. Mixtures of exponential distributions. Annals of Statistics, 10(2):479–484, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345789.

Janssen:1985:RCO

[JJV85] P. Janssen, J. Jureckova, and N. Veraverbeke. Rate of convergence of one- and two-step M-estimators with applications to maximum likelihood and Pitman estimators. Annals of Statistics, 13(3): 1222–1229, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349666.

James:1980:NCD

[JM80] Ian R. James and James E. Mosimann. A new characterization of the Dirichlet distribution through neutrality. Annals of Statistics, 8(1):183–189, January 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176344900.

Jupp:1982:CMP

[JM82] P. E. Jupp and K. V. Mardia. A characterization of the multivariate Pareto distribution. Annals of Statistics, 10(3):1021–1024, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345894.

Jacroux:1983:OSB

[JN83] Mike Jacroux and William Notz. On the optimality of spring balance weighing designs. Annals of Statistics, 11(3):970–978, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346262.

Jockel:1986:FSP

[Joc86] Karl-Heinz Jockel. Finite sample properties and asymptotic efficiency of Monte Carlo tests. Annals of Statistics, 14(1):336–347, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349860.

Joe:1988:MEP

[Joe88] Harry Joe. Majorization, entropy and paired comparisons. Annals of Statistics, 16(2):915–925, June 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350843.

Johnstone:1984:ADE

[Joh84] Iain Johnstone. Admissibility, difference equations and recurrence in estimating a Poisson mean. Annals of Statistics, 12(4): 1173–1198, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346786.

Jones:1985:DPP

[Jon85] M. C. Jones. Discussion: Projection pursuit. Annals of Statistics, 13(2):508–510, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349529.

Jones:1987:CHC

[Jon87] Lee K. Jones. On a conjecture of Huber concerning the convergence of projection pursuit regression. Annals of Statistics, 15(2):880–882, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350382.

Joshi:1975:CBR

[Jos75] V. M. Joshi. A conjecture of Berry regarding a Bernoulli twoarmed bandit. Annals of Statistics, 3(1):189–202, January 1975.

CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343007. See

correction [Jos85].

Joshi:1985:CCB

[Jos85] V. M. Joshi. Correction: “A Conjecture of Berry Regarding a Bernoulli Two-Armed Bandit”. Annals of Statistics, 13(3):1249, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349670. See [Jos75].

John:1983:EEP

[JR83] R. D. John and J. Robinson. Edgeworth expansions for the power of permutation tests. Annals of Statistics, 11(2):625–631, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346167.

Jupp:1983:STS

[JS83] P. E. Jupp and B. D. Spurr. Sobolev tests for symmetry of directional data. Annals of Statistics, 11(4):1225–1231, December 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346335.

James:1984:CRL

[JS84] I. R. James and P. J. Smith. Consistency results for linear regression with censored data. Annals of Statistics, 12(2):590–600, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346507.

Jupp:1985:STI

[JS85] P. E. Jupp and B. D. Spurr. Sobolev tests for independence of directions. Annals of Statistics, 13(3):1140–1155, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349661.

Johnstone:1989:HFV

[JS89] Iain Johnstone and David Siegmund. On Hotelling’s formula for the volume of tubes and Naiman’s inequality. Annals of Statistics, 17(1):184–194, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347010.

Janssen:1984:ANG

[JSV84] Paul Janssen, Robert Serfling, and Noel Veraverbeke. Asymptotic normality for a general class of statistical functions and applications to measures of spread. Annals of Statistics, 12(4): 1369–1379, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346797.

Jureckova:1981:TBL

[Jur81] Jana Jureckova. Tail-behavior of location estimators. Annals of Statistics, 9(3):578–585, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345461.

Kohn:1983:SPB

[KA83] Robert Kohn and Craig F. Ansley. On the smoothness properties of the best linear unbiased estimate of a stochastic process observed with noise. Annals of Statistics, 11(3):1011–1017, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346270.

Kohn:1989:DLS

[KA89] Robert Kohn and Craig F. Ansley. Discussion: Linear smoothers and additive models. Annals of Statistics, 17(2):535–540, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176347122.

Kabaila:1980:OPL

[Kab80] Paul V. Kabaila. An optimality property of the least-squares estimate of the parameter of the spectrum of a purely nondeterministic time series. Annals of Statistics, 8(5):1082–1092, September 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345145.

Kakigi:1983:NDF

[Kak83] Richard Kakigi. A note on discounted future two-armed bandits. Annals of Statistics, 11(2):707–711, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346176.

Kallenberg:1981:SLM

[Kal81] Wilbert C. M. Kallenberg. The shortcoming of locally most powerful tests in curved exponential families. Annals of Statistics, 9(3):673–677, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345472.

Kallenberg:1982:CED

[Kal82] Wilbert C. M. Kallenberg. Chernoff efficiency and deficiency. Annals of Statistics, 10(2):583–594, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345799.

Kallenberg:1983:IET

[Kal83a] Wilbert C. M. Kallenberg. Intermediate efficiency, theory and examples. Annals of Statistics, 11(1):170–182, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346067.

Kallenberg:1983:MDT

[Kal83b] Wilbert C. M. Kallenberg. On moderate deviation theory in estimation. Annals of Statistics, 11(2):498–504, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346156.

Kallenberg:1985:MLD

[Kal85] Wilbert C. M. Kallenberg. On moderate and large deviations in multinomial distributions. Annals of Statistics, 13(4):1554–1580, December 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349755.

Kariya:1980:LRT

[Kar80] Takeaki Kariya. Locally robust tests for serial correlation in least squares regression. Annals of Statistics, 8(5):1065–1070, September 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345143.

Kariya:1981:RMT

[Kar81a] Takeaki Kariya. Robustness of multivariate tests. Annals of Statistics, 9(6):1267–1275, November 1981. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:/

/projecteuclid.org/euclid.aos/1176345643. See correction

[Kar89a].

Kariya:1981:RPH

[Kar81b] Takeaki Kariya. A robustness property of Hotelling’s T^{2}-test. Annals of Statistics, 9(1):211–214, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345350.

Kariya:1981:TIB

[Kar81c] Takeaki Kariya. Tests for the independence between two seemingly unrelated regression equations. Annals of Statistics, 9(2):381–390, March 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345403.

Karr:1987:MLE

[Kar87] Alan F. Karr. Maximum likelihood estimation in the multiplicative intensity model via sieves. Annals of Statistics, 15(2):473–490, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350356.

Karunamuni:1988:EBT

[Kar88] Rohana J. Karunamuni. On empirical Bayes testing with sequential components. Annals of Statistics, 16(3):1270–1282, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350961.

Kariya:1989:CRM

[Kar89a] Takeaki Kariya. Correction: “Robustness of Multivariate Tests”. Annals of Statistics, 17(2):948, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347156. See [Kar81a].

Kariya:1989:EEM

[Kar89b] Takeaki Kariya. Equivariant estimation in a model with an ancillary statistic. Annals of Statistics, 17(2):920–928, June 1989. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347151.

Kaufmann:1987:RMN

[Kau87] Heinz Kaufmann. Regression models for nonstationary categorical time series: Asymptotic estimation theory. Annals of Statistics, 15(1):79–98, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350254.

Kawashima:1980:PEA

[Kaw80] Hironao Kawashima. Parameter estimation of autoregressive integrated processes by least squares. Annals of Statistics, 8(2):423– 435, March 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344962.

Koenker:1985:BE

[KB85] Roger Koenker and Gilbert Bassett. On Boscovich’s estimator. Annals of Statistics, 13(4):1625–1628, December 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349759.

Koul:1983:MDE

[KD83] H. Koul and T. DeWet. Minimum distance estimation in a linear regression model. Annals of Statistics, 11(3):921–932, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346258.

Kedem:1984:SLS

[Ked84] Benjamin Kedem. On the sinusoidal limit of stationary time series. Annals of Statistics, 12(2):665–674, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346513.

Keenan:1983:LBF

[Kee83] Daniel MacRae Keenan. Limiting behavior of functionals of the sample spectral distribution. Annals of Statistics, 11(4):1206–1217, December 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346333.

Keener:1984:SOE

[Kee84] Robert Keener. Second order efficiency in the sequential design of experiments. Annals of Statistics, 12(2):510–532, June 1984.

CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346503.

Keenan:1985:APM

[Kee85a] Daniel MacRae Keenan. Asymptotic properties of minimization estimators for time series parameters. Annals of Statistics, 13 (1):369–382, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346598.

Keener:1985:FCT

[Kee85b] Robert Keener. Further contributions to the “Two-Armed Bandit” problem. Annals of Statistics, 13(1):418–422, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346603.

Keenan:1987:LBF

[Kee87a] Daniel MacRae Keenan. Limiting behavior of functionals of higher-order sample cumulant spectra. Annals of Statistics, 15 (1):134–151, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350257.

Keener:1987:NVS

[Kee87b] Robert Keener. A note on the variance of a stopping time. Annals of Statistics, 15(4):1709–1712, December 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350620.

Kelly:1981:MAB

[Kel81] F. P. Kelly. Multi-armed bandits with discount factor near one: The Bernoulli case. Annals of Statistics, 9(5):987–1001, September 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345578.

Kelly:1984:IFE

[Kel84] Gabrielle Kelly. The influence function in the errors in variables problem. Annals of Statistics, 12(1):87–100, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346394.

Kelly:1989:SRL

[Kel89] Robert E. Kelly. Stochastic reduction of loss in estimating normal means by isotonic regression. Annals of Statistics, 17(2):937–940, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347153.

Kempthorne:1987:DWA

[Kem87a] Oscar Kempthorne. Discussion: What is an analysis of variance? Annals of Statistics, 15(3):925–929, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350479.

Kempthorne:1987:EMN

[Kem87b] Peter J. Kempthorne. Estimating the mean of a normal distribution with loss equal to squared error plus complexity cost. Annals of Statistics, 15(4):1389–1400, December 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350600.

Kempthorne:1988:CRU

[Kem88] Peter J. Kempthorne. Controlling risks under different loss functions: The compromise decision problem. Annals of Statistics, 16 (4):1594–1608, December 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176351055.

Kent:1983:IFM

[Ken83] John T. Kent. Identifiability of finite mixtures for directional data. Annals of Statistics, 11(3):984–988, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346264.

Kershaw:1985:APE

[Ker85] Christopher D. Kershaw. Asymptotic properties of w̄, an estimator of the ED50 suggested for use in up-and-down experiments in bioassay. Annals of Statistics, 13(1):85–94, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346578.

Kneip:1988:CCR

[KG88a] Alois Kneip and Theo Gasser. Convergence and consistency results for self-modeling nonlinear regression. Annals of Statistics, 16(1):82–112, March 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350692.

Ko:1988:RED

[KG88b] Daijin Ko and Peter Guttorp. Robustness of estimators for directional data. Annals of Statistics, 16(2):609–618, June 1988. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350822.

Khmaladze:1988:IAG

[Khm88] E. V. Khmaladze. An innovation approach to goodness-of-fit tests in R^{m}. Annals of Statistics, 16(4):1503–1516, December 1988. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176351051.

Krieger:1981:WCE

[KI81] A. M. Krieger and J. Pickands III. Weak convergence and efficient density estimation at a point. Annals of Statistics, 9(5): 1066–1078, September 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345586.

Kildea:1981:BMT

[Kil81] D. G. Kildea. Brown-mood type median estimators for simple regression models. Annals of Statistics, 9(2):438–442, March 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345410.

King:1980:RTS

[Kin80] M. L. King. Robust tests for spherical symmetry and their application to least squares regression. Annals of Statistics, 8(6): 1265–1271, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345199.

Kulkarni:1986:OPB

[KJ86] Radhika V. Kulkarni and Christopher Jennison. Optimal properties of the Bechhofer–Kulkarni Bernoulli selection procedure. Annals of Statistics, 14(1):298–314, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349857.

Kester:1986:LDE

[KK86] A. D. M. Kester and W. C. M. Kallenberg. Large deviations of estimators. Annals of Statistics, 14(2):648–664, June 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349944.

Kallenberg:1987:LNM

[KL87] Wilbert C. M. Kallenberg and Teresa Ledwina. On local and nonlocal measures of efficiency. Annals of Statistics, 15(4):1401–1420, December 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350601.

Koul:1989:WCR

[KL89] Hira L. Koul and Shlomo Levental. Weak convergence of the residual empirical process in explosive autoregression. Annals of Statistics, 17(4):1784–1794, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347395.

Klaassen:1984:LES

[Kla84] Chris A. J. Klaassen. Location estimators and spread. Annals of Statistics, 12(1):311–321, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346409.

Klaassen:1987:CEI

[Kla87] Chris A. J. Klaassen. Consistent estimation of the influence function of locally asymptotically linear estimators. Annals of Statistics, 15(4):1548–1562, December 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350609.

Klaassen:1989:ES

[Kla89] Chris A. J. Klaassen. Estimators and spread. Annals of Statistics, 17(2):859–867, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347147.

Klonias:1982:CTN

[Klo82] V. K. Klonias. Consistency of two nonparametric maximum penalized likelihood estimators of the probability density function.

Annals of Statistics, 10(3):811–824, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345873.

Klonias:1984:CND

[Klo84] V. K. Klonias. On a class of nonparametric density and regression estimators. Annals of Statistics, 12(4):1263–1284, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346791.

Kushner:1980:EEV

[KM80] H. B. Kushner and Morris Meisner. Eigenfunctions of expected value operators in the Wishart distribution. Annals of Statistics, 8 (5):977–988, September 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345136.

Kunert:1987:OFW

[KM87] J. Kunert and R. J. Martin. On the optimality of finite Williams II(a) designs. Annals of Statistics, 15(4):1604–1628, December 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350613.

Kulp:1984:AEN

[KN84] R. W. Kulp and B. N. Nagarsenker. An asymptotic expansion of the nonnull distribution of Wilks criterion for testing the multivariate linear hypothesis. Annals of Statistics, 12(4):1576–1583, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346816.

Knight:1989:CAI

[Kni89a] Keith Knight. Consistency of Akaike’s information criterion for infinite variance autoregressive processes. Annals of Statistics, 17(2):824–840, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347145.

Knight:1989:BSM

[Kni89b] Keith Knight. On the bootstrap of the sample mean in the infinite variance case. Annals of Statistics, 17(3):1168–1175, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347262.

Koenker:1987:DTM

[Koe87] Roger Koenker. Discussion: The trimmed mean in the linear model. Annals of Statistics, 15(1):39–44, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350250.

Kogure:1987:AOC

[Kog87] Atsuyuki Kogure. Asymptotically optimal cells for a histogram. Annals of Statistics, 15(3):1023–1030, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350490.

Koschat:1987:CFS

[Kos87] Martin A. Koschat. A characterization of the Fieller solution. Annals of Statistics, 15(1):462–468, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350282.

Kourouklis:1984:BOS

[Kou84a] Stavros Kourouklis. Bahadur optimality of sequential experiments for exponential families. Annals of Statistics, 12(4):1522–1527, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346808.

Kourouklis:1984:LDR

[Kou84b] Stavros Kourouklis. A large deviation result for the likelihood ratio statistic in exponential families. Annals of Statistics, 12(4): 1510–1521, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346807.

Koul:1986:MDE

[Kou86] Hira L. Koul. Minimum distance estimation and goodness-of-fit tests in first-order autoregression. Annals of Statistics, 14(3): 1194–1213, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350059.

Koyak:1987:MID

[Koy87] Robert A. Koyak. On measuring internal dependence in a set of random variables. Annals of Statistics, 15(3):1215–1228, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350501.

Kozek:1982:TCA

[Koz82] Andrzej Kozek. Towards a calculus for admissibility. Annals of Statistics, 10(3):825–837, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345874.

Kitchin:1981:GBS

[KP81] John Kitchin and Frank Proschan. Generalization of Block–Savits’ convolution result. Annals of Statistics, 9(2):437, March 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345409.

Krasker:1986:DCB

[KP86] William S. Krasker and John W. Pratt. Discussion: On the consistency of Bayes estimates. Annals of Statistics, 14(1):55–58, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349839.

Karlin:1981:TPP

[KR81a] Samuel Karlin and Yosef Rinott. Total positivity properties of absolute value multinormal variables with applications to confidence interval estimates and related probabilistic inequalities. Annals of Statistics, 9(5):1035–1049, September 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345583.

Krewski:1981:ISS

[KR81b] D. Krewski and J. N. K. Rao. Inference from stratified samples: Properties of the linearization, jackknife and balanced repeated replication methods. Annals of Statistics, 9(5):1010–1019, September 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345580.

Karlin:1982:AAT

[KR82] Samuel Karlin and Yosef Rinott. Applications of anova type decompositions for comparisons of conditional variance statistics including jackknife estimates. Annals of Statistics, 10(2):485–501, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345790.

Kunte:1984:RRM

[KR84] S. Kunte and R. N. Rattihalli. Rectangular regions of maximum probability content. Annals of Statistics, 12(3):1106–1108, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346727.

Krasker:1984:NSP

[Kra84] William S. Krasker. A note on selecting parametric models in Bayesian inference. Annals of Statistics, 12(2):751–757, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346521.

Kredler:1986:BTO

[Kre86] Christian Kredler. Behaviour of third order terms in quadratic approximations of LR-statistics in multivariate generalized linear models. Annals of Statistics, 14(1):326–335, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349859.

Kreiss:1987:AES

[Kre87] Jens-Peter Kreiss. On adaptive estimation in stationary ARMA processes. Annals of Statistics, 15(1):112–133, March 1987. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350256.

Keener:1987:DP

[KRS87] Robert Keener, Edward Rothman, and Norman Starr. Distributions on partitions. Annals of Statistics, 15(4):1466–1481, December 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350604.

Kadane:1977:OSC

[KS77] Joseph B. Kadane and Herbert A. Simon. Optimal strategies for a class of constrained sequential problems. Annals of Statistics, 5(2):237–255, March 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343791. See correction [KS83].

Kedem:1982:TSD

[KS82] Benjamin Kedem and Eric Slud. Time series discrimination by higher order crossings. Annals of Statistics, 10(3):786–794, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345871.

Kadane:1983:COS

[KS83] Joseph B. Kadane and Herbert A. Simon. Corrections: “Optimal Strategies for a Class of Constrained Sequential Problems”. Annals of Statistics, 11(1):346, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346087. See [KS77].

Krafft:1984:KCR

[KS84] O. Krafft and M. Schaefer. On Karlin’s conjecture for random replacement sampling plans. Annals of Statistics, 12(4): 1528–1535, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346809.

Kariya:1985:NOR

[KS85] Takeaki Kariya and Bimal Kumar Sinha. Nonnull and optimality robustness of some tests. Annals of Statistics, 13(3):1182–1197, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349663.

Koul:1981:RAR

[KSV81] H. Koul, V. Susarla, and J. Van Ryzin. Regression analysis with randomly right-censored data. Annals of Statistics, 9(6): 1276–1288, November 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345644.

Kubokawa:1987:AME

[Kub87] Tatsuya Kubokawa. Admissible minimax estimation of a common mean of two normal populations. Annals of Statistics, 15(3): 1245–1256, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350503.

Kumazawa:1987:TWN

[Kum87] Yoshiki Kumazawa. On testing whether new is better than used using randomly censored data. Annals of Statistics, 15(1):420–426, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350276.

Kunert:1983:ODR

[Kun83] Joachim Kunert. Optimal design and refinement of the linera model with applications to repeated measurements designs. Annals of Statistics, 11(1):247–257, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346075.

Kunert:1984:OBU

[Kun84a] Joachim Kunert. Optimality of balanced uniform repeated measurements designs. Annals of Statistics, 12(3):1006–1017, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346717.

Kunsch:1984:IRA

[Kun84b] H. Kunsch. Infinitesimal robustness for autoregressive processes. Annals of Statistics, 12(3):843–863, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346706.

Kunsch:1986:DIF

[Kun86] Hans R. Kunsch. Discussion: Influence functionals for time series. Annals of Statistics, 14(3):824–826, September 1986. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350030.

Kunsch:1989:JBG

[Kun89] Hans R. Kunsch. The jackknife and the bootstrap for general stationary observations. Annals of Statistics, 17(3):1217–1241, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347265.

Kuo:1983:BBD

[Kuo83] Lynn Kuo. Bayesian bioassay design. Annals of Statistics, 11(3): 886–895, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346254.

Kiefer:1981:OBB

[KW81] J. Kiefer and H. P. Wynn. Optimum balanced block and Latin square designs for correlated observations. Annals of Statistics, 9(4):737–757, July 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345515.

Kiefer:1984:OME

[KW84] J. Kiefer and H. P. Wynn. Optimum and minimax exact treatment designs for one-dimensional autoregressive error processes. Annals of Statistics, 12(2):431–450, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346498.

Lai:1981:AOI

[Lai81] Tze Leung Lai. Asymptotic optimality of invariant sequential probability ratio tests. Annals of Statistics, 9(2):318–333, March 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345398.

Lai:1987:ATA

[Lai87] Tze Leung Lai. Adaptive treatment allocation and the multiarmed bandit problem. Annals of Statistics, 15(3):1091–1114, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350495.

Lai:1988:NOS

[Lai88] Tze Leung Lai. Nearly optimal sequential tests of composite hypotheses. Annals of Statistics, 16(2):856–886, June 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350840.

LaMotte:1982:ALE

[LaM82] Lynn Roy LaMotte. Admissibility in linear estimation. Annals of Statistics, 10(1):245–255, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345707.

Lambert:1985:RTS

[Lam85] Diane Lambert. Robust two-sample permutation tests. Annals of Statistics, 13(2):606–625, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349542.

LaRiccia:1982:APW

[LaR82] Vincent N. LaRiccia. Asymptotic properties of weighted L^{2} quantile distance estimators. Annals of Statistics, 10(2):621–624, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345803.

Lu:1989:ENM

[LB89] K. L. Lu and James O. Berger. Estimation of normal means: Frequentist estimation of loss. Annals of Statistics, 17(2):890–906, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347149.

Lo:1987:BPR

[LC87] Albert Y. Lo and Javier Cabrera. Bayes procedures for rotationally symmetric models on the sphere. Annals of Statistics, 15(3): 1257–1268, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350504.

Li:1989:RAU

[LD89] Ker-Chau Li and Naihua Duan. Regression analysis under link violation. Annals of Statistics, 17(3):1009–1052, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347254.

LeCam:1986:DCB

[Le 86] L. Le Cam. Discussion: On the consistency of Bayes estimates. Annals of Statistics, 14(1):59–60, March 1986. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349840.

Lee:1981:QLI

[Lee81] Chu-In Charles Lee. The quadratic loss of isotonic regression under normality. Annals of Statistics, 9(3):686–688, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345475.

Lee:1983:MMA

[Lee83] Chu-In Charles Lee. The min-max algorithm and isotonic regression. Annals of Statistics, 11(2):467–477, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346153.

Lee:1988:QLO

[Lee88] Chu-In Charles Lee. Quadratic loss of order restricted estimators for treatment means with a control. Annals of Statistics, 16(2):751–758, June 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350833.

Lehmann:1988:CLE

[Leh88] E. L. Lehmann. Comparing location experiments. Annals of Statistics, 16(2):521–533, June 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350818.

Leonard:1985:DTI

[Leo85] Tom Leonard. Discussion: Testing for independence in a twoway table: New interpretations of the chi-square statistic. Annals of Statistics, 13(3):893–898, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349640.

Lerche:1986:SBT

[Ler86] Hans Rudolf Lerche. The shape of Bayes tests of power one. Annals of Statistics, 14(3):1030–1048, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350048.

Letac:1981:ISS

[Let81] Gerard Letac. Isotropy and sphericity: Some characterisations of the normal distribution. Annals of Statistics, 9(2):408–417, March 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345406.

Leurgans:1982:ADS

[Leu82] Sue Leurgans. Asymptotic distributions of slope-of-greatest-convex-minorant estimators. Annals of Statistics, 10(1):287–296, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345711.

Leurgans:1984:ABT

[Leu84] Sue Leurgans. Asymptotic behavior of two-sample rank tests in the presence of random censoring. Annals of Statistics, 12 (2):572–589, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346506.

Levin:1981:RMC

[Lev81] Bruce Levin. A representation for multinomial cumulative distribution functions. Annals of Statistics, 9(5):1123–1126, September 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345593.

Lambert:1982:ALV

[LH82] Diane Lambert and W. J. Hall. Asymptotic lognormality of P - values. Annals of Statistics, 10(1):44–64, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345689. See cor-

rection [LH83].

Lambert:1983:CAL

[LH83] D. Lambert and W. J. Hall. Corrections: Asymptotic lognormality of P -values. Annals of Statistics, 11(1):348, March 1983. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346089. See

[LH82].

Li:1984:DSA

[LH84] Ker-Chau Li and Jiunn Tzon Hwang. The data-smoothing aspect of Stein estimates. Annals of Statistics, 12(3):887–897, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346709.

Li:1982:MMR

[Li82] Ker-Chau Li. Minimaxity of the method of regularization of stochastic processes. Annals of Statistics, 10(3):937–942, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345883.

Li:1983:MRD

[Li83] Ker-Chau Li. Minimaxity for randomized designs: Some general results. Annals of Statistics, 11(1):225–239, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346073.

Li:1984:CCV

[Li84a] Ker-Chau Li. Consistency for cross-validated nearest neighbor estimates in nonparametric regression. Annals of Statistics, 12 (1):230–240, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346403.

Li:1984:RMI

[Li84b] Ker-Chau Li. Regression models with infinitely many parameters: Consistency of bounded linear functionals. Annals of Statistics, 12(2):601–611, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346508.

Li:1984:RRD

[Li84c] Ker-Chau Li. Robust regression designs when the design space consists of finitely many points. Annals of Statistics, 12(1):269–282, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346406.

Li:1985:SUR

[Li85] Ker-Chau Li. From Stein’s unbiased risk estimates to the method of generalized cross validation. Annals of Statistics, 13(4):1352– 1377, December 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349742.

Li:1986:AOG

[Li86] Ker-Chau Li. Asymptotic optimality of C_{L} and generalized crossvalidation in ridge regression with application to spline smoothing. Annals of Statistics, 14(3):1101–1112, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350052.

Li:1987:AOC

[Li87] Ker-Chau Li. Asymptotic optimality for C_{p}, C_{L}, cross-validation and generalized cross-validation: Discrete index set. Annals of Statistics, 15(3):958–975, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350486.

Li:1989:HCR

[Li89] Ker-Chau Li. Honest confidence regions for nonparametric regression. Annals of Statistics, 17(3):1001–1008, September 1989. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347253.

Liang:1988:CRE

[Lia88] TaChen Liang. On the convergence rates of empirical Bayes rules for two-action problems: Discrete case. Annals of Statistics, 16(4): 1635–1642, December 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176351058.

Lindley:1980:JSH

[Lin80] D. V. Lindley. L. J. Savage — his work in probability and statistics. Annals of Statistics, 8(1):1–24, January 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344889.

Lindsay:1983:ECS

[Lin83a] B. G. Lindsay. Efficiency of the conditional score in a mixture setting. Annals of Statistics, 11(2):486–497, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346155.

Lindsay:1983:GMLa

[Lin83b] Bruce G. Lindsay. The geometry of mixture likelihoods: A general theory. Annals of Statistics, 11(1):86–94, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346059.

Lindsay:1983:GMLb

[Lin83c] Bruce G. Lindsay. The geometry of mixture likelihoods, Part II: The exponential family. Annals of Statistics, 11(3):783–792, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346245.

Lindsay:1985:UEP

[Lin85] Bruce G. Lindsay. Using empirical partially Bayes inference for increased efficiency. Annals of Statistics, 13(3):914–931, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349646.

Lindley:1986:DCB

[Lin86a] Dennis V. Lindley. Discussion: On the consistency of Bayes estimates. Annals of Statistics, 14(1):60–63, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349841.

Lindsay:1986:EFM

[Lin86b] Bruce G. Lindsay. Exponential family mixture models (with least-squares estimators). Annals of Statistics, 14(1):124–137, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349845.

Lindsay:1989:MMA

[Lin89a] Bruce G. Lindsay. Moment matrices: Applications in mixtures. Annals of Statistics, 17(2):722–740, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347138.

Lindsay:1989:DMM

[Lin89b] Bruce G. Lindsay. On the determinants of moment matrices. Annals of Statistics, 17(2):711–721, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347137.

Liu:1988:BPU

[Liu88] Regina Y. Liu. Bootstrap procedures under some non-I.I.D. models. Annals of Statistics, 16(4):1696–1708, December 1988. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176351062.

LaRiccia:1985:OGF

[LM85] Vincent N. LaRiccia and David M. Mason. Optimal goodness-of-fit tests for location/scale families of distributions based on the sum of squares of L-statistics. Annals of Statistics, 13(1):315–330, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346595.

Leung:1987:EPM

[LM87] Pui Lam Leung and Robb J. Muirhead. Estimation of parameter matrices and eigenvalues in MANOVA and canonical correlation analysis. Annals of Statistics, 15(4):1651–1666, December 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350616.

Lindley:1981:REI

[LN81] D. V. Lindley and Melvin R. Novick. The role of exchangeability in inference. Annals of Statistics, 9(1):45–58, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345331.

Lo:1982:NME

[Lo82] Albert Y. Lo. A note on the minimax estimation of the Poisson intensity function. Annals of Statistics, 10(3):1009–1011, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345891.

Lo:1984:CLM

[Lo84a] Albert Y. Lo. Consistency in the location model: The undominated case. Annals of Statistics, 12(4):1584–1587, December 1984.

CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346817.

Lo:1984:CBN

[Lo84b] Albert Y. Lo. On a class of Bayesian nonparametric estimates: I. Density estimates. Annals of Statistics, 12(1):351–357, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346412.

Lo:1985:ESD

[Lo85] Shaw-Hwa Lo. Estimation of a symmetric distribution. Annals of Statistics, 13(3):1097–1113, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349658.

Lo:1986:BSI

[Lo86a] Albert Y. Lo. Bayesian statistical inference for sampling a finite population. Annals of Statistics, 14(3):1226–1233, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350061.

Lo:1986:EUD

[Lo86b] Shaw-Hwa Lo. Estimation of a unimodal distribution function. Annals of Statistics, 14(3):1132–1138, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350054.

Lo:1987:LSS

[Lo87] Albert Y. Lo. A large sample study of the Bayesian bootstrap. Annals of Statistics, 15(1):360–375, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350271.

Lo:1988:BBF

[Lo88] Albert Y. Lo. A Bayesian bootstrap for a finite population. Annals of Statistics, 16(4):1684–1695, December 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176351061.

Loh:1984:BAR

[Loh84a] Wei-Yin Loh. Bounds on AREs for restricted classes of distributions defined via tail-orderings. Annals of Statistics, 12(2):685– 701, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346515.

Loh:1984:EED

[Loh84b] Wei-Yin Loh. Estimating an endpoint of a distribution with resampling methods. Annals of Statistics, 12(4):1543–1550, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346811.

Loh:1988:DTC

[Loh88] Wei-Yin Loh. Discussion: Theoretical comparison of bootstrap confidence intervals. Annals of Statistics, 16(3):972–976, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350940.

Loh:1989:BST

[Loh89] Wei-Yin Loh. Bounds on the size of the χ^{2}-test of independence in a contingency table. Annals of Statistics, 17(4):1709–1722, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347389.

Lopuhaa:1989:RBE

[Lop89] Hendrik P. Lopuhaa. On the relation between S-estimators and M-estimators of multivariate location and covariance. Annals of Statistics, 17(4):1662–1683, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347386.

Lorden:1983:AET

[Lor83] Gary Lorden. Asymptotic efficiency of three-stage hypothesis tests. Annals of Statistics, 11(1):129–140, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346064.

Loynes:1980:EDF

[Loy80] R. M. Loynes. The empirical distribution function of residuals from generalised regression. Annals of Statistics, 8(2):285–298, March 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344954.

Langberg:1981:EDL

[LPQ81] N. Langberg, F. Proschan, and A. J. Quinzi. Estimating dependent life lengths, with applications to the theory of competing risks. Annals of Statistics, 9(1):157–167, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345342.

Lii:1982:DET

[LR82] K. S. Lii and M. Rosenblatt. Deconvolution and estimation of transfer function phase and coefficients for non-Gaussian linear processes. Annals of Statistics, 10(4):1195–1208, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345984.

Lange:1989:ANR

[LR89] Nicholas Lange and Louise Ryan. Assessing normality in random effects models. Annals of Statistics, 17(2):624–642, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347130.

Laycock:1980:PRM

[LS80] P. J. Laycock and E. Seiden. On a problem of repeated measurement design with treatment additivity. Annals of Statistics, 8(6): 1284–1292, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345201.

Lai:1983:FAE

[LS83a] T. L. Lai and D. Siegmund. Fixed accuracy estimation of an autoregressive parameter. Annals of Statistics, 11(2):478–485, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346154.

Lane:1983:CCI

[LS83b] David A. Lane and William D. Sudderth. Coherent and continuous inference. Annals of Statistics, 11(1):114–120, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346062.

Lane:1985:CPS

[LS85a] David A. Lane and William D. Sudderth. Coherent predictions are strategic. Annals of Statistics, 13(3):1244–1248, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349669.

Lau:1985:ODT

[LS85b] Tai-Shing Lau and W. J. Studden. Optimal designs for trigonometric and polynomial regression using canonical moments. Annals of Statistics, 13(1):383–394, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346599.

Liu:1987:PCB

[LS87] Regina Y. Liu and Kesar Singh. On a partial correction by the bootstrap. Annals of Statistics, 15(4):1713–1718, December 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350621.

Li:1988:MDT

[LS88a] L. A. Li and N. Sedransk. Mixtures of distributions: A topological approach. Annals of Statistics, 16(4):1623–1634, December 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176351057.

Lim:1988:EOD

[LS88b] Yong B. Lim and W. J. Studden. Efficient D_{s}-optimal designs for multivariate polynomial regression on the q-cube. Annals of Statistics, 16(3):1225–1240, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350957.

Liu:1988:DTC

[LS88c] Regina Y. Liu and Kesar Singh. Discussion: Theoretical comparison of bootstrap confidence intervals. Annals of Statistics, 16(3): 978–979, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350942.

Letac:1989:EFE

[LS89] Gerard Letac and Vanamamalai Seshadri. The expectation of X^{1} as a function of E(X) for an exponential family on the positive line. Annals of Statistics, 17(4):1735–1741, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347391.

Leslie:1986:ADS

[LSF86] J. R. Leslie, M. A. Stephens, and S. Fotopoulos. Asymptotic distribution of the Shapiro-wilk W for testing for normality. Annals of Statistics, 14(4):1497–1506, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350172.

Liu:1983:PEQ

[LT83] T. P. Liu and M. E. Thompson. Properties of estimators of quadratic finite population functions: The batch approach. Annals of Statistics, 11(1):275–285, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346078.

Lambert:1984:APM

[LT84] Diane Lambert and Luke Tierney. Asymptotic properties of maximum likelihood estimates in the mixed Poisson model. Annals of Statistics, 12(4):1388–1399, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346799.

Liu:1985:HEH

[LV85] Regina Y. C. Liu and John Van Ryzin. A histogram estimator of the hazard rate with censored data. Annals of Statistics, 13(2):592–605, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349541.

Lai:1982:LSE

[LW82] Tze Leung Lai and Ching Zong Wei. Least squares estimates in stochastic regression models with applications to identification and control of dynamic systems. Annals of Statistics, 10 (1):154–166, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345697.

Lauritzen:1989:CGM

[LW89a] S. L. Lauritzen and N. Wermuth. Correction: “Graphical Models for Associations Between Variables, Some of which are Qualitative and Some Quantitative”. Annals of Statistics, 17(4):1916, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347402. See [LW89b].

Lauritzen:1989:GMA

[LW89b] S. L. Lauritzen and N. Wermuth. Graphical models for associations between variables, some of which are qualitative and some quantitative. Annals of Statistics, 17(1):31–57, March 1989. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347003. See

correction [LW89a].

LeCam:1988:PLA

[LY88] Lucien Le Cam and Grace L. Yang. On the preservation of local asymptotic normality under information loss. Annals of Statistics, 16(2):483–520, June 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350817.

MacGillivray:1986:SAM

[Mac86] H. L. MacGillivray. Skewness and asymmetry: Measures and orderings. Annals of Statistics, 14(3):994–1011, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350046. See

correction [Mac87].

MacGillivray:1987:CSA

[Mac87] H. L. MacGillivray. Correction: “Skewness and Asymmetry: Measures and Orderings”. Annals of Statistics, 15(2):884, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350384. See

[Mac86].

Mallows:1980:STN

[Mal80] C. L. Mallows. Some theory of nonlinear smoothers. Annals of Statistics, 8(4):695–715, July 1980. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345067.

Malley:1983:SAI

[Mal83] James D. Malley. Statistical and algebraic independence. Annals of Statistics, 11(1):341–345, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346086.

Mallows:1985:D

[Mal85] C. L. Mallows. Discussion. Annals of Statistics, 13(1):204–205, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346586.

Mallik:1987:NBP

[Mal87] Ashim K. Mallik. A note on the Buyer’s problem. Annals of Statistics, 15(3):1329–1331, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350510.

Mammen:1986:SIC

[Mam86] Enno Mammen. The statistical information contained in additional observations. Annals of Statistics, 14(2):665–678, June 1986. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349945.

Mammen:1989:AID

[Mam89] Enno Mammen. Asymptotics with increasing dimension for robust regression with applications to the bootstrap. Annals of Statistics, 17(1):382–400, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347023.

Manski:1981:LDM

[Man81] Charles F. Manski. Learning and decision making when subjective probabilities have subjective domains. Annals of Statistics, 9(1):59–65, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345332.

Mandelbaum:1984:AAL

[Man84] Avi Mandelbaum. All admissible linear estimators of the mean of a Gaussian distribution on a Hilbert space. Annals of Statistics, 12 (4):1448–1466, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346803.

Marden:1981:ITC

[Mar81] John I. Marden. Invariant tests on covariance matrices. Annals of Statistics, 9(6):1258–1266, November 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345642.

Marden:1982:CIN

[Mar82a] John I. Marden. Combining independent noncentral chi squared or F tests. Annals of Statistics, 10(1):266–277, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345709.

Marden:1982:MCC

[Mar82b] John I. Marden. Minimal complete classes of tests of hypotheses with multivariate one-sided alternatives. Annals of Statistics, 10 (3):962–970, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345886.

Marden:1983:AIT

[Mar83a] John I. Marden. Admissibility of invariant tests in the general multivariate analysis of variance problem. Annals of Statistics, 11(4): 1086–1099, December 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346323.

Marron:1983:ORC

[Mar83b] James Stephen Marron. Optimal rates of convergence to Bayes risk in nonparametric discrimination. Annals of Statistics, 11(4): 1142–1155, December 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346328.

Martinsek:1983:SOA

[Mar83c] Adam T. Martinsek. Second order approximation to the risk of a sequential procedure. Annals of Statistics, 11(3):827–836, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.

aos/1176346249. See corrections [Mar86, Mar91].

Martinsek:1984:SDE

[Mar84] Adam T. Martinsek. Sequential determination of estimator as well as sample size. Annals of Statistics, 12(2):533–550, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346504.

Marden:1985:CIO

[Mar85a] John I. Marden. Combining independent one-sided noncentral t or normal mean tests. Annals of Statistics, 13(4):1535–1553, December 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349754.

Marron:1985:AES

[Mar85b] James Stephen Marron. An asymptotically efficient solution to the bandwidth problem of kernel density estimation. Annals of Statistics, 13(3):1011–1023, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349653.

Martinsek:1986:CSO

[Mar86] Adam T. Martinsek. Correction: “Second Order Approximation to the Risk of a Sequential Procedure”. Annals of Statistics, 14(1):359, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349864. See [Mar83c].

Marron:1987:CCV

[Mar87] J. S. Marron. A comparison of cross-validation techniques in density estimation. Annals of Statistics, 15(1):152–162, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350258.

Martinsek:1991:CSO

[Mar91] Adam T. Martinsek. Correction: “Second Order Approximation to the Risk of a Sequential Procedure”. Annals of Statistics, 19 (4):2283, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348404. See [Mar83c].

Masry:1980:DTS

[Mas80] E. Masry. Discrete-time spectral estimation of continuoustime processes — the orthogonal series method. Annals of Statistics, 8(5):1100–1109, September 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345147.

Mason:1981:ANL

[Mas81a] David M. Mason. Asymptotic normality of linear combinations of order statistics with a smooth score function. Annals of Statistics, 9(4):899–908, July 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345531.

Mason:1981:USB

[Mas81b] David M. Mason. On the use of a statistic based on sequential ranks to prove limit theorems for simple linear rank statistics. Annals of Statistics, 9(2):424–436, March 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345408.

Mason:1983:MCC

[Mas83] David M. Mason. A minimax criterion for choosing weight functions for L-estimates of location. Annals of Statistics, 11(1):317– 325, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346082.

Mathew:1984:NQU

[Mat84] Thomas Mathew. On nonnegative quadratic unbiased estimability of variance components. Annals of Statistics, 12(4):1566–1569, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346814.

Mauro:1985:CAK

[Mau85] David Mauro. A combinatoric approach to the Kaplan–Meier estimator. Annals of Statistics, 13(1):142–149, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346582.

McCullagh:1986:ILR

[MC86] P. McCullagh and D. R. Cox. Invariants and likelihood ratio statistics. Annals of Statistics, 14(4):1419–1430, December 1986. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350167.

Maatta:1987:CPI

[MC87] Jon M. Maatta and George Casella. Conditional properties of interval estimators of the normal variance. Annals of Statistics, 15 (4):1372–1388, December 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350599.

Morgan:1988:BDF

[MC88] John P. Morgan and I. M. Chakravarti. Block designs for first and second order neighbor correlations. Annals of Statistics, 16(3): 1206–1224, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350956.

McCullagh:1983:QLF

[McC83] Peter McCullagh. Quasi-likelihood functions. Annals of Statistics, 11(1):59–67, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346056.

McKeague:1984:SBE

[McK84] Ian W. McKeague. On the stability of Bayes estimators for Gaussian processes. Annals of Statistics, 12(4):1310–1323, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346794.

McKeague:1986:ESR

[McK86] Ian W. McKeague. Estimation for a semimartingale regression model using the method of sieves. Annals of Statistics, 14(2):579– 589, June 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349939.

Mara:1984:EBC

[MD84] M. K. Mara and J. J. Deely. Empirical Bayes with a changing prior. Annals of Statistics, 12(3):1071–1078, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346722.

Milasevic:1987:USM

[MD87] P. Milasevic and G. R. Ducharme. Uniqueness of the spatial median. Annals of Statistics, 15(3):1332–1333, September 1987. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350511.

Meyer:1982:TCT

[Mey82] Michael M. Meyer. Transforming contingency tables. Annals of Statistics, 10(4):1172–1181, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345982.

Macpherson:1983:CLS

[MF83] Brian D. Macpherson and Wayne A. Fuller. Consistency of the least squares estimator of the first order moving average parameter. Annals of Statistics, 11(1):326–329, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346083.

Matthews:1985:ASS

[MFP85] D. E. Matthews, V. T. Farewell, and R. Pyke. Asymptotic scorestatistic processes and tests for constant hazard against a changepoint alternative. Annals of Statistics, 13(2):583–591, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349540.

Muller-Funk:1989:ACR

[MFPW89] Ulrich Muller-Funk, Friedrich Pukelsheim, and Hermann Witting. On the attainment of the Cramér–Rao bound in L_{r}-differentiable families of distributions. Annals of Statistics, 17(4):1742–1748, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347392.

Meeden:1981:AFP

[MG81] Glen Meeden and Malay Ghosh. Admissibility in finite problems. Annals of Statistics, 9(4):846–852, July 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345524.

Meeden:1983:CBE

[MG83] Glen Meeden and Malay Ghosh. Choosing between experiments: Applications to finite population sampling. Annals of Statistics, 11(1):296–305, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346080.

Meeden:1989:AKM

[MGSV89] G. Meeden, M. Ghosh, C. Srinivasan, and S. Vardeman. The admissibility of the Kaplan–Meier and other maximum likelihood estimators in the presence of censoring. Annals of Statistics, 17(4): 1509–1531, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347379.

Meeden:1985:SAN

[MGV85] Glen Meeden, Malay Ghosh, and Stephen Vardeman. Some admissible nonparametric and related finite population sampling estimators. Annals of Statistics, 13(2):811–817, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349559.

Miescke:1981:MSP

[Mie81] Klaus J. Miescke. Γ-minimax selection procedures in simultaneous testing problems. Annals of Statistics, 9(1):215–220, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345351.

Mielniczuk:1986:SAP

[Mie86] Jan Mielniczuk. Some asymptotic properties of kernel estimators of a density function in case of censored data. Annals of Statistics, 14(2):766–773, June 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349954.

Miller:1980:NEN

[Mil80] Rupert G. Miller. Note: Editorial note on Pao-Lu Hsu. Annals of Statistics, 8(2):456, March 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344966.

Millar:1982:OEG

[Mil82] P. W. Millar. Optimal estimation of a general regression function. Annals of Statistics, 10(3):717–740, September 1982. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345867.

Miller:1985:DPP

[Mil85] Rupert G. Miller. Discussion: Projection pursuit. Annals of Statistics, 13(2):510–513, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349530.

Miller:1989:EPB

[Mil89] Stephen M. Miller. Empirical processes based upon residuals from errors-in-variables regressions. Annals of Statistics, 17(1):282–292, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347016.

Mittal:1981:SSM

[Mit81] Yashaswini Mittal. Smoothing of samples for maxima. Annals of Statistics, 9(1):66–77, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345333.

Miller:1986:DIF

[ML86] Robert B. Miller and Jae June Lee. Discussion: Influence functionals for time series. Annals of Statistics, 14(3):827–829, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350031.

Moolgavkar:1984:GAN

[MLV84] Suresh H. Moolgavkar, Edward D. Lustbader, and David J. Venzon. A geometric approach to nonlinear regression diagnostics with applications to matched case-control studies. Annals of Statistics, 12(3):816–826, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346704.

Mihalko:1980:CST

[MM80] Daniel P. Mihalko and David S. Moore. Chi-square tests of fit for type II censored data. Annals of Statistics, 8(3):625–644, May 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345013.

Mazloum:1987:USB

[MM87] Reda Mazloum and Glen Meeden. Using the stepwise Bayes technique to choose between experiments. Annals of Statistics, 15 (1):269–277, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350265.

Majumdar:1983:OIB

[MN83] Dibyen Majumdar and William I. Notz. Optimal incomplete block designs for comparing treatments with a control. Annals of Statistics, 11(1):258–266, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346076.

Mitchell-Olds:1986:DJB

[MO86] Thomas Mitchell-Olds. Discussion: Jackknife, bootstrap and other resampling methods in regression analysis. Annals of Statistics, 14 (4):1316–1318, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350151.

Moore:1982:EDC

[Moo82] David S. Moore. The effect of dependence on chi squared tests of fit. Annals of Statistics, 10(4):1163–1171, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345981.

Moore:1984:ECS

[Moo84] Marc Moore. On the estimation of a convex set. Annals of Statistics, 12(3):1090–1099, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346725.

Morris:1982:NEF

[Mor82] Carl N. Morris. Natural exponential families with quadratic variance functions. Annals of Statistics, 10(1):65–80, March 1982. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345690.

Morris:1983:DCI

[Mor83a] Carl N. Morris. Discussion: Construction of improved estimators in multiparameter estimation for discrete exponential families. Annals of Statistics, 11(2):372–374, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346146. See [GHT83].

Morris:1983:NEF

[Mor83b] Carl N. Morris. Natural exponential families with quadratic variance functions: Statistical theory. Annals of Statistics, 11(2):515– 529, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346158.

Morgenthaler:1986:ACL

[Mor86] Stephan Morgenthaler. Asymptotics for configural location estimators. Annals of Statistics, 14(1):174–187, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349848.

Moustakides:1986:OST

[Mou86] George V. Moustakides. Optimal stopping times for detecting changes in distributions. Annals of Statistics, 14(4):1379–1387, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350164.

Marden:1980:ITM

[MP80] John Marden and Michael D. Perlman. Invariant tests for means with covariates. Annals of Statistics, 8(1):25–63, January 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344890.

Mitra:1984:NSR

[MP84] S. K. Mitra and P. K. Pathak. The nature of simple random sampling. Annals of Statistics, 12(4):1536–1542, December 1984. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346810.

McCullagh:1985:DTI

[MP85] Peter McCullagh and Daryl Pregibon. Discussion: Testing for independence in a two-way table: New interpretations of the chi-square statistic. Annals of Statistics, 13(3):898–900, September 1985. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349641.

Marron:1987:AOB

[MP87a] J. S. Marron and W. J. Padgett. Asymptotically optimal bandwidth selection for kernel density estimators from randomly rightcensored samples. Annals of Statistics, 15(4):1520–1535, December 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350607.

McCullagh:1987:SDE

[MP87b] Peter McCullagh and Daryl Pregibon. k-statistics and dispersion effects in regression. Annals of Statistics, 15(1):202–219, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350261.

Mandelbaum:1987:CSC

[MR87a] Avi Mandelbaum and Ludger Ruschendorf. Complete and symmetrically complete families of distributions. Annals of Statistics, 15(3):1229–1244, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350502.

Melkman:1987:MEM

[MR87b] Avraham A. Melkman and Ya’acov Ritov. Minimax estimation of the mean of a general distribution when the parameter space is restricted. Annals of Statistics, 15(1):432–442, March 1987. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350278.

Major:1988:SEE

[MR88] P. Major and L. Rejto. Strong embedding of the estimator of the distribution function under random censorship. Annals of Statistics, 16(3):1113–1132, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350949.

Mukerjee:1986:PIE

[MRW86] Hari Mukerjee, Tim Robertson, and F. T. Wright. A probability inequality for elliptically contoured densities with applications in order restricted inference. Annals of Statistics, 14(4): 1544–1554, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350175.

Mason:1983:MKS

[MS83] David M. Mason and John H. Schuenemeyer. A modified Kolmogorov–Smirnov test sensitive to tail alternatives. Annals of Statistics, 11(3):933–946, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:/

/projecteuclid.org/euclid.aos/1176346259. See correction

[MS92].

Mandelbaum:1987:AT

[MS87a] Avi Mandelbaum and L. A. Shepp. Admissibility as a touchstone. Annals of Statistics, 15(1):252–268, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350264.

Muller:1987:EHR

[MS87b] Hans-Georg Muller and Ulrich Stadtmuller. Estimation of heteroscedasticity in regression analysis. Annals of Statistics, 15 (2):610–625, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350364.

Muller:1987:VBK

[MS87c] Hans-Georg Muller and Ulrich Stadtmuller. Variable bandwidth kernel estimators of regression curves. Annals of Statistics, 15 (1):182–201, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350260.

Mathew:1988:OTU

[MS88] Thomas Mathew and Bimal Kumar Sinha. Optimum tests in unbalanced two-way models without interaction. Annals of Statistics, 16(4):1727–1740, December 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176351065.

Mason:1992:CMK

[MS92] David M. Mason and John H. Schuenemeyer. Correction: “A Modified Kolmogorov–Smirnov Test Sensitive to Tail Alternatives”. Annals of Statistics, 20(1):620–621, March 1992. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176348549. See

[MS83].

Makelainen:1981:EUM

[MSS81] Timo Makelainen, Klaus Schmidt, and George P. H. Styan. On the existence and uniqueness of the maximum likelihood estimate of a vector-valued parameter in fixed-size samples. Annals of Statistics, 9(4):758–767, July 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345516.

Mandelbaum:1984:IPS

[MT84] Avi Mandelbaum and Murad S. Taqqu. Invariance principle for symmetric statistics. Annals of Statistics, 12(2):483–496, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346501.

Mocks:1984:THN

[MTG84] Joachim Mocks, Pham Dinh Tuan, and Theo Gasser. Testing for homogeneity of noisy signals evoked by repeated stimuli. Annals of Statistics, 12(1):193–209, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346401.

Mukerjee:1981:SAO

[Muk81] H. G. Mukerjee. A stochastic approximation by observations on a discrete lattice using isotonic regression. Annals of Statistics, 9(5): 1020–1025, September 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345581.

Mukerjee:1988:MNR

[Muk88] Hari Mukerjee. Monotone nonparametric regression. Annals of Statistics, 16(2):741–750, June 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350832.

Muller:1984:SOK

[Mul84] Hans-Georg Muller. Smooth optimum kernel estimators of densities, regression curves and modes. Annals of Statistics, 12 (2):766–774, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346523.

Muller:1989:ANP

[Mul89] Hans-Georg Muller. Adaptive nonparametric peak estimation. Annals of Statistics, 17(3):1053–1069, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347255.

Moolgavkar:1987:CRC

[MV87] Suresh H. Moolgavkar and David J. Venzon. Confidence regions in curved exponential families: Application to matched case-control and survival studies with general relative risk function. Annals of Statistics, 15(1):346–359, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350270.

Mallik:1984:BBR

[MY84] Ashim Mallik and Yi-Ching Yao. Bounds for the Bayes risk for testing sequentially the sign of the drift parameter of a Wiener process. Annals of Statistics, 12(3):1117–1123, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346729.

Martin:1986:IFT

[MY86a] R. Douglas Martin and Victor J. Yohai. Influence functionals for time series. Annals of Statistics, 14(3):781–818, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350027.

Martin:1986:RIF

[MY86b] R. Douglas Martin and Victor J. Yohai. Rejoinder: Influence functionals for time series. Annals of Statistics, 14(3):840–855, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350037.

Martin:1989:MMB

[MYZ89] R. D. Martin, V. J. Yohai, and R. H. Zamar. Min-max bias robust regression. Annals of Statistics, 17(4):1608–1630, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347384.

Nagaraja:1982:SNL

[Nag82] H. N. Nagaraja. Some nondegenerate limit laws for the selection differential. Annals of Statistics, 10(4):1306–1310, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345997.

Nair:1980:VDG

[Nai80] K. Aiyappan Nair. Variance and distribution of the Graybill–Deal estimator of the common mean of two normal populations. Annals of Statistics, 8(1):212–216, January 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344904.

Naiman:1984:AWO

[Nai84a] Daniel Q. Naiman. Average width optimality of simultaneous confidence bounds. Annals of Statistics, 12(4):1199–1214, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346787.

Naiman:1984:OSC

[Nai84b] Daniel Q. Naiman. Optimal simultaneous confidence bounds. Annals of Statistics, 12(2):702–715, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346516.

Naiman:1986:CCB

[Nai86] Daniel Q. Naiman. Conservative confidence bands in curvilinear regression. Annals of Statistics, 14(3):896–906, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350040.

Navidi:1989:EEB

[Nav89] William Navidi. Edgeworth expansions for bootstrapping regression models. Annals of Statistics, 17(4):1472–1478, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347375.

Nychka:1989:CRR

[NC89] Douglas W. Nychka and Dennis D. Cox. Convergence rates for regularized solutions of integral equations from discrete noisy data.

Annals of Statistics, 17(2):556–572, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347125.

Neill:1988:TLF

[Nei88] James W. Neill. Testing for lack of fit in nonlinear regression. Annals of Statistics, 16(2):733–740, June 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350831.

Nelson:1980:NSC

[Nel80] Paul I. Nelson. A note on strong consistency of least squares estimators in regression models with martingale difference errors. Annals of Statistics, 8(5):1057–1064, September 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345142.

Nelder:1987:DWA

[Nel87] J. A. Nelder. Discussion: What is an analysis of variance? Annals of Statistics, 15(3):930–931, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350480.

Neuhaus:1982:CNT

[Neu82] Georg Neuhaus. H_{0}-contiguity in nonparametric testing problems and sample Pitman efficiency. Annals of Statistics, 10(2):575–582, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345798.

Neuhaus:1987:LAL

[Neu87] Georg Neuhaus. Local asymptotics for linear rank statistics with estimated score functions. Annals of Statistics, 15(2):491–512, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350357.

Neuhaus:1988:ALA

[Neu88] Georg Neuhaus. Addendum to: “Local Asymptotics for Linear Rank Statistics with Estimated Score Functions”. Annals of Statistics, 16(3):1342–1343, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350967.

Neyman:1981:EPA

[Ney81] Jerzy Neyman. Egon S. Pearson (August 11, 1895–June 12, 1980). An appreciation. Annals of Statistics, 9(1):1–2, January 1981. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345328.

Niederhausen:1981:SPC

[Nie81] Heinrich Niederhausen. Sheffer polynomials for computing exact Kolmogorov–Smirnov and Rényi type distributions. Annals of Statistics, 9(5):923–944, September 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345574.

Niederhausen:1983:SPC

[Nie83] Heinrich Niederhausen. Sheffer polynomials for computing Takacs’s goodness-of-fit distributions. Annals of Statistics, 11 (2):600–606, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346165.

Nishii:1984:APC

[Nis84] Ryuei Nishii. Asymptotic properties of criteria for selection of variables in multiple regression. Annals of Statistics, 12(2):758–765, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346522.

Neill:1985:TLR

[NJ85] James W. Neill and Dallas E. Johnson. Testing linear regression function adequacy without replication. Annals of Statistics, 13(4): 1482–1489, December 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349749.

Nogami:1988:CRE

[Nog88] Yoshiko Nogami. Convergence rates for empirical Bayes estimation in the uniform U(0, θ) distribution. Annals of Statistics, 16(3): 1335–1341, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350966.

Novic:1980:BSE

[Nov80] Bradley Novic. Bayes sequential estimation of a Poisson rate: A discrete time approach. Annals of Statistics, 8(4):840–844, July 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345076.

Nolan:1987:PRC

[NP87] Deborah Nolan and David Pollard. U-processes: Rates of convergence. Annals of Statistics, 15(2):780–799, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350374.

Natarajan:1985:TSS

[NS85] Jayalakshmi Natarajan and William E. Strawderman. Twostage sequential estimation of a multivariate normal mean under quadratic loss. Annals of Statistics, 13(4):1509–1522, December 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349752.

Nabeya:1988:ATT

[NT88] Seiji Nabeya and Katsuto Tanaka. Asymptotic theory of a test for the constancy of regression coefficients against the random walk alternative. Annals of Statistics, 16(1):218–235, March 1988. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350701.

Nussbaum:1985:SSR

[Nus85] Michael Nussbaum. Spline smoothing in regression models and asymptotic efficiency in L_{2}. Annals of Statistics, 13(3):984–997, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349651.

Oberhofer:1982:CNR

[Obe82] Walter Oberhofer. The consistency of nonlinear regression minimizing the L_{1}-norm. Annals of Statistics, 10(1):316–319, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345716.

Olshen:1989:GAB

[OBWS89] Richard A. Olshen, Edmund N. Biden, Marilynn P. Wyatt, and David H. Sutherland. Gait analysis and the bootstrap. Annals of Statistics, 17(4):1419–1440, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347372.

OReilly:1980:CCS

[OD80] Federico J. O’Reilly and Joaquin Diaz. Correction to “On a Criterion for Simultaneous Extrapolation in Nonfull Rank Normal Regression”. Annals of Statistics, 8(5):1177–1178, September 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345158. See

[O’R76].

Oehlert:1985:RAM

[Oeh85] Gary W. Oehlert. The random average mode estimator. Annals of Statistics, 13(4):1418–1431, December 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349745.

Olshen:1986:DJB

[Ols86] Richard A. Olshen. Discussion: Jackknife, bootstrap and other resampling methods in regression analysis. Annals of Statistics, 14 (4):1318–1320, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350152.

Oman:1988:CRM

[Oma88] Samuel D. Oman. Confidence regions in multivariate calibration. Annals of Statistics, 16(1):174–187, March 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350698.

OReilly:1976:CSE

[O’R76] Federico J. O’Reilly. On a criterion for simultaneous extrapolation in nonfull rank normal regression. Annals of Statistics, 4(3):625–628, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343469. See correction [OD80].

Olkin:1981:NCM

[OT81] Ingram Olkin and Jack L. Tomsky. A new class of multivariate tests based on the union-intersection principle. Annals of Statistics, 9(4):792–802, July 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345519.

Parke:1986:PML

[Par86] William R. Parke. Pseudo maximum likelihood estimation: The asymptotic distribution. Annals of Statistics, 14(1):355–357, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349862.

Pons:1988:CPR

[PdT88] O. Pons and E. de Turckheim. Cox’s periodic regression model. Annals of Statistics, 16(2):678–693, June 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350828.

Pedersen:1981:CEH

[Ped81] Bo V. Pedersen. A comparison of the Efron–Hinkley ancillary and the likelihood ratio ancillary in a particular example. Annals of Statistics, 9(6):1328–1333, November 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345649.

Perlman:1980:ULR

[Per80] Michael D. Perlman. Unbiasedness of the likelihood ratio tests for equality of several covariance matrices and equality of several multivariate normal populations. Annals of Statistics, 8 (2):247–263, March 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176344951.

Pesotchinsky:1982:ORD

[Pes82] L. Pesotchinsky. Optimal robust designs: Linear regression in R^{k}. Annals of Statistics, 10(2):511–525, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345792.

Petruccelli:1980:BCP

[Pet80] Joseph D. Petruccelli. On a best choice problem with partial information. Annals of Statistics, 8(5):1171–1174, September 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345156.

Pfaff:1982:QCQ

[Pfa82] Thomas Pfaff. Quick consistency of quasi maximum likelihood estimators. Annals of Statistics, 10(3):990–1005, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345889.

Phadia:1980:NEB

[Pha80] E. G. Phadia. A note on empirical Bayes estimation of a distribution function based on censored data. Annals of Statistics, 8 (1):226–229, January 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176344907.

Pierce:1982:AES

[Pie82] Donald A. Pierce. The asymptotic effect of substituting estimators for parameters in certain types of statistics. Annals of Statistics, 10(2):475–478, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345788.

Piegorsch:1985:AOC

[Pie85a] Walter W. Piegorsch. Admissible and optimal confidence bands in simple linear regression. Annals of Statistics, 13(2):801–810, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349558.

Pierce:1985:DTI

[Pie85b] Donald A. Pierce. Discussion: Testing for independence in a twoway table: New interpretations of the chi-square statistic. Annals of Statistics, 13(3):901–902, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349642.

Portnoy:1989:AEL

[PK89] Stephen Portnoy and Roger Koenker. Adaptive L-estimation for linear models. Annals of Statistics, 17(1):362–381, March 1989.

CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347022. See

correction [PK90].

Portnoy:1990:CAE

[PK90] Stephen Portnoy and Roger Koenker. Correction: “Adaptive L-Estimation for Linear Models”. Annals of Statistics, 18(2):986, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347638. See [PK89].

Plackett:1985:DTI

[Pla85] R. L. Plackett. Discussion: Testing for independence in a twoway table: New interpretations of the chi-square statistic. Annals of Statistics, 13(3):903, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349643.

Perlman:1980:UIT

[PO80] Michael D. Perlman and Ingram Olkin. Unbiasedness of invariant tests for manova and other multivariate problems. Annals of Statistics, 8(6):1326–1341, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345204.

Pollard:1981:SCM

[Pol81] David Pollard. Strong consistency of K-means clustering. Annals of Statistics, 9(1):135–140, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345339.

Pollak:1985:ODC

[Pol85] Moshe Pollak. Optimal detection of a change in distribution. Annals of Statistics, 13(1):206–227, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346587.

Pollak:1986:AFP

[Pol86] Moshe Pollak. On the asymptotic formula for the probability of a Type I error of mixture type power one tests. Annals of Statistics, 14(3):1012–1029, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350047.

Pollak:1987:ARL

[Pol87] Moshe Pollak. Average run lengths of an optimal method of detecting a change in distribution. Annals of Statistics, 15(2):749–779, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350373.

Poor:1986:DIF

[Poo86] H. Vincent Poor. Discussion: Influence functionals for time series. Annals of Statistics, 14(3):829–831, September 1986. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350032.

Portnoy:1984:ABE

[Por84] Stephen Portnoy. Asymptotic behavior of M-estimators of p regression parameters when p^{2}/n is large. I. Consistency. Annals of Statistics, 12(4):1298–1309, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346793.

Portnoy:1985:ABE

[Por85] Stephen Portnoy. Asymptotic behavior of M estimators of p regression parameters when p^{2}/n is large; II. normal approximation. Annals of Statistics, 13(4):1403–1417, December 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349744. See cor-

rection [Por91].

Portnoy:1986:ABE

[Por86] Stephen Portnoy. Asymptotic behavior of the empiric distribution of M-estimated residuals from a regression model with many parameters. Annals of Statistics, 14(3):1152–1170, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350056.

Portnoy:1988:ABL

[Por88] Stephen Portnoy. Asymptotic behavior of likelihood methods for exponential families when the number of parameters tends to infinity. Annals of Statistics, 16(1):356–366, March 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350710.

Porteous:1989:SIR

[Por89] B. T. Porteous. Stochastic inequalities relating a class of loglikelihood ratio statistics to their asymptotic χ^{2} distribution. Annals of Statistics, 17(4):1723–1734, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347390.

Portnoy:1991:CAB

[Por91] Stephen Portnoy. Correction: “Asymptotic Behavior of M Estimators of p Regression Parameters when p^{2}/n is Large: II. Normal Approximation”. Annals of Statistics, 19(4):2282, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176348403. See

[Por85].

Poskitt:1986:BPI

[Pos86] D. S. Poskitt. A Bayes procedure for the identification of univariate time series models. Annals of Statistics, 14(2):502–516, June 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349935.

Potscher:1983:OEA

[Pot83] B. M. Potscher. Order estimation in ARMA-models by Lagrangian multiplier tests. Annals of Statistics, 11(3):872–885, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.

aos/1176346253. See corrections [Pot84].

Potscher:1984:COE

[Pot84] B. M. Potscher. Corrections: “Order Estimation in ARMA-Models by Lagrangian Multiplier Tests”. Annals of Statistics, 12(2):785, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346529. See [Pot83].

Potscher:1989:MSU

[Pot89] B. M. Potscher. Model selection under nonstationarity: Autoregressive models and stochastic linear regression models. Annals of Statistics, 17(3):1257–1274, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347267.

Pesotan:1981:FRI

[PR81] H. Pesotan and B. L. Raktoe. Further results on invariance and randomization in fractional replication. Annals of Statistics, 9 (2):418–423, March 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345407.

Praagman:1988:BER

[Pra88] Jaap Praagman. Bahadur efficiency of rank tests for the changepoint problem. Annals of Statistics, 16(1):198–217, March 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350700.

Pregibon:1981:LRD

[Pre81] Daryl Pregibon. Logistic regression diagnostics. Annals of Statistics, 9(4):705–724, July 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345513.

Pregibon:1984:RPM

[Pre84] Daryl Pregibon. Review: P. McCullagh, J. A. Nelder, Generalized Linear Models. Annals of Statistics, 12(4):1589–1596, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346819.

Prentice:1983:ADT

[PS83] Ross L. Prentice and Steven G. Self. Asymptotic distribution theory for Cox-type regression models with general relative risk form. Annals of Statistics, 11(3):804–813, September 1983. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346247.

Poskitt:1981:ATL

[PT81] D. S. Poskitt and A. R. Tremayne. An approach to testing linear times series models. Annals of Statistics, 9(5):974–986, September 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345577.

Poskitt:1982:DTM

[PT82] D. S. Poskitt and A. R. Tremayne. Diagnostic tests for multiple time series models. Annals of Statistics, 10(1):114–120, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345694.

Pukelsheim:1983:GDL

[PT83] F. Pukelsheim and D. M. Titterington. General differential and Lagrangian theory for optimal experimental design. Annals of Statistics, 11(4):1060–1068, December 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346321.

Pukelsheim:1981:EUN

[Puk81] Friedrich Pukelsheim. On the existence of unbiased nonnegative estimates of variance convariance components. Annals of Statistics, 9(2):293–299, March 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345395.

Pukelsheim:1984:NNT

[Puk84] Friedrich Pukelsheim. A note on nonparametric trend conformity. Annals of Statistics, 12(2):775–777, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346524.

Phadia:1980:NCR

[PV80] E. G. Phadia and J. Van Ryzin. A note on convergence rates for the product limit estimator. Annals of Statistics, 8(3):673–678, May 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345017.

Quine:1985:ECS

[QR85] M. P. Quine and J. Robinson. Efficiencies of chi-square and likelihood ratio goodness-of-fit tests. Annals of Statistics, 13 (2):727–742, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349550.

Quang:1985:RST

[Qua85] Pham Xuan Quang. Robust sequential testing. Annals of Statistics, 13(2):638–649, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349544.

Ramkaran:1983:RST

[Ram83] Ramkaran. The robustness of Stein’s two-stage procedure. Annals of Statistics, 11(4):1251–1256, December 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346338.

Randles:1982:ANS

[Ran82] Ronald H. Randles. On the asymptotic normality of statistics with estimated parameters. Annals of Statistics, 10(2):462–474, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345787.

Rao:1978:ADE

[Rao78] M. M. Rao. Asymptotic distribution of an estimator of the boundary parameter of an unstable process. Annals of Statistics, 6 (1):185–190, January 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176344077. See correction [Rao80].

Rao:1980:CAD

[Rao80] M. M. Rao. Correction to “Asymptotic Distribution of an Estimator of the Boundary Parameter of an Unstable Process”. Annals of Statistics, 8(6):1403, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345217. See [Rao78].

Rasmussen:1980:BAP

[Ras80] Shelley L. Rasmussen. A Bayesian approach to a problem in sequential estimation. Annals of Statistics, 8(6):1229–1243, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345196.

Richardson:1986:CEN

[RB86] G. D. Richardson and B. B. Bhattacharyya. Consistent estimators in nonlinear regression for a noncompact parameter space. Annals of Statistics, 14(4):1591–1596, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350179.

Redner:1981:NCM

[Red81] Richard Redner. Note on the consistency of the maximum likelihood estimate for nonidentifiable distributions. Annals of Statistics, 9(1):225–228, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345353.

Reeds:1985:ANR

[Ree85] James A. Reeds. Asymptotic number of roots of Cauchy location likelihood equations. Annals of Statistics, 13(2):775–784, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349554.

Regazzini:1987:FCS

[Reg87] Eugenio Regazzini. De Finetti’s coherence and statistical inference. Annals of Statistics, 15(2):845–864, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350379.

Reiss:1980:EQC

[Rei80] R.-D. Reiss. Estimation of quantiles in certain nonparametric models. Annals of Statistics, 8(1):87–105, January 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344893.

Reid:1981:IFC

[Rei81] Nancy Reid. Influence functions for censored data. Annals of Statistics, 9(1):78–92, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345334.

Ramlau-Hansen:1983:SCP

[RH83] Henrik Ramlau-Hansen. Smoothing counting process intensities by means of kernel functions. Annals of Statistics, 11(2):453–466, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346152.

Rice:1984:BCN

[Ric84] John Rice. Bandwidth choice for nonparametric regression. Annals of Statistics, 12(4):1215–1230, December 1984. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346788.

Rieder:1980:EDR

[Rie80] Helmut Rieder. Estimates derived from robust tests. Annals of Statistics, 8(1):106–115, January 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344894.

Rieder:1981:LAM

[Rie81a] Helmut Rieder. On local asymptotic minimaxity and admissibility in robust estimation. Annals of Statistics, 9(2):266–277, March 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345393.

Rieder:1981:ROT

[Rie81b] Helmut Rieder. Robustness of one-and two-sample rank tests against Gross errors. Annals of Statistics, 9(2):245–265, March 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176345392.

Rieder:1982:QRR

[Rie82] Helmut Rieder. Qualitative robustness of rank tests. Annals of Statistics, 10(1):205–211, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345703.

Rissanen:1983:UPI

[Ris83] Jorma Rissanen. A universal prior for integers and estimation by minimum description length. Annals of Statistics, 11(2):416–431, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346150.

Rissanen:1986:SCM

[Ris86] Jorma Rissanen. Stochastic complexity and modeling. Annals of Statistics, 14(3):1080–1100, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350051.

Ritov:1985:RBD

[Rit85] Ya’acov Ritov. Robust Bayes decision procedures: Gross error in the data distribution. Annals of Statistics, 13(2):626–637, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349543.

Rivest:1984:IMS

[Riv84a] Louis-Paul Rivest. On the information matrix for symmetric distributions on the hypersphere. Annals of Statistics, 12(3): 1085–1089, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346724.

Rivest:1984:SDD

[Riv84b] Louis-Paul Rivest. Symmetric distributions for dependent unit vectors. Annals of Statistics, 12(3):1050–1057, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346720.

Rivest:1989:SRC

[Riv89] Louis-Paul Rivest. Spherical regression for concentrated Fisher– von Mises distributions. Annals of Statistics, 17(1):307–317, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347018.

Robinson:1976:PSB

[Rob76] G. K. Robinson. Properties of Student’s t and of the Behrens– Fisher solution to the two means problem. Annals of Statistics, 4 (5):963–971, September 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343594. See correction [Rob82a].

Robinson:1980:AEP

[Rob80] J. Robinson. An asymptotic expansion for permutation tests with several samples. Annals of Statistics, 8(4):851–864, July 1980. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345078.

Robinson:1982:CPS

[Rob82a] G. K. Robinson. Corrections: “Properties of Student’s t and of the Behrens–Fisher Solution to the Two Means Problem”. Annals of Statistics, 10(1):321, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345718. See [Rob76].

Robinson:1982:ATS

[Rob82b] P. M. Robinson. Analysis of time series from mixed distributions. Annals of Statistics, 10(3):915–925, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345881.

Robbins:1983:STE

[Rob83] Herbert Robbins. Some thoughts on empirical Bayes estimation. Annals of Statistics, 11(3):713–723, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346239.

Robinson:1986:DIF

[Rob86] P. M. Robinson. Discussion: Influence functionals for time series. Annals of Statistics, 14(3):832–834, September 1986. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350033.

Robinson:1988:DTC

[Rob88] J. Robinson. Discussion: Theoretical comparison of bootstrap confidence intervals. Annals of Statistics, 16(3):976–977, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350941.

Romano:1988:WCO

[Rom88] Joseph P. Romano. On weak convergence and optimality of kernel density estimates of the mode. Annals of Statistics, 16 (2):629–647, June 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350824.

Romano:1989:BRT

[Rom89] Joseph P. Romano. Bootstrap and randomization tests of some nonparametric hypotheses. Annals of Statistics, 17(1):141–159, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347007.

Rosenblatt:1975:QMD

[Ros75] M. Rosenblatt. A quadratic measure of deviation of twodimensional density estimates and a test of independence. Annals of Statistics, 3(1):1–14, January 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:/

/projecteuclid.org/euclid.aos/1176342996. See correction

[Ros82].

Rosenblatt:1982:CQM

[Ros82] M. Rosenblatt. Corrections: “A Quadratic Measure of Deviation of Two-Dimension Density Estimates and a Test of Independence”. Annals of Statistics, 10(2):646, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345809. See [Ros75].

Rosenbaum:1989:PTH

[Ros89] Paul R. Rosenbaum. On permutation tests for hidden biases in observational studies: An application of Holley’s inequality to the Savage lattice. Annals of Statistics, 17(2):643–653, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347131.

Rothe:1981:SPA

[Rot81] Gunter Rothe. Some properties of the asymptotic relative Pitman efficiency. Annals of Statistics, 9(3):663–669, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345470.

Roth:1985:VKM

[Rot85] Arthur J. Roth. Variance of the Kaplan–Meier estimator and its quantiles under certain fixed censoring models. Annals of Statistics, 13(3):1230–1238, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349667.

Rinott:1980:SOI

[RP80] Yosef Rinott and Moshe Pollak. A stochastic ordering induced by a concept of positive dependence and monotonicity of asymptotic test sizes. Annals of Statistics, 8(1):190–198, January 1980. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344901.

Rao:1986:DJB

[RP86] J. N. K. Rao and N. G. N. Prasad. Discussion: Jackknife, bootstrap and other resampling methods in regression analysis. Annals of Statistics, 14(4):1320–1322, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350153.

Ralescu:1981:CNA

[RR81] Dan Ralescu and Stefan Ralescu. A class of nonlinear admissible estimators in the one-parameter exponential family. Annals of Statistics, 9(1):177–183, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345344.

Rice:1983:SSR

[RR83] John Rice and Murray Rosenblatt. Smoothing splines: Regression, derivatives and deconvolution. Annals of Statistics, 11 (1):141–156, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346065.

Rao:1984:CST

[RS84] J. N. K. Rao and A. J. Scott. On chi-squared tests for multiway contingency tables with Cell proportions estimated from survey data. Annals of Statistics, 12(1):46–60, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346391.

Rosenblatt:1985:JRB

[RS85] M. Rosenblatt and F. J. Samaniego. Julius R. Blum 1922–1982. Annals of Statistics, 13(1):1–9, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346575.

Rao:1987:SAC

[RS87] J. N. K. Rao and A. J. Scott. On simple adjustments to chisquare tests with sample survey data. Annals of Statistics, 15 (1):385–397, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350273.

Rinott:1987:COS

[RSC87] Yosef Rinott and Ester Samuel-Cahn. Comparisons of optimal stopping values and prophet inequalities for negatively dependent random variables. Annals of Statistics, 15(4):1482–1490, December 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350605.

Robertson:1984:GKM

[RU84] James B. Robertson and V. R. R. Uppuluri. A generalized Kaplan– Meier estimator. Annals of Statistics, 12(1):366–371, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346414.

Rubin:1981:BB

[Rub81] Donald B. Rubin. The Bayesian bootstrap. Annals of Statistics, 9 (1):130–134, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345338.

Rubin:1984:BJR

[Rub84] Donald B. Rubin. Bayesianly justifiable and relevant frequency calculations for the applies statistician. Annals of Statistics, 12(4): 1151–1172, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346785.

Rukhin:1982:APM

[Ruk82] Andrew L. Rukhin. Adaptive procedures in multiple decision problems and hypothesis testing. Annals of Statistics, 10(4): 1148–1162, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345980.

Rukhin:1983:CRE

[Ruk83] Andrew L. Rukhin. Convergence rates of estimators of a finite parameter: How small can error probabilities be? Annals of Statistics, 11(1):202–207, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346070.

Rukhin:1985:ERN

[Ruk85] Andrew Rukhin. Estimating a ratio of normal parameters. Annals of Statistics, 13(4):1616–1624, December 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349758.

Rukhin:1986:AMR

[Ruk86] Andrew L. Rukhin. Admissibility and minaxity results in the estimation problem of exponential quantiles. Annals of Statistics, 14(1):220–237, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349851.

Rukhin:1988:LFL

[Ruk88] Andrew L. Rukhin. Loss functions for loss estimation. Annals of Statistics, 16(3):1262–1269, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350960.

Ruppert:1981:SAI

[Rup81] David Ruppert. Stochastic approximation of an implicitly defined function. Annals of Statistics, 9(3):555–566, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345459.

Ruppert:1985:NRV

[Rup85] David Ruppert. A Newton–Raphson version of the multivariate Robbins–Monro procedure. Annals of Statistics, 13(1):236–245, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346589.

Ruschendorf:1976:ADM

[Rus76] Ludger Ruschendorf. Asymptotic distributions of multivariate rank order statistics. Annals of Statistics, 4(5):912–923, September 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.

aos/1176343588. See correction [Rus82].

Ruschendorf:1982:CAD

[Rus82] L. Ruschendorf. Corrections: “Asymptotic Distributions of Multivariate Rank Order Statistics”. Annals of Statistics, 10(4):1311, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345998. See [Rus76].

Rubin:1980:ADS

[RV80] H. Rubin and R. A. Vitale. Asymptotic distribution of symmetric statistics. Annals of Statistics, 8(1):165–170, January 1980. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344898.

Robertson:1980:AOR

[RW80] Tim Robertson and F. T. Wright. Algorithms in order restricted statistical inference and the Cauchy mean value property. Annals of Statistics, 8(3):645–651, May 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345014.

Robertson:1981:LRT

[RW81] Tim Robertson and F. T. Wright. Likelihood ratio tests for and against a stochastic ordering between multinomial populations. Annals of Statistics, 9(6):1248–1257, November 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345641.

Robertson:1982:BMD

[RW82a] Tim Robertson and F. T. Wright. Bounds on mixtures of distributions arising in order restricted inference. Annals of Statistics, 10(1):302–306, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345713.

Robertson:1982:MCP

[RW82b] Tim Robertson and F. T. Wright. On measuring the conformity of a parameter set to a trend, with applications. Annals of Statistics, 10(4):1234–1245, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345988.

Saxena:1982:ENC

[SA82] K. M. Lal Saxena and Khursheed Alam. Estimation of the noncentrality parameter of a chi squared distribution. Annals of Statistics, 10(3):1012–1016, September 1982. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345892.

Sadooghi-Alvandi:1986:AEB

[SA86] S. M. Sadooghi-Alvandi. Admissible estimation of the binomial parameter n. Annals of Statistics, 14(4):1634–1641, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350185.

Sacks:1984:JCK

[Sac84] Jerome Sacks. Jack Carl Kiefer 1924–1981. Annals of Statistics, 12(2):403–405, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346494.

Sager:1982:NML

[Sag82] Thomas W. Sager. Nonparametric maximum likelihood estimation of spatial patterns. Annals of Statistics, 10(4):1125–1136, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345978.

Sakai:1983:CMC

[Sak83] Hideaki Sakai. Covariance matrices characterization by a set of scalar partial autocorrelation coefficients. Annals of Statistics, 11 (1):337–340, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346085.

Samn:1980:RDC

[Sam80] S. Samn. Ratios of off-diagonal C-Wishart. Annals of Statistics, 8 (1):199–204, January 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176344902.

Samarov:1987:RSR

[Sam87] Alexander M. Samarov. Robust spectral regression. Annals of Statistics, 15(1):99–111, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350255.

Sauermann:1989:BML

[Sau89] Wilhelm Sauermann. Bootstrapping the maximum likelihood estimator in high-dimensional log-linear models. Annals of Statistics, 17(3):1198–1216, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347264.

Seely:1980:EPL

[SB80] Justus Seely and David Birkes. Estimability in partitioned linear models. Annals of Statistics, 8(2):399–406, March 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344960.

Sivaganesan:1989:RPM

[SB89] S. Sivaganesan and James O. Berger. Ranges of posterior measures for priors with unimodal contaminations. Annals of Statistics, 17(2):868–889, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347148.

Chang:1979:DOC

[sC79] Der shin Chang. Design of optimal control for a regression problem. Annals of Statistics, 7(5):1078–1085, September 1979. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344791. See

correction [CW80a].

Schoenfelder:1982:RDE

[SC82] Carol Schoenfelder and Stamatis Cambanis. Random designs for estimating integrals of stochastic processes. Annals of Statistics, 10(2):526–538, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345793.

Stefanski:1985:CME

[SC85] Leonard A. Stefanski and Raymond J. Carroll. Covariate measurement error in logistic regression. Annals of Statistics, 13(4): 1335–1351, December 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349741.

Schoenfeld:1980:TBL

[Sch80] David Schoenfeld. Tests based on linear combinations of the orthogonal components of the Cramér–von Mises statistic when parameters are estimated. Annals of Statistics, 8(5):1017–1022, September 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345139.

Schervish:1981:AEC

[Sch81a] Mark J. Schervish. Asymptotic expansions for correct classification rates in discriminant analysis. Annals of Statistics, 9(5): 1002–1009, September 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345579.

Schwarz:1981:AAS

[Sch81b] Gideon Schwarz. Applying asymptotic shapes to nonexponential families. Annals of Statistics, 9(2):461–464, March 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345415.

Schilling:1983:GFT

[Sch83a] Mark F. Schilling. Goodness of fit testing in R^{m} based on the weighted empirical distribution of certain nearest neighbor statistics. Annals of Statistics, 11(1):1–12, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346051.

Schilling:1983:IDA

[Sch83b] Mark F. Schilling. An infinite-dimensional approximation for nearest neighbor goodness of fit tests. Annals of Statistics, 11(1):13–24, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346052.

Schafer:1985:NDA

[Sch85a] Helmut Schafer. A note on data-adaptive kernel estimation of the hazard and density function in the random censorship situation. Annals of Statistics, 13(2):818–820, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349560.

Schervish:1985:DCB

[Sch85b] Mark J. Schervish. Discussion: Calibration-based empirical probability. Annals of Statistics, 13(4):1274–1282, December 1985. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349737.

Schafer:1986:LCE

[Sch86a] Helmut Schafer. Local convergence of empirical measures in the random censorship situation with application to density and rate estimators. Annals of Statistics, 14(3):1240–1245, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350063.

Schick:1986:AEE

[Sch86b] Anton Schick. On asymptotically efficient estimation in semiparametric models. Annals of Statistics, 14(3):1139–1151, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350055.

Schriever:1987:OPD

[Sch87a] B. F. Schriever. An ordering for positive dependence. Annals of Statistics, 15(3):1208–1214, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350500.

Schuster:1987:ICS

[Sch87b] Eugene F. Schuster. Identifying the closest symmetric distribution or density function. Annals of Statistics, 15(2):865–874, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350380.

Schervish:1989:GMC

[Sch89a] Mark J. Schervish. A general method for comparing probability assessors. Annals of Statistics, 17(4):1856–1879, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347398.

Schneller:1989:EEL

[Sch89b] Walter Schneller. Edgeworth expansions for linear rank statistics. Annals of Statistics, 17(3):1103–1123, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347258.

Scott:1985:ASH

[Sco85] David W. Scott. Averaged shifted histograms: Effective nonparametric density estimators in several dimensions. Annals of Statistics, 13(3):1024–1040, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349654.

Simpson:1987:EDD

[SCR87] Douglas G. Simpson, Raymond J. Carroll, and David Ruppert. M-estimation for discrete data: Asymptotic distribution theory and implications. Annals of Statistics, 15(2):657–669, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350367.

Sieders:1987:LDR

[SD87] Arthur Sieders and Kacha Dzhaparidze. A large deviation result for parameter estimators and its application to nonlinear regression analysis. Annals of Statistics, 15(3):1031–1049, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350491.

Seely:1983:AWV

[SEB83] Justus F. Seely and Yahia El-Bassiouni. Applying Wald’s variance component test. Annals of Statistics, 11(1):197–201, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346069.

Sen:1980:ASL

[Sen80] Pranab Kumar Sen. On almost sure linearity theorems for signed rank order statistics. Annals of Statistics, 8(2):313–321, March 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344956.

Sen:1981:CRM

[Sen81] Pranab Kumar Sen. The Cox regression model, invariance principles for some induced quantile processes and some repeated significance tests. Annals of Statistics, 9(1):109–121, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345336.

Sen:1982:IPR

[Sen82] Pranab Kumar Sen. Invariance principles for recursive residuals. Annals of Statistics, 10(1):307–312, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345714.

Sen:1988:FJR

[Sen88] Pranab Kumar Sen. Functional jackknifing: Rationality and general asymptotics. Annals of Statistics, 16(1):450–469, March 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350714.

Serfling:1984:GS

[Ser84] Robert J. Serfling. Generalized L-, M-, and R-statistics. Annals of Statistics, 12(1):76–86, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346393.

Siegmund:1980:SCT

[SG80] D. Siegmund and P. Gregory. A sequential clinical trial for testing p_{1} = p_{2}. Annals of Statistics, 8(6):1219–1228, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345195.

Stenger:1981:CCF

[SG81] H. Stenger and S. Gabler. On the completeness of the class of fixed size sampling strategies. Annals of Statistics, 9(1):229–232, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345354.

Shaffer:1980:CDE

[Sha80] Juliet Popper Shaffer. Control of directional errors with stagewise multiple test procedures. Annals of Statistics, 8(6):1342–1347, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345205.

Shafer:1982:BTA

[Sha82] Glenn Shafer. Bayes’s two arguments for the rule of conditioning. Annals of Statistics, 10(4):1075–1089, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345974.

Shao:1986:DJB

[Sha86] Jun Shao. Discussion: Jackknife, bootstrap and other resampling methods in regression analysis. Annals of Statistics, 14(4): 1322–1326, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350154.

Shao:1988:RMV

[Sha88a] Jun Shao. On resampling methods for variance and bias estimation in linear models. Annals of Statistics, 16(3):986–1008, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350945.

Shaw:1988:QAI

[Sha88b] J. E. H. Shaw. A quasirandom approach to integration in Bayesian statistics. Annals of Statistics, 16(2):895–914, June 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350842.

Shapiro:1989:APS

[Sha89] Alexander Shapiro. Asymptotic properties of statistical estimators in stochastic programming. Annals of Statistics, 17(2):841–858, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347146.

Shepp:1985:DPP

[She85] L. A. Shepp. Discussion: Projection pursuit. Annals of Statistics, 13(2):513–515, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349531.

Shibata:1980:AES

[Shi80] Ritei Shibata. Asymptotically efficient selection of the order of the model for estimating parameters of a linear process. Annals of Statistics, 8(1):147–164, January 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344897.

Shibata:1981:OAS

[Shi81] Ritei Shibata. An optimal autoregressive spectral estimate. Annals of Statistics, 9(2):300–306, March 1981. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345396.

Shinozaki:1984:SEL

[Shi84] Nobuo Shinozaki. Simultaneous estimation of location parameters under quadratic loss. Annals of Statistics, 12(1):322–335, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346410.

Siegmund:1986:BCP

[Sie86] David Siegmund. Boundary crossing probabilities and statistical applications. Annals of Statistics, 14(2):361–404, June 1986. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349928.

Silverman:1980:AWS

[Sil80] Bernard W. Silverman. Addendum to weak and strong uniform consistency of the kernel estimate of a density and its derivatives. Annals of Statistics, 8(5):1175–1176, September 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345157.

Silverman:1982:EPD

[Sil82] B. W. Silverman. On the estimation of a probability density function by the maximum penalized likelihood method. Annals of Statistics, 10(3):795–810, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345872.

Silverman:1984:SSE

[Sil84] B. W. Silverman. Spline smoothing: The equivalent variable kernel method. Annals of Statistics, 12(3):898–916, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346710.

Silvapulle:1985:ABR

[Sil85a] Mervyn J. Silvapulle. Asymptotic behavior of robust estimators of regression and scale parameters with fixed carriers. Annals of Statistics, 13(4):1490–1497, December 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349750.

Silverman:1985:TBD

[Sil85b] B. W. Silverman. Two books on density estimation. Annals of Statistics, 13(4):1630–1638, December 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349761.

Simons:1980:ESO

[Sim80] Gordon Simons. Extensions of the stochastic ordering property of likelihood ratios. Annals of Statistics, 8(4):833–839, July 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345075.

Simonoff:1983:PFA

[Sim83] Jeffrey S. Simonoff. A penalty function approach to smoothing large sparse contingency tables. Annals of Statistics, 11 (1):208–218, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346071.

Simons:1986:BRC

[Sim86] Gordon Simons. Bayes rules for a clinical-trials model with dichotomous responses. Annals of Statistics, 14(3):954–970, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350044.

Singh:1981:AAE

[Sin81a] Kesar Singh. On the asymptotic accuracy of Efron’s bootstrap. Annals of Statistics, 9(6):1187–1195, November 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345636.

Singh:1981:EAB

[Sin81b] R. S. Singh. On the exact asymptotic behavior of estimators of a density and its derivatives. Annals of Statistics, 9(2):453–456, March 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345413.

Sinha:1984:DMO

[Sin84] Bimal Kumar Sinha. Detection of multivariate outliers in elliptically symmetric distributions. Annals of Statistics, 12(4):

1558–1565, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346813.

Singh:1986:DJB

[Sin86] Kesar Singh. Discussion: Jackknife, bootstrap and other resampling methods in regression analysis. Annals of Statistics, 14(4): 1328–1330, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350156.

Speed:1986:GMD

[SK86] T. P. Speed and H. T. Kiiveri. Gaussian Markov distributions over finite graphs. Annals of Statistics, 14(1):138–150, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349846.

Skovgaard:1985:SOI

[Sko85] Ib M. Skovgaard. A second-order investigation of asymptotic ancillarity. Annals of Statistics, 13(2):534–551, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349537.

Sen:1989:SPS

[SKS89] Pranab Kumar Sen, Tatsuya Kubokawa, and A. K. Md. Ehsanes Saleh. The Stein paradox in the sense of the Pitman measure of closeness. Annals of Statistics, 17(3):1375–1386, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347276.

Slud:1982:CPS

[Slu82] Eric V. Slud. A characterization problem in stationary time series. Annals of Statistics, 10(2):630–633, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345805.

Slud:1984:SLR

[Slu84] Eric V. Slud. Sequential linear rank tests for two-sample censored survival data. Annals of Statistics, 12(2):551–571, June 1984. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346505.

Szatrowski:1980:EML

[SM80] Ted H. Szatrowski and John J. Miller. Explicit maximum likelihood estimates from balanced data in the mixed model of the analysis of variance. Annals of Statistics, 8(4):811–819, July 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345073.

Schwager:1982:DMN

[SM82] Steven J. Schwager and Barry H. Margolin. Detection of multivariate normal outliers. Annals of Statistics, 10(3):943–954, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345884.

Small:1988:GAC

[SM88] Christopher G. Small and D. L. McLeish. Generalizations of ancillarity, completeness and sufficiency in an inference function space. Annals of Statistics, 16(2):534–551, June 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350819.

Small:1983:CTM

[Sma83] Christopher G. Small. Characterization of type from maximal invariant spectra. Annals of Statistics, 11(3):979–983, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346263.

Smith:1984:PBC

[Smi84] Richard L. Smith. Properties of biased coin designs in sequential clinical trials. Annals of Statistics, 12(3):1018–1034, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346718.

Smith:1987:ETP

[Smi87] Richard L. Smith. Estimating tails of probability distributions. Annals of Statistics, 15(3):1174–1207, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350499.

Smith:1989:IDS

[Smi89] J. Q. Smith. Influence diagrams for statistical modelling. Annals of Statistics, 17(2):654–672, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347132.

Smythe:1988:CIR

[Smy88] R. T. Smythe. Conditional inference for restricted randomization designs. Annals of Statistics, 16(3):1155–1161, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350952.

Snijders:1981:RTB

[Sni81] Tom Snijders. Rank tests for bivariate symmetry. Annals of Statistics, 9(5):1087–1095, September 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345588.

Solo:1981:SOP

[Sol81a] V. Solo. The second order properties of a time series recursion. Annals of Statistics, 9(2):307–317, March 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345397.

Solo:1981:SCL

[Sol81b] V. Solo. Strong consistency of least squares estimators in regression with correlated disturbances. Annals of Statistics, 9 (3):689–693, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345476.

Song:1982:CST

[Son82] C. C. Song. Covariance stabilizing transformations and a conjecture of Holland. Annals of Statistics, 10(1):313–315, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345715.

Self:1982:CAG

[SP82] Steven G. Self and Ross L. Prentice. Commentary on Andersen and Gill’s “Cox’s Regression Model for Counting Processes: A Large Sample Study”. Annals of Statistics, 10(4):1121–1124, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345977.

Self:1988:ADT

[SP88] Steven G. Self and Ross L. Prentice. Asymptotic distribution theory and efficiency results for case-cohort studies. Annals of Statistics, 16(1):64–81, March 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350691.

Speckman:1985:SSO

[Spe85] Paul Speckman. Spline smoothing and optimal rates of convergence in nonparametric regression models. Annals of Statistics, 13 (3):970–983, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349650.

Speed:1987:RWA

[Spe87a] T. P. Speed. Rejoinder: What is an analysis of variance? Annals of Statistics, 15(3):937–941, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350484.

Speed:1987:WAV

[Spe87b] T. P. Speed. What is an analysis of variance? Annals of Statistics, 15(3):885–910, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350472.

Spruill:1980:ODS

[Spr80] Carl Spruill. Optimal designs for second order processes with general linear means. Annals of Statistics, 8(3):652–663, May 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345015.

Sproule:1985:SNF

[Spr85] Raymond N. Sproule. Sequential nonparametric fixed-width confidence intervals for U-statistics. Annals of Statistics, 13(1):228–235, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346588.

Srinivasan:1984:SNS

[Sri84] C. Srinivasan. A sharp necessary and sufficient condition for inadmissibility of estimators in a control problem. Annals of Statistics, 12(3):927–944, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346712.

Srivastava:1986:DJB

[Sri86] M. S. Srivastava. Discussion: Jackknife, bootstrap and other resampling methods in regression analysis. Annals of Statistics, 14 (4):1331–1335, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350157.

Sriram:1987:SEM

[Sri87] T. N. Sriram. Sequential estimation of the mean of a first-order stationary autoregressive process. Annals of Statistics, 15(3): 1079–1090, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350494.

Spiegelman:1980:CWE

[SS80] C. Spiegelman and J. Sacks. Consistent window estimation in nonparametric regression. Annals of Statistics, 8(2):240–246, March 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344950.

Sethuraman:1981:LSE

[SS81] J. Sethuraman and Nozer D. Singpurwalla. Large sample estimates and uniform confidence bounds for the failure rate function based on a naive estimator. Annals of Statistics, 9(3):628–632, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345466.

Sinha:1984:ICS

[SS84] B. K. Sinha and S. K. Sarkar. Invariant confidence sequences for some parameters in a multivariate linear regression model. Annals of Statistics, 12(1):301–310, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346408.

Sen:1985:SSE

[SS85] Pranab Kumar Sen and A. K. MD. Ehsanes Saleh. On some shrinkage estimators of multivariate location. Annals of Statistics, 13 (1):272–281, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346592.

Sen:1987:PTS

[SS87] Pranab Kumar Sen and A. K. M. Ehsanes Saleh. On preliminary test and shrinkage M-estimation in linear models. Annals of Statistics, 15(4):1580–1592, December 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350611.

Sathe:1989:OWD

[SS89a] Y. S. Sathe and R. G. Shenoy. A-optimal weighing designs when N ≡ 3(mod4). Annals of Statistics, 17(4):1906–1915, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347401.

Stine:1989:FPC

[SS89b] Robert A. Stine and Paul Shaman. A fixed point characterization for bias of autoregressive estimators. Annals of Statistics, 17(3): 1275–1284, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347268.

Sverdrup-Thygeson:1981:SLL

[ST81] Harald Sverdrup-Thygeson. Strong law of large numbers for measures of central tendency and dispersion of random variables in compact metric spaces. Annals of Statistics, 9(1):141–145, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345340.

Sager:1982:MLE

[ST82] Thomas W. Sager and Ronald A. Thisted. Maximum likelihood estimation of isotonic modal regression. Annals of Statistics, 10(3): 690–707, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345865.

Simonoff:1986:DJB

[ST86] Jeffrey S. Simonoff and Chih-Ling Tsai. Discussion: Jackknife, bootstrap and other resampling methods in regression analysis. Annals of Statistics, 14(4):1326–1328, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350155.

Stadje:1985:EPS

[Sta85] Wolfgang Stadje. Estimation problems for samples with measurement errors. Annals of Statistics, 13(4):1592–1615, December 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349757.

Stein:1981:EMM

[Ste81] Charles M. Stein. Estimation of the mean of a multivariate normal distribution. Annals of Statistics, 9(6):1135–1151, November 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345632.

Steele:1986:ESI

[Ste86a] J. Michael Steele. An Efron–Stein inequality for nonsymmetric statistics. Annals of Statistics, 14(2):753–758, June 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349952.

Stefanov:1986:ESE

[Ste86b] Valeri T. Stefanov. Efficient sequential estimation in exponentialtype processes. Annals of Statistics, 14(4):1606–1611, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350181.

Stein:1988:AEP

[Ste88] Michael L. Stein. Asymptotically efficient prediction of a random field with a misspecified covariance function. Annals of Statistics, 16(1):55–63, March 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350690.

Stein:1989:ADM

[Ste89a] Michael Stein. Asymptotic distributions of minimum norm quadratic estimators of the covariance function of a Gaussian random field. Annals of Statistics, 17(3):980–1000, September 1989.

CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347252.

Stenger:1989:AAM

[Ste89b] Horst Stenger. Asymptotic analysis of minimax strategies in survey sampling. Annals of Statistics, 17(3):1301–1314, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347270.

Stigler:1980:EC

[Sti80] Stephen M. Stigler. An Edgeworth curiosum. Annals of Statistics, 8(4):931–934, July 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345085.

Stigler:1981:GIL

[Sti81] Stephen M. Stigler. Gauss and the invention of least squares. Annals of Statistics, 9(3):465–474, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345451.

Stijnen:1985:ABE

[Sti85] Theo Stijnen. On the asymptotic behaviour of empirical Bayes tests for the continuous one-parameter exponential family. Annals of Statistics, 13(1):403–412, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346601.

Stone:1980:ORC

[Sto80] Charles J. Stone. Optimal rates of convergence for nonparametric estimators. Annals of Statistics, 8(6):1348–1360, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345206.

Stone:1981:ASA

[Sto81] Charles J. Stone. Admissible selection of an accurate and parsimonious normal linear regression model. Annals of Statistics, 9(3):475–485, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345452.

Stone:1982:OGR

[Sto82] Charles J. Stone. Optimal global rates of convergence for nonparametric regression. Annals of Statistics, 10(4):1040–1053, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345969.

Stone:1984:AOW

[Sto84] Charles J. Stone. An asymptotically optimal window selection rule for kernel density estimates. Annals of Statistics, 12(4): 1285–1297, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346792.

Stone:1985:ARO

[Sto85] Charles J. Stone. Additive regression and other nonparametric models. Annals of Statistics, 13(2):689–705, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349548.

Stone:1986:DRP

[Sto86] Charles J. Stone. The dimensionality reduction principle for generalized additive models. Annals of Statistics, 14(2):590–606, June 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349940.

Strasser:1981:CML

[Str81] Helmut Strasser. Consistency of maximum likelihood and Bayes estimates. Annals of Statistics, 9(5):1107–1113, September 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345590.

Scott:1980:NPD

[STT80] D. W. Scott, R. A. Tapia, and J. R. Thompson. Nonparametric probability density estimation by discrete maximum penalizedlikelihood criteria. Annals of Statistics, 8(4):820–832, July 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345074.

Studden:1980:ODP

[Stu80] W. J. Studden. D_{s}-optimal designs for polynomial regression using continued fractions. Annals of Statistics, 8(5):1132–1141, September 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345150.

Stute:1984:ANN

[Stu84] Winfried Stute. Asymptotic normality of nearest neighbor regression function estimates. Annals of Statistics, 12(3):917–926, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346711.

Stute:1986:CEP

[Stu86] Winfried Stute. Conditional empirical processes. Annals of Statistics, 14(2):638–647, June 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349943.

Stufken:1987:OBD

[Stu87] John Stufken. A-optimal block designs for comparing test treatments with a control. Annals of Statistics, 15(4):1629–1638, December 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350614.

Studden:1989:NSO

[Stu89] W. J. Studden. Note on some ϕ_{p}-optimal designs for polynomial regression. Annals of Statistics, 17(2):618–623, June 1989. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347129.

Sundberg:1985:DTI

[Sun85] Rolf Sundberg. Discussion: Testing for independence in a twoway table: New interpretations of the chi-square statistic. Annals of Statistics, 13(3):903–904, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349644.

Susarla:1980:LST

[SV80a] V. Susarla and J. Van Ryzin. Large sample theory for an estimator of the mean survival time from censored samples. Annals of Statistics, 8(5):1002–1016, September 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345138.

Susarla:1980:ALS

[SV80b] V. Susarla and John Van Ryzin. Addendum to “Large Sample Theory for a Bayesian Nonparametric Survival Curve Estimator Based on Censored Data”. Annals of Statistics, 8(3):693, May 1980. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345021.

Svensson:1981:GFT

[Sve81] Ake Svensson. On a goodness-of-fit test for multiplicative Poisson models. Annals of Statistics, 9(4):697–704, July 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345512.

Svensson:1989:ESC

[Sve89] Ake Svensson. Estimation in some counting process models with multiplicative structure. Annals of Statistics, 17(4):1501–1508, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347378.

Shapiro:1980:DIA

[SW80] C. P. Shapiro and Robert L. Wardrop. Dynkin’s identity applied to Bayes sequential estimation of a Poisson process rate. Annals of Statistics, 8(1):171–182, January 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344899.

Schweizer:1981:NMD

[SW81a] B. Schweizer and E. F. Wolff. On nonparametric measures of dependence for random variables. Annals of Statistics, 9(4):879–885, July 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345528.

Susarla:1981:EMD

[SW81b] V. Susarla and G. Walter. Estimation of a multivariate density function using delta sequences. Annals of Statistics, 9(2):347–355, March 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345400.

Shirahata:1984:ANC

[SW84] Shingo Shirahata and Kazumasa Wakimoto. Asymptotic normality of a class of nonlinear rank tests for independence. Annals of Statistics, 12(3):1124–1129, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346730.

Shao:1987:HRJ

[SW87] Jun Shao and C. F. J. Wu. Heteroscedasticity-robustness of jackknife variance estimators in linear models. Annals of Statistics, 15 (4):1563–1579, December 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350610.

Schenker:1988:ARM

[SW88] Nathaniel Schenker and A. H. Welsh. Asymptotic results for multiple imputation. Annals of Statistics, 16(4):1550–1566, December 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176351053.

Shao:1989:GTJ

[SW89] Jun Shao and C. F. J. Wu. A general theory for jackknife variance estimation. Annals of Statistics, 17(3):1176–1197, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347263.

Sweeting:1980:UAN

[Swe80a] T. J. Sweeting. Uniform asymptotic normality of the maximum likelihood estimator. Annals of Statistics, 8(6):1375–1381, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.

aos/1176345208. See correction [Swe82].

Swensen:1980:DBL

[Swe80b] Anders Rygh Swensen. Deficiencies between linear normal experiments. Annals of Statistics, 8(5):1142–1155, September 1980. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345151.

Sweeting:1982:CUA

[Swe82] T. J. Sweeting. Corrections: “Uniform Asymptotic Normality of the Maximum Likelihood Estimator”. Annals of Statistics, 10(1):320, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345717. See [Swe80a].

Sweeting:1986:ACI

[Swe86a] T. J. Sweeting. Asymptotic conditional inference for the offspring mean of a supercritical Galton–Watson process. Annals of Statistics, 14(3):925–933, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350042.

Sweeting:1986:AIS

[Swe86b] T. J. Sweeting. Asymptotically independent scale-free spacings with applications to discordancy testing. Annals of Statistics, 14 (4):1485–1496, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350171.

Sweeting:1986:CST

[Swe86c] T. J. Sweeting. On a converse to Scheffe’s theorem. Annals of Statistics, 14(3):1252–1256, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350065.

Switzer:1985:DPP

[Swi85] Paul Switzer. Discussion: Projection pursuit. Annals of Statistics, 13(2):515–517, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349532.

Stepniak:1984:CLE

[SWW84] Czeslaw Stepniak, Song-Gui Wang, and C. F. Jeff Wu. Comparison of linear experiments with known covariances. Annals of Statistics, 12(1):358–365, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346413.

Sacks:1981:AOK

[SY81] Jerome Sacks and Donald Ylvisaker. Asymptotically optimum kernels for density estimation at a point. Annals of Statistics, 9(2):334–346, March 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345399.

Sacks:1982:EMP

[SY82] Jerome Sacks and Donald Ylvisaker. L- and R-estimation and the minimax property. Annals of Statistics, 10(2):643–645, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345808.

Sacks:1984:SMR

[SY84] Jerome Sacks and Donald Ylvisaker. Some model robust designs in regression. Annals of Statistics, 12(4):1324–1348, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346795.

Stengle:1989:SNV

[SY89] G. Stengle and J. E. Yukich. Some new Vapnik–Chervonenkis classes. Annals of Statistics, 17(4):1441–1446, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347373.

Simons:1989:STD

[SYW89] Gordon Simons, Yi-Ching Yao, and Xizhi Wu. Sequential tests for the drift of a Wiener process with a smooth prior, and the heat equation. Annals of Statistics, 17(2):783–792, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347142.

Szatrowski:1980:NSC

[Sza80] Ted H. Szatrowski. Necessary and sufficient conditions for explicit solutions in the multivariate normal estimation problem for patterned means and covariances. Annals of Statistics, 8(4):802– 810, July 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345072.

Szatrowski:1983:MDO

[Sza83] Ted H. Szatrowski. Missing data in the one-population multivariate normal patterned mean and covariance matrix testing and estimation problem. Annals of Statistics, 11(3):947–958, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346260.

Szkutnik:1988:MPI

[Szk88] Zbigniew Szkutnik. Most powerful invariant tests for binormality. Annals of Statistics, 16(1):292–301, March 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350706.

Tahir:1989:ALB

[Tah89] Mohamed Tahir. An asymptotic lower bound for the local minimax regret in sequential point estimation. Annals of Statistics, 17(3): 1335–1346, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347273.

Takada:1981:RBI

[Tak81] Yoshikazu Takada. Relation of the best invariant predictor and the best unbiased predictor in location and scale families. Annals of Statistics, 9(4):917–921, July 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345534.

Takada:1982:CBI

[Tak82] Yoshikazu Takada. A comment on best invariant predictors. Annals of Statistics, 10(3):971–978, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345887.

Takahashi:1987:AEA

[Tak87] Hajime Takahashi. Asymptotic expansions in Anscombe’s theorem for repeated significance tests and estimation after sequential testing. Annals of Statistics, 15(1):278–295, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350266.

Taniguchi:1983:SOA

[Tan83a] Masanobu Taniguchi. On the second order asymptotic efficiency of estimators of Gaussian ARMA processes. Annals of Statistics, 11(1):157–169, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346066.

Tanner:1983:NVK

[Tan83b] Martin A. Tanner. A note on the variable kernel estimator of the hazard function from randomly censored data. Annals of Statistics, 11(3):994–998, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346266.

Tardif:1981:ASC

[Tar81] Serge Tardif. On the almost sure convergence of the permutation distribution for aligned rank test statistics in randomized block designs. Annals of Statistics, 9(1):190–193, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345346.

Tarter:1965:PMO

[TC65] Michael E. Tarter and Virginia A. Clark. Properties of the median and other order statistics of logistic variates. Annals of Mathematical Statistics, 36(6):1779–1786, December 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177699806. See correction

[TC80].

Tarter:1980:COS

[TC80] Michael E. Tarter and Virginia A. Clark. Correction to “Order Statistics of Logistic Variates”. Annals of Statistics, 8(4):935, July 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345086. See [TC65].

Tsai:1985:LSS

[TC85] Wei-Yann Tsai and John Crowley. A large sample study of generalized maximum likelihood estimators from incomplete data via selfconsistency. Annals of Statistics, 13(4):1317–1334, December 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349740. See

correction [TC90].

Tsai:1990:CLS

[TC90] Wei-Yann Tsai and John Crowley. Correction: “A Large Sample Study of Generalized Maximum Likelihood Estimators from Incomplete Data via Self-Consistency”. Annals of Statistics, 18(1): 470, March 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347514. See [TC85].

Tibshirani:1986:DJB

[Tib86] Robert Tibshirani. Discussion: Jackknife, bootstrap and other resampling methods in regression analysis. Annals of Statistics, 14 (4):1335–1339, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350158.

Tibshirani:1988:CDJ

[Tib88] Robert Tibshirani. Correction: Discussion of “Jackknife, Bootstrap and Other Resampling Methods in Regression Analysis” by c. f. j. Wu. Annals of Statistics, 16(1):479, March 1988. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350716. See

[Wu86a].

Tierney:1987:ARC

[Tie87] Luke Tierney. An alternative regularity condition for Hajek’s representation theorem. Annals of Statistics, 15(1):427–431, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350277.

Tierney:1988:APA

[Tie88] Luke Tierney. Acknowledgement of priority: An alternative regularity condition for Hajek’s representation theorem. Annals of Statistics, 16(2):926, June 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350844.

Titterington:1989:DLS

[Tit89] D. M. Titterington. Discussion: Linear smoothers and additive models. Annals of Statistics, 17(2):540–543, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347123.

Tjur:1987:DWA

[Tju87] Tue Tjur. Discussion: What is an analysis of variance? Annals of Statistics, 15(3):931–932, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350481.

Toyooka:1986:AUB

[TK86] Yasuyuki Toyooka and Takeaki Kariya. An approach to upper bound problems for risks of generalized least squares estimators. Annals of Statistics, 14(2):679–690, June 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349946.

Tierney:1984:AEE

[TL84] Luke Tierney and Diane Lambert. Asymptotic efficiency of estimators of functionals of mixed distributions. Annals of Statistics, 12 (4):1380–1387, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346798.

Tsai:1986:NEB

[TLC86] Wei-Yann Tsai, Sue Leurgans, and John Crowley. Nonparametric estimation of a bivariate survival function in the presence of censoring. Annals of Statistics, 14(4):1351–1365, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350162.

Tobias:1987:DWA

[Tob87] Randall D. Tobias. Discussion: What is an analysis of variance? Annals of Statistics, 15(3):932–936, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350482.

Tong:1982:REP

[Ton82] Y. L. Tong. Rectangular and elliptical probability inequalities for Schur-concave random variables. Annals of Statistics, 10 (2):637–642, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345807.

Tong:1989:ICP

[Ton89] Y. L. Tong. Inequalities for a class of positively dependent random variables with a common marginal. Annals of Statistics, 17 (1):429–435, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347026.

Torgersen:1981:MIB

[Tor81] Erik N. Torgersen. Measures of information based on comparison with total information and with total ignorance. Annals of Statistics, 9(3):638–657, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345468.

Toyooka:1986:SOR

[Toy86] Yasuyuki Toyooka. Second-order risk structure of GLSE and MLE in a regression with a linear process. Annals of Statistics, 14(3): 1214–1225, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350060.

Tsui:1982:SES

[TP82] Kam-Wah Tsui and S. James Press. Simultaneous estimation of several Poisson parameters under K-normalized squared error loss. Annals of Statistics, 10(1):93–100, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345692.

Truong:1989:APK

[Tru89] Young K. Truong. Asymptotic properties of kernel estimators based on local medians. Annals of Statistics, 17(2):606–617, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347128.

Terrell:1980:ICR

[TS80] George R. Terrell and David W. Scott. On improving convergence rates for nonnegative kernel density estimators. Annals of Statistics, 8(5):1160–1163, September 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345153.

Tsay:1984:OSN

[Tsa84] Ruey S. Tsay. Order selection in nonstationary autoregressive models. Annals of Statistics, 12(4):1425–1433, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346801.

Tsai:1986:ESC

[Tsa86a] Wei-Yann Tsai. Estimation of survival curves from dependent censorship models via a generalized self-consistent property with nonparametric Bayesian estimation application. Annals of Statistics, 14(1):238–249, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349852.

Tsay:1986:DIF

[Tsa86b] Ruey S. Tsay. Discussion: Influence functionals for time series. Annals of Statistics, 14(3):835–836, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350034.

Tsiatis:1981:LSS

[Tsi81] Anastasios A. Tsiatis. A large sample study of Cox’s regression model. Annals of Statistics, 9(1):93–108, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345335.

Tiao:1983:CPL

[TT83] George C. Tiao and Ruey S. Tsay. Consistency properties of least squares estimates of autoregressive parameters in ARMA models. Annals of Statistics, 11(3):856–871, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346252.

Tuan:1981:ESP

[Tua81] Pham Dinh Tuan. Estimation of the spectral parameters of a stationary point process. Annals of Statistics, 9(3):615–627, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345465.

Tukey:1985:PPD

[Tuk85a] J. Tukey. Projection pursuit — discussion. Annals of Statistics, 13(2):517–518, ???? 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Tukey:1985:DPP

[Tuk85b] John Tukey. Discussion: Projection pursuit. Annals of Statistics, 13(2):517–518, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349533.

Tukey:1987:WAV

[Tuk87a] J. W. Tukey. What is an analysis of variance — discussion. Annals of Statistics, 15(3):936–937, ???? 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

Tukey:1987:DWA

[Tuk87b] John W. Tukey. Discussion: What is an analysis of variance? Annals of Statistics, 15(3):936–937, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350483.

Tanner:1983:EHF

[TW83] Martin A. Tanner and Wing Hung Wong. The estimation of the hazard function from randomly censored data by the kernel method. Annals of Statistics, 11(3):989–993, September 1983. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346265.

Tsui:1980:IBF

[TWZ80] Kam-Wah Tsui, Samaradasa Weerahandi, and Jim Zidek. Inadmissibility of the best fully equivariant estimator of the generalized residual variance. Annals of Statistics, 8(5):1156–1159, September 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345152.

Tyler:1981:AIE

[Tyl81] David E. Tyler. Asymptotic inference for eigenvectors. Annals of Statistics, 9(4):725–736, July 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345514.

Tyler:1983:ADP

[Tyl83a] David E. Tyler. The asymptotic distribution of principal component roots under local alternatives to multiple roots. Annals of Statistics, 11(4):1232–1242, December 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346336.

Tyler:1983:CAT

[Tyl83b] David E. Tyler. A class of asymptotic tests for principal component vectors. Annals of Statistics, 11(4):1243–1250, December 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346337.

Tyler:1987:DFE

[Tyl87] David E. Tyler. A distribution-free M-estimator of multivariate scatter. Annals of Statistics, 15(1):234–251, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350263.

VanDerPlas:1983:EPM

[Van83] Adriaan P. Van Der Plas. On the estimation of the parameters of Markov probability models using macro data. Annals of Statistics, 11(1):78–85, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346058.

VanDeGeer:1987:NAL

[Van87] Sara Van De Geer. A new approach to least-squares estimation, with applications. Annals of Statistics, 15(2):587–602, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350362.

Vardi:1982:NER

[Var82a] Y. Vardi. Nonparametric estimation in renewal processes. Annals of Statistics, 10(3):772–785, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345870.

Vardi:1982:NEP

[Var82b] Y. Vardi. Nonparametric estimation in the presence of length bias. Annals of Statistics, 10(2):616–620, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345802.

Vardi:1985:EDS

[Var85] Y. Vardi. Empirical distributions in selection bias models. Annals of Statistics, 13(1):178–203, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346585.

Voelkel:1984:NIC

[VC84] Joseph G. Voelkel and John Crowley. Nonparametric inference for a class of semi-Markov processes with censored observations. Annals of Statistics, 12(1):142–160, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346398.

vanDawen:1985:NCO

[vD85] R. van Dawen. A note on the characterization of optimal return functions and optimal strategies for gambling problems. Annals of Statistics, 13(2):832–835, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349563.

Voss:1987:CCS

[VD87] D. T. Voss and A. M. Dean. A comparison of classes of single replicate factorial designs. Annals of Statistics, 15(1):376–384, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350272.

vanderVaart:1988:ERP

[vdV88] A. W. van der Vaart. Estimating a real parameter in a class of semiparametric models. Annals of Statistics, 16(4):1450–1474, December 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176351048.

vanderVaart:1989:AIB

[vdV89] Aad van der Vaart. On the asymptotic information bound. Annals of Statistics, 17(4):1487–1500, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347377.

vanEeden:1983:ARB

[vE83] Constance van Eeden. On the asymptotic relation between Lestimators and M-estimators and their asymptotic efficiency relative to the Cramér–Rao lower bound. Annals of Statistics, 11 (2):674–690, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346172.

Veall:1988:DTC

[Vea88] Michael R. Veall. Discussion: Theoretical comparison of bootstrap confidence intervals. Annals of Statistics, 16(3):979–981, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350943.

vanHouwelingen:1987:MEB

[vH87] J. C. van Houwelingen. Monotone empirical Bayes test for uniform distributions using the maximum likelihood estimator of a decreasing density. Annals of Statistics, 15(2):875–879, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350381.

Villegas:1981:ISI

[Vil81] C. Villegas. Inner statistical inference II. Annals of Statistics, 9(4):768–776, July 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345517.

Villegas:1982:MLL

[Vil82] C. Villegas. Maximum likelihood and least squares estimation in linear and affine functional models. Annals of Statistics, 10 (1):256–265, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345708.

Viollaz:1980:ADN

[Vio80] Aldo Jose Viollaz. Asymptotic distribution of L_{2} norms of the deviations of density function estimates. Annals of Statistics, 8 (2):322–346, March 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176344957.

Verrill:1987:AES

[VJ87] Steve Verrill and Richard A. Johnson. The asymptotic equivalence of some modified Shapiro-wilk statistics– complete and censored sample cases. Annals of Statistics, 15(1):413–419, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350275.

Vardeman:1984:AET

[VM84] Stephen Vardeman and Glen Meeden. Admissible estimators for the total of a stratified population that employ prior information. Annals of Statistics, 12(2):675–684, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346514.

Voshaar:1980:MSL

[Vos80] J. H. Oude Voshaar. (k − 1)-mean significance levels of nonparametric multiple comparisons procedures. Annals of Statistics, 8 (1):75–86, January 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176344892.

Voss:1988:SGG

[Vos88] D. T. Voss. Single-generator generalized cyclic factorial designs as pseudofactor designs. Annals of Statistics, 16(4):1723–1726, December 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176351064.

Wachter:1980:LEM

[Wac80] Kenneth W. Wachter. The limiting empirical measure of multiple discriminant ratios. Annals of Statistics, 8(5):937–957, September 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345134.

Wahba:1981:DBO

[Wah81] Grace Wahba. Data-based optimal smoothing of orthogonal series density estimates. Annals of Statistics, 9(1):146–156, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345341.

Wahba:1985:CGG

[Wah85a] Grace Wahba. A comparison of GCV and GML for choosing the smoothing parameter in the generalized spline smoothing problem. Annals of Statistics, 13(4):1378–1402, December 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349743.

Wahba:1985:DPP

[Wah85b] Grace Wahba. Discussion: Projection pursuit. Annals of Statistics, 13(2):518–521, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349534.

Wakker:1981:APM

[Wak81] Peter Wakker. Agreeing probability measures for comparative probability structures. Annals of Statistics, 9(3):658–662, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345469.

Walter:1980:NAP

[Wal80] Gilbert G. Walter. Note: Addendum to “Properties of Hermite Series Estimation of Probability Density”. Annals of Statistics, 8(2):454–455, March 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176344965.

Walley:1987:BFR

[Wal87] Peter Walley. Belief function representations of statistical evidence. Annals of Statistics, 15(4):1439–1465, December 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350603.

Wang:1981:RAT

[Wan81] Paul C. C. Wang. Robust asymptotic tests of statistical hypotheses involving nuisance parameters. Annals of Statistics, 9(5): 1096–1106, September 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345589.

Wang:1985:SCA

[Wan85] Jane-Ling Wang. Strong consistency of approximate maximum likelihood estimators with applications in nonparametrics. Annals of Statistics, 13(3):932–946, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349647.

Wang:1986:MDE

[Wan86a] Chamont W. H. Wang. A minimum distance estimator for firstorder autoregressive processes. Annals of Statistics, 14(3):1180– 1193, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350058.

Wang:1986:AME

[Wan86b] Jane-Ling Wang. Asymptotically minimax estimators for distributions with increasing failure rate. Annals of Statistics, 14(3): 1113–1131, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350053.

Wang:1987:NUC

[Wan87] Jia-Gang Wang. A note on the uniform consistency of the Kaplan– Meier estimator. Annals of Statistics, 15(3):1313–1316, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350507.

Wasserman:1989:RBI

[Was89] Larry Alan Wasserman. A robust Bayesian interpretation of likelihood regions. Annals of Statistics, 17(3):1387–1393, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347277.

Watson:1982:WGC

[Wat82] G. S. Watson. William Gemmell Cochran 1909–1980. Annals of Statistics, 10(1):1–10, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345687.

Walter:1984:SSN

[WB84] Gilbert G. Walter and Julius R. Blum. A simple solution to a nonparametric maximum likelihood estimation problem. Annals of Statistics, 12(1):372–379, March 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346415.

Weber:1986:DJB

[Web86] Neville Weber. Discussion: Jackknife, bootstrap and other resampling methods in regression analysis. Annals of Statistics, 14(4): 1339–1340, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350159.

Wegman:1986:DIF

[Weg86] Edward J. Wegman. Discussion: Influence functionals for time series. Annals of Statistics, 14(3):836–837, September 1986. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350035.

Wei:1985:APL

[Wei85] C. Z. Wei. Asymptotic properties of least-squares estimates in stochastic regression models. Annals of Statistics, 13(4):1498– 1508, December 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349751.

Wei:1987:APL

[Wei87a] C. Z. Wei. Adaptive prediction by least squares predictors in stochastic regression models with applications to time series. Annals of Statistics, 15(4):1667–1682, December 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350617.

Wei:1987:MAS

[Wei87b] C. Z. Wei. Multivariate adaptive stochastic approximation. Annals of Statistics, 15(3):1115–1130, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350496.

Wellner:1982:AOP

[Wel82] Jon A. Wellner. Asymptotic optimality of the product limit estimator. Annals of Statistics, 10(2):595–602, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345800.

Wellner:1985:HCL

[Wel85] Jon A. Wellner. A heavy censoring limit theorem for the product limit estimator. Annals of Statistics, 13(1):150–162, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346583.

Welsh:1986:BRR

[Wel86] A. H. Welsh. Bahadur representations for robust scale estimators based on regression residuals. Annals of Statistics, 14(3): 1246–1251, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350064.

Welsh:1987:OSE

[Wel87a] A. H. Welsh. One-step L-estimators for the linear model. Annals of Statistics, 15(2):626–641, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:/

/projecteuclid.org/euclid.aos/1176350365. See correction

[Wel88a].

Welsh:1987:RTM

[Wel87b] A. H. Welsh. Rejoinder: The trimmed mean in the linear model. Annals of Statistics, 15(1):44–45, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350251.

Welsh:1987:TML

[Wel87c] A. H. Welsh. The trimmed mean in the linear model. Annals of Statistics, 15(1):20–36, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:/

/projecteuclid.org/euclid.aos/1176350248. See correction

[Wel88b].

Welsh:1988:COS

[Wel88a] A. H. Welsh. Correction: “One-Step L-Estimators for the Linear Model”. Annals of Statistics, 16(1):481, March 1988. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350718. See

[Wel87a].

Welsh:1988:CTM

[Wel88b] A. H. Welsh. Correction: “The Trimmed Mean in the Linear Model”. Annals of Statistics, 16(1):480, March 1988. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350717. See

[Wel87c].

Welsh:1989:PE

[Wel89] A. H. Welsh. On M-processes and M-estimation. Annals of Statistics, 17(1):337–361, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:/

/projecteuclid.org/euclid.aos/1176347021. See correction

[Wel90].

Welsh:1990:CPE

[Wel90] A. H. Welsh. Correction: “On M-Processes and M-Estimation”. Annals of Statistics, 18(3):1500, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347766. See

[Wel89].

Weng:1989:SOA

[Wen89] Chung-Sing Weng. On a second-order asymptotic property of the Bayesian bootstrap mean. Annals of Statistics, 17(2):705–710, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347136.

West:1986:DIF

[Wes86a] Mike West. Discussion: Influence functionals for time series. Annals of Statistics, 14(3):838–840, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176350036.

Westfall:1986:ANA

[Wes86b] Peter H. Westfall. Asymptotic normality of the anova estimates of components of variance in the nonnormal, unbalanced hierarchal mixed model. Annals of Statistics, 14(4):1572–1582, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350177.

Westfall:1989:PCI

[Wes89] Peter H. Westfall. Power comparisons for invariant variance ratio tests in mixed anova models. Annals of Statistics, 17(1):318–326, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347019.

Walley:1982:TFT

[WF82a] Peter Walley and Terrence L. Fine. Towards a frequentist theory of upper and lower probability. Annals of Statistics, 10(3): 741–761, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345868.

Wolter:1982:ENE

[WF82b] Kirk M. Wolter and Wayne A. Fuller. Estimation of nonlinear errors-in-variables models. Annals of Statistics, 10(2):539–548, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345794.

Wiens:1986:MVE

[Wie86] Doug Wiens. Minimax variance M-estimators of location in Kolmogorov neighbourhoods. Annals of Statistics, 14(2):724–732, June 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349949.

Wijsman:1983:MNP

[Wij83] Robert A. Wijsman. Monotonicity in the noncentrality parameter of the ratio of two noncentral t-densities. Annals of Statistics, 11(3):1008–1010, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346269.

Wijsman:1984:RTB

[Wij84] Robert A. Wijsman. Review: Two books on multivariate analysis: Robb J. Muirhead, Aspects of Multivariate Statistical Theory; Morris L. Eaton, Multivariate Statistics. A Vector Space Approach. Annals of Statistics, 12(3):1145–1150, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346733.

Wijsman:1985:PAS

[Wij85] Robert A. Wijsman. Proper action in steps, with application to density ratios of maximal invariants. Annals of Statistics, 13 (1):395–402, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346600. See correction [Wij93].

Wijsman:1993:CPA

[Wij93] Robert A. Wijsman. Correction: “Proper Action in Steps, with Application to Density Ratios of Maximal Invariants”. Annals of Statistics, 21(4):2168–2169, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349418. See [Wij85].

Withers:1983:EDQ

[Wit83] C. S. Withers. Expansions for the distribution and quantiles of a regular functional of the empirical distribution with applications to nonparametric confidence intervals. Annals of Statistics, 11(2):577–587, June 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346163.

Witmer:1986:BMD

[Wit86] Jeffrey A. Witmer. Bayesian multistage decision problems. Annals of Statistics, 14(1):283–297, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349856.

Wang:1986:APP

[WJT86] Mei-Cheng Wang, Nicholas P. Jewell, and Wei-Yann Tsai. Asymptotic properties of the product limit estimate under random truncation. Annals of Statistics, 14(4):1597–1605, December 1986. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350180.

Wild:1981:NPF

[WK81] C. J. Wild and J. D. Kalbfleisch. A note on a paper by Ferguson and Phadia. Annals of Statistics, 9(5):1061–1065, September 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345585.

Wynn:1986:DJB

[WO86] H. P. Wynn and S. M. Ogbonmwan. Discussion: Jackknife, bootstrap and other resampling methods in regression analysis. Annals of Statistics, 14(4):1340–1343, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350160.

Wong:1983:CCV

[Won83] Wing Hung Wong. On the consistency of cross-validation in kernel nonparametric regression. Annals of Statistics, 11(4):1136–1141, December 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346327.

Wong:1986:TPL

[Won86] Wing Hung Wong. Theory of partial likelihood. Annals of Statistics, 14(1):88–123, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349844.

Woodroofe:1982:MSA

[Woo82] Michael Woodroofe. On model selection and the ARC sine laws. Annals of Statistics, 10(4):1182–1194, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345983.

Woodroofe:1985:ALM

[Woo85a] Michael Woodroofe. Asymptotic local minimaxity in sequential point estimation. Annals of Statistics, 13(2):676–688, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349547. See

correction [Woo89a].

Woodroofe:1985:EDF

[Woo85b] Michael Woodroofe. Estimating a distribution function with truncated data. Annals of Statistics, 13(1):163–177, March 1985. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346584. See

correction [Woo87].

Woodroofe:1986:VWE

[Woo86] Michael Woodroofe. Very weak expansions for sequential confidence levels. Annals of Statistics, 14(3):1049–1067, September 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350049.

Woodroofe:1987:CED

[Woo87] M. Woodroofe. Correction: “Estimating a Distribution Function with Truncated Data”. Annals of Statistics, 15(2):883, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176350383. See

[Woo85b].

Woodroofe:1989:CAL

[Woo89a] Michael Woodroofe. Correction: “Asymptotic Local Minimaxity in Sequential Point Estimation”. Annals of Statistics, 17(1):452, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347031. See [Woo85a].

Woodroofe:1989:VWE

[Woo89b] Michael Woodroofe. Very weak expansions for sequentially designed experiments: Linear models. Annals of Statistics, 17(3): 1087–1102, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347257.

Wright:1981:ABM

[Wri81a] F. T. Wright. The asymptotic behavior of monotone regression estimates. Annals of Statistics, 9(2):443–448, March 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345411.

Wright:1981:SRV

[Wri81b] F. T. Wright. Sums of random variables indexed by a partially ordered set and the estimation of integral regression functions. Annals of Statistics, 9(2):449–452, March 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345412.

Wright:1982:MRE

[Wri82] F. T. Wright. Monotone regression estimates for grouped observations. Annals of Statistics, 10(1):278–286, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345710.

Wang:1980:CER

[WS80] Y. H. Wang and R. C. Srivastava. A characterization of the exponential and related distributions by linear regression. Annals of Statistics, 8(1):217–220, January 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344905.

Wei:1986:TCR

[WSS86] L. J. Wei, R. T. Smythe, and R. L. Smith. K-treatment comparisons with restricted randomization rules in clinical trials. Annals of Statistics, 14(1):265–274, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349854.

Walster:1980:END

[WT80] G. William Walster and Marietta J. Tretter. Exact noncentral distributions of Wilks’ Λ and Wilks–Lawley U criteria as mixtures of incomplete beta functions: for three tests. Annals of Statistics, 8(6):1388–1390, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345210.

Woodroofe:1982:AEE

[WT82] Michael Woodroofe and Hajime Takahashi. Asymptotic expansions for the error probabilities of some repeated significance tests. Annals of Statistics, 10(3):895–908, September 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345879. See cor-

rection [WT85].

Woodroofe:1985:CAE

[WT85] Michael Woodroofe and Hajime Takahashi. Correction: “Asymptotic Expansions for the Error Probabilities of Some Repeated Significance Tests”. Annals of Statistics, 13(2):837, June 1985. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349565. See

[WT82].

Wu:1980:CCD

[Wu80] Chien-Fu Wu. Characterizing the consistent directions of least squares estimates. Annals of Statistics, 8(4):789–801, July 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345071.

Wu:1981:ATN

[Wu81a] Chien-Fu Wu. Asymptotic theory of nonlinear least squares estimation. Annals of Statistics, 9(3):501–513, May 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345455.

Wu:1981:RES

[Wu81b] Chien-Fu Wu. On the robustness and efficiency of some randomized designs. Annals of Statistics, 9(6):1168–1177, November 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345634.

Wu:1983:CPE

[Wu83] C. F. Jeff Wu. On the convergence properties of the EM algorithm. Annals of Statistics, 11(1):95–103, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346060.

Wu:1986:JBO

[Wu86a] C. F. J. Wu. Jackknife, bootstrap and other resampling methods in regression analysis. Annals of Statistics, 14(4):1261–1295, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.

aos/1176350142. See discussion [Tib88].

Wu:1986:RJB

[Wu86b] C. F. J. Wu. Rejoinder: Jackknife, bootstrap and other resampling methods in regression analysis. Annals of Statistics, 14(4): 1343–1350, December 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350161.

Wu:1986:LDR

[Wu86c] Tiee-Jian Wu. A large deviation result for signed linear rank statistics under the symmetry hypothesis. Annals of Statistics, 14(2):774–780, June 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349955.

Wu:1989:CDG

[Wu89] C. F. J. Wu. Construction of 2^{m}4^{n} designs via a grouping scheme. Annals of Statistics, 17(4):1880–1885, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347399.

Wright:1980:ICR

[WW80] Ian W. Wright and Edward J. Wegman. Isotonic, convex and related splines. Annals of Statistics, 8(5):1023–1035, September 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345140.

Wang:1983:FRC

[WW83] Song-Gui Wang and C. F. J. Wu. Further results on the consistent directions of least squares estimators. Annals of Statistics, 11(4):

1257–1262, December 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346339.

Wynn:1984:JKC

[Wyn84] Henry P. Wynn. Jack Kiefer’s contributions to experimental design. Annals of Statistics, 12(2):416–423, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346496.

Weerahandi:1983:EMB

[WZ83] S. Weerahandi and J. V. Zidek. Elements of multi-Bayesian decision theory. Annals of Statistics, 11(4):1032–1046, December 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176346319.

Yajima:1988:ERM

[Yaj88] Yoshihiro Yajima. On estimation of a regression model with longmemory stationary errors. Annals of Statistics, 16(2):791–807, June 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350837.

Yandell:1983:NIR

[Yan83] Brian S. Yandell. Nonparametric inference for rates with censored survival data. Annals of Statistics, 11(4):1119–1135, December 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346326.

Yao:1984:END

[Yao84] Yi-Ching Yao. Estimation of a noisy discrete-time step function: Bayes and empirical Bayes approaches. Annals of Statistics, 12(4): 1434–1447, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346802.

Yao:1987:ADM

[Yao87] Yi-Ching Yao. Approximating the distribution of the maximum likelihood estimate of the change-point in a sequence of independent random variables. Annals of Statistics, 15(3):1321–1328, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350509.

Yatracos:1985:RCM

[Yat85] Yannis G. Yatracos. Rates of convergence of minimum distance estimators and Kolmogorov’s entropy. Annals of Statistics, 13 (2):768–774, June 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349553.

Yatracos:1988:LBE

[Yat88] Yannis G. Yatracos. A lower bound on the error in nonparametric regression type problems. Annals of Statistics, 16(3): 1180–1187, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350954.

Yatracos:1989:RTP

[Yat89] Yannis G. Yatracos. A regression type problem. Annals of Statistics, 17(4):1597–1607, December 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347383.

Yohai:1980:CVO

[YB80] V. J. Yohai and M. S. Garcia Ben. Canonical variables as optimal predictors. Annals of Statistics, 8(4):865–869, July 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345079.

Ylvisaker:1987:PD

[Ylv87] Donald Ylvisaker. Prediction and design. Annals of Statistics, 15(1):1–19, March 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176350247.

Yohai:1987:HBP

[Yoh87] Victor J. Yohai. High breakdown-point and high efficiency robust estimates for regression. Annals of Statistics, 15(2):642–656, June 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176350366.

Yu:1989:IED

[Yu89] Qiqing Yu. Inadmissibility of the empirical distribution function in continuous invariant problems. Annals of Statistics, 17(3): 1347–1359, September 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347274.

Zabell:1982:WJS

[Zab82] Sandy L. Zabell. W. E. Johnson’s “sufficientness” postulate. Annals of Statistics, 10(4):1090–1099, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345975.

Zaman:1981:CCT

[Zam81] Asad Zaman. A complete class theorem for the control problem and further results on admissibility and inadmissibility. Annals of Statistics, 9(4):812–821, July 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345521.

Zehnwirth:1981:NAO

[Zeh81] Benjamin Zehnwirth. A note on the asymptotic optimality of the empirical Bayes distribution function. Annals of Statistics, 9 (1):221–224, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345352.

Zerdy:1980:RAO

[Zer80] Gloria C. Zerdy. Risk of asymptotically optimum sequential tests. Annals of Statistics, 8(5):1110–1122, September 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345148.

Zeytinoglu:1984:OFS

[ZM84] Mehmet Zeytinoglu and Max Mintz. Optimal fixed size confidence procedures for a restricted parameter space. Annals of Statistics, 12(3):945–957, September 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176346713.

Zeytinoglu:1988:RFS

[ZM88] Mehmet Zeytinoglu and Max Mintz. Robust fixed size confidence procedures for a restricted parameter space. Annals of Statistics, 16(3):1241–1253, September 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350958.